### v3

# Reliability, and Global Privacy

Abstract
This paper studies privacy-sensitive federated learning (FL) under unreliable communication, with a focus on secure aggregation and straggler mitigation. To preserve user privacy without compromising the utility of the global model, secure aggregation emerges as a promising approach by coordinating the use of privacy-preserving noise (secret keys) across participating clients. However, the unreliable communication will randomly disrupt the key coordination and disable the exact recovery of the global model in secure aggregation. Furthermore, unreliable communication can distort the optimization trajectory, causing the global model to deviate further from the intended global optimum. To address these challenges, we propose Secure Cooperative Gradient Coding (SecCoGC), a practical solution that achieves accurate aggregation with arbitrarily strong privacy guarantees and is inherently robust to communication uncertainties. SecCoGC operates natively in the real field, making it directly deployable in real-world FL systems. To ensure fairness in privacy protection, we further introduce Fair-SecCoGC, an extension of SecCoGC that enforces equitable privacy preservation across all clients. Notably, Fair-SecCoGC achieves optimal privacy under a per-key total power constraint. We formally formulate the problem of secure aggregation in the real field and present both general and computationally efficient methods for secret key construction. Our privacy analysis covers both Local Mutual Information Privacy (LMIP) and Local Differential Privacy (LDP) across all protocol layers, accounting for intermittent networks and correlation among secret keys. In addition, we characterize the system reliability and convergence properties of the proposed scheme. Experimental results demonstrate that SecCoGC achieves strong resilience to unreliable communication while maintaining arbitrarily strong privacy guarantees, yielding test accuracy improvements of $20\%$ to $70\%$ over existing privacy-preserving methods, without requiring any prior information.
Index Terms:
Federated learning, Unreliable communication, Secure Aggregation, Straggler Mitigation, Fairness, Local Mutual Information Privacy, Local Differential Privacy.
I Introduction
I-A Literature Review
Distributed learning (DL) mitigates the computational bottleneck in conventional centralized machine learning (ML) by leveraging edge device (client) capabilities. However, decentralization incurs significant communication overhead during the transmission of datasets. To overcome the communication bottleneck in DL, federated learning (FL) enables clients to collaboratively train models based on their local datasets, thus significantly reducing communication costs and inherently enhancing user privacy by keeping training dataset local [1, 2].
Despite the enhanced privacy provided by restricting raw data sharing, FL systems remain vulnerable to indirect model-based privacy leakage. Even without direct access to the raw datasets, adversaries can still infer sensitive information from the local model updates, making FL susceptible to various model-based attacks, such as inversion attacks (IA) and membership inference attacks (MIA). In IA, adversaries aim to reconstruct the original input from its observation. In MIA, adversaries attempt to infer whether a specific data patch is included in the training datasets by analyzing the model’s behavior. To defend against these and other attacks, a widely adopted strategy is to inject artificial privacy noise into the local models to obscure sensitive information, through e.g., the well-known Gaussian mechanism [3]. To quantify the information leakage of each client’s contribution, local mutual information privacy (LMIP) and local differential privacy (LDP) are commonly employed as effective privacy measures that capture the information disclosure of a mechanism from different perspectives [4, 5, 6]. While a number of studies [7, 8] confirm the effectiveness of Gaussian mechanisms with provable convergence guarantees, their practical applicability remains limited by the inherent utility-privacy trade-off. Strong privacy requires substantial noise, whose accumulation at the server can severely impair model accuracy and hinder convergence. Conversely, maintaining stable convergence and satisfactory training performance requires the injected noise to remain sufficiently small. This inherent tension creates a fundamental conflict between convergence and privacy in FL.
The aforementioned issue can be effectively addressed through secure aggregation [9, 10], in which privacy noises (secret keys) are designed with coordinated structures such that they can be canceled out during the server aggregation. This approach preserves privacy while maintaining convergence speed, making secure aggregation an appealing solution to privacy-sensitive FL applications, where both privacy and model utility are preserved. However, the effectiveness of secure aggregation highly relies on stable communication, as the random link disruptions will break the carefully designed coordination of secret keys required for accurate reconstruction of the global model. Recent advances in dropout-resilient secure aggregation primarily build upon Shamir’s Secret Sharing (SSS) [9, 11], and the use of maximum distance separable (MDS) random variables in secret key construction [12]. While these techniques offer perfect privacy by guaranteeing zero information leakage, they encounter several limitations in practical FL scenarios. These include high encoding complexity and the lack of a clear, non-trivial mapping between real-valued model parameters and finite-field arithmetic, raising concerns about their practicality in real-world deployments.
Another challenge posed by unreliable communication is the persistent, strict sub-optimality of federated optimization. As the data generated at the network edge is inherently uncontrollable and typically follows non-independent and identically distributed (non-IID) patterns across clients, local models tend to deviate significantly from the global objective, known as client drift[13]. As a result, the global model may converge in unstable or misaligned directions. Unreliable communication further amplifies this issue by causing statistically uneven local contributions, potentially leading the model to a stationary point far from the true optimum. This phenomenon, known as objective inconsistency [14], results in sub-optimality and undermines model reliability. The presence of privacy noise further intensifies the problem: The injected noise will enlarge the step size, and when combined with distorted optimization trajectories caused by unreliable communication, it can prevent the global model from approaching the true global optimum, and may even cause divergence.
Existing approaches to mitigate communication stragglers include heterogeneity-aware client sampling [15], adaptive resource allocation[16], heterogeneity-aware aggregation [17, 18], collaborative relaying [19, 20], etc. But they all rely on precise prior information to ensure optimality, thus increasing the implementation complexity in practical deployments. In contrast, [21] eliminates this need by appropriate accumulation mechanisms. Further improvements are made to handle unknown client participation statistics[22] and to adapt to non-stationary participation patterns[23]. However, these methods are prone to divergence, as the accumulated gradient after several consecutive transmission failures can grow excessively. Another line of work applies coding strategies for straggler mitigation, eliminating the need for prior knowledge. For example, [24] applies the MDS code to reduce the outage frequency in FL, but a consistent objective is not guaranteed in heterogeneous networks. In contrast, gradient coding (GC) has emerged as an effective approach that guarantees the optimality of FL without requiring prior knowledge of network conditions. Particularly, GC employs cyclic gradient codes to enable exact recovery of the global model despite arbitrary communication patterns [25]. Several variations of GC have been proposed to address various practical constraints and scenarios [26, 27, 28, 29]. However, these methods typically impose heavy computational burdens due to dataset replication across clients. To overcome this limitation, [30] proposed cooperative GC (CoGC), a gradient-sharing framework to compute the exact sum, as shown in Fig. 1. The CoGC greatly improves the communication-computation-efficiency and ensures convergence to the optimal solution over heterogeneous networks without requiring prior knowledge of network statistics.
Any two: $\bm{\Theta}_{1}+\bm{\Theta}_{2}+\bm{\Theta}_{3}$ $\mathcal{D}_{1}$$\mathcal{D}_{2}$$\mathcal{D}_{3}$$\frac{1}{2}\bm{\Theta}_{1}+\bm{\Theta}_{2}$$\bm{\Theta}_{2}-\bm{\Theta}_{3}$$\frac{1}{2}\bm{\Theta}_{1}+\bm{\Theta}_{3}$$\bm{\Theta}_{1}$$\bm{\Theta}_{2}$$\bm{\Theta}_{3}$
Figure 1: An illustration of the CoGC framework [31] with $K=3$ clients, where all communication links are unreliable. Each client $k$ computes its local model $\bm{\Theta}_{k}$ based on its local datasets $\mathcal{D}_{k}$, and shares $\bm{\Theta}_{k}$ with $s=1$ neighbor. Then, clients compute and transmit the partial sums to the server.
I-B Motivations
On one hand, CoGC still confronts unresolved privacy risks, especially during the gradient-sharing phase. On the other hand, the binary nature of the GC decoding mechanism provides a promising solution to the problem of unreliable communication in secure aggregation, as it allows either the exact recovery of the global model or nothing meaningful at all in the presence of random user dropouts. This property ensures that the structure of aggregation remains intact whenever the number of received updates reaches the threshold. Following this intuition, we arrive at the first fundamental research question explored in this paper:
[…]
    Is it possible to keep exact recovery of the global model while preserving privacy globally, despite unreliable communication? (C1)
Since the coding structure enforces the coefficients in the final sums, this challenge can be addressed by incorporating secure aggregation into the CoGC framework with the careful design of the additive secret keys. However, flexibility in secret key design introduces a new issue: inconsistent and uneven privacy protection among clients. This leads to the second key research question, concerning fairness in privacy preservation:
[…]
    How to design secret keys to ensure fair privacy preservation among clients in secure aggregation? (C2)
We seek to systematically address these questions by formally defining them and analyzing them from the perspective of secret key construction. The significance of our work is three-fold: (i) We ultimately resolve the model privacy issues in CoGC at all levels, while preserving its original advantages; (ii) Meanwhile, the proposed scheme offers an alternative secure aggregation approach that is resilient to random client dropout, without sophisticated encoding or mapping operations or multiple rounds of communication with the server; (iii) We formally formulate and address the secure aggregation problem for FL in the real field, making our solution practical and directly applicable. Notably, we introduce and formalize fairness as a real-field privacy metric, providing both a clear motivation and a principled foundation.
I-C Our Contributions
We apply rigorous theoretical analysis and extensive experimental validations to support our findings. The key contributions are detailed below.
(C1) To achieve both resilience and secure aggregation against unreliable communication, we propose a novel Secure CoGC scheme, termed SecCoGC. The SecCoGC guarantees optimal convergence performance in heterogeneous federated networks, enables exact reconstruction of the global model at the server, and provides arbitrarily strong user privacy globally. The construction conditions and implementation procedures are tailored to be directly compatible with the most existing FL frameworks. Furthermore, the proposed construction ensures privacy protection against any decoding mechanisms, providing robust security against various model attacks.
(C2) To address fairness concerns in privacy preservation, we propose a Fair and Secure CoGC (Fair-SecCoGC) scheme that ensures equitable privacy protection across all participating clients, achieves optimal privacy under a total power constraint, and retains the full benefits of the original SecCoGC. We establish rigorous theoretical foundations along with a general construction approach for fair secret keys. In addition, we present an extremely simplified yet effective construction method that requires almost no computation. To the best of our knowledge, this work is the first to identify user fairness as a privacy metric in FL.
(C3) To ensure a thorough evaluation of model privacy, we provide multi-view and in-depth theoretical privacy analysis for SecCoGC and Fair-SecCoGC, using both LMIP and LDP, across intermittent networks and throughout all stages of the training process. These results, in turn, ensure that the key constructions are both theoretically sound and practically deployable. Furthermore, our analysis of real-field secure aggregation, where privacy noises are correlated, offers valuable insights and a foundation for future research.
(C4) To ensure a tractable analysis of the proposed methods, we evaluate the system reliability of SecCoGC and Fair-SecCoGC under arbitrary network conditions. Building upon this, we analyze their convergence behavior with an arbitrary choice of local solver in non-convex settings. Our theoretical results yield a standard convergence rate of $\mathcal{O}\left(\frac{1}{\sqrt{T}}\right)$ with probability $99.74\%$ under arbitrary privacy levels. Notably, the convergence analysis also applies to CoGC and GC-based distributed training, while accommodating arbitrary local solvers for broader applicability.
I-D Paper Organization
The remainder of this paper is organized as follows. Section II introduces the system model and necessary preliminaries. Section III presents the formulation of secure aggregation in the real field together with the proposed SecCoGC scheme. Section IV extends this result by incorporating fairness concerns into the SecCoGC design. Privacy guarantees at each stage in FL are evaluated in Section V and Section VI, using LMIP and LDP, respectively. In Section VII, we analyze the system reliability of the proposed methods, while Section VIII examines its convergence behavior. The performance of the proposed methods is demonstrated through numerical comparisons with various benchmarks on multiple datasets in Section IX. Finally, conclusions are drawn in Section X, and supporting proofs are provided in the appendices.
II System Model and Preliminaries
II-A Machine Learning (ML) Model
Consider a distributed system consisting of $K$ clients, each holding local datasets $\mathcal{D}_{k}$ on it, coordinated by a central server. The goal is to learn a global model $\bm{\Theta}\in\mathbb{R}^{D}$ that minimizes the global objective function (GOF) defined in (1):
$\min_{\bm{\Theta}}\mathcal{L}(\bm{\Theta})\triangleq\sum_{k=1}^{K}w_{k}\mathcal{L}_{k}(\bm{\Theta},\mathcal{D}_{k}),$ (1)
through iterative training and communication. The learning weight $w_{k}$ measures the importance of clients, satisfying $\sum_{k=1}^{K}w_{k}=1$. W.L.O.G., this paper assumes $w_{k}=\frac{1}{K}$. The local objective function (LOF) that measures the accuracy of the model $\bm{\Theta}$ on local datasets $\mathcal{D}_{k}$ is defined by the average loss on all training samples in $\mathcal{D}_{k}$, that is,
$\mathcal{L}_{k}(\bm{\Theta},\mathcal{D}_{k})\triangleq\frac{1}{\lvert\mathcal{D}_{k}\rvert}\sum_{\xi\in\mathcal{D}_{k}}l(\bm{\Theta},\xi),$ (2)
where $l(\bm{\Theta},\xi)$ is the sample loss of the model $\bm{\Theta}$ evaluated on a single training sample $\xi$.
In the $t$-th training round, where $t\in[T]$, all clients are initialized with the latest global model, i.e., $\bm{\Theta}_{k}^{t,0}=\bm{\Theta}^{t-1}$. The local training on client $k$ at the $i$-th iteration is given by
$\bm{\Theta}_{k}^{t,i}=\bm{\Theta}_{k}^{t,i-1}-\eta a_{i}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t,i-1},\bm{\xi}_{k}^{t,i}),i\in[I]$ (3)
where $\eta$ is the learning rate, $I$ is the number of local iterations, $a_{i}$ characterizes the type of local solver, $\bm{\xi}_{k}^{t,i}$ is the data patch containing $n_{p}$ data samples extracted from $\mathcal{D}_{k}$ with $\lvert\bm{\xi}_{k}^{t,i}\rvert=n_{p}$. The gradient $\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t,i-1},\bm{\xi}_{k}^{t,i})$ computed on the data patch $\bm{\xi}_{k}^{t,i}$ is defined by the averaged gradients over all data samples $\xi\in\bm{\xi}_{k}^{t,i}$. In matrix form, the local model update $\Delta\bm{\Theta}_{k}^{t,I}=\bm{\Theta}_{k}^{t,I}-\bm{\Theta}^{t}$ can be expressed by
[…]
With full participation of clients, the server updates $\bm{\Theta}^{t}$ as
$\bm{\Theta}^{t}=\bm{\Theta}^{t-1}+\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t,I}.$ (5)
II-B Communication Model
The inter-user and uplink communication channels are assumed to be orthogonal and modeled as independent Bernoulli r.v.s. Specifically, the link from client $k$ to the sever, denoted by $\tau_{k}$, is modeled by $\tau_{k}\sim\mathrm{Ber}(1-p_{k})$, the link from client $m$ to client $k$, denoted by $\tau_{k,m}$ is modeled by $\tau_{k,m}\sim\mathrm{Ber}(1-p_{k,m})$, where $p_{k}$ and $p_{k,m}$ are the probabilities when the channel is in outage. The $\tau_{k}=1$ and $\tau_{k,m}=1$ implies a successful communication and perfect recovery, instead, $\tau_{k}=0$ and $\tau_{k,m}=0$ implies a complete failure. For all $k$, $p_{k,k}=0$, $\tau_{k,k}=1$, since client $k$ has access to its own local model and requires no transmission. For $k\neq m$, if no transmission occurs from client $m$ to client $k$, then $p_{k,m}=1$, $\tau_{k,m}=0$, as clients do not have direct access to each other’s local models. Moreover, any link is statistically independent from each other, i.e., $\mathrm{(i)}\;\forall k_{1}\neq k_{2}:\tau_{k_{1}}\perp\tau_{k_{2}}$; $\mathrm{(ii)}\;\forall(k_{1},m_{1})\neq(k_{2},m_{2}):\tau_{k_{1},m_{1}}\perp\tau_{k_{2},m_{2}}$; and $\mathrm{(iii)}\;\forall k_{1},k_{2},m:\tau_{k_{1}}\perp\tau_{k_{2},m}$.
II-C Cyclic GC Codes
The GC codes consists of pair-wise cyclic matrices $\bm{G}$ and $\bm{C}$ that are carefully devised such that $\bm{C}\bm{G}=\mathbf{1}_{f\times K}$, where $f={K\choose s}$. In particular, $\bm{G}$ is the allocation matrix of size $K\times K$ in a cyclic pattern with $s$ nonzero entries per row, and $\bm{C}$ is the combination matrix of size $f\times K$, where each row represents a unique pattern with exactly $s$ zeros per row, encompassing all such possible patterns. The structure ensures that, even if up to $s$ rows are missing from $\bm{G}$, its resulting product with $\bm{C}$ still includes at least one all-one row. The construction for GC codes can be found in [25, Algo. II].
II-D CoGC Framework
Instead of replicating datasets across clients according to the nonzero patterns in $\bm{G}$ [25, Algo. II], CoGC enables gradient sharing among clients. Let $g_{k,m}$ denote the $(k,m)$-th element in $\bm{G}$, the following phases are carried out:
[…]
    Gradient-Sharing Phase: Client $k$ shares its local model update $\Delta\bm{\Theta}_{k}^{t,I}$ with its neighbors in $\mathcal{V}_{k}=\{m:g_{m,k}\neq 0\}$ (including client $k$ itself).
[…]
    Relaying: Client $k$ should listen to its neighbors in $\mathcal{U}_{k}=\{m:g_{k,m}\neq 0\}$ (including client $k$ itself). If client $k$ successfully receives $\Delta\bm{\Theta}_{m}^{t,I}$ from client $m$, it computes the weighted update $g_{k,m}\Delta\bm{\Theta}_{m}^{t,I}$ and aggregates these to form partial sums. If client $k$ successfully hears from all of its $s$ neighbors, the aggregated partial sum is called complete partial sum, otherwise, it is incomplete.
[…]
    Server Aggregation: If the server receives at least KsK-s complete partial sums, it can reconstruct the exact global model in (5) without any distortion, by combining these complete partial sums according to $\bm{C}$.
More details about CoGC can be found in [31].
III The Proposed Method: Secure Cooperative Gradient Coding (SecCoGC)
III-A Method
Consider $K$ clients, where each client $k\in\{1,\cdots,K\}\triangleq[K]$ holds a raw local update $\Delta\bm{\Theta}_{k}^{t,I}$ and a secret key $\bm{N}_{k}^{t}$. To protect data against exposure during transmission, each client locally obscures its raw local model update by applying additive masking, i.e.,
$\bm{Y}_{k}^{t}=\Delta\bm{\Theta}_{k}^{t,I}+\bm{N}_{k}^{t}.$ (6)
Then, each client $k$ transmits $\bm{Y}_{k}^{t}$ to its neighbors $m\in\mathcal{V}_{k}$ through unreliable link $\tau_{m,k}^{t}$.
In the meanwhile, client $k$ also acts as a relay, listening to its neighboring client $m\in\mathcal{U}_{k}$. Client $k$ may or may not be able to receive information from client $m$, depending on the realization of link $\tau_{k,m}^{t}$, where
[…]
Let $\mathcal{U}_{k}^{t}$ denote the set of messages that client $k$ can successfully decode at the $t$-th round, i.e., $\mathcal{U}_{k}^{t}=\{m:\tau_{k,m}^{t}=1\}$. Then, client $k$ computes the product $g_{k,m}\bm{Y}_{m}^{t}$ for each $m\in\mathcal{U}_{k}^{t}$, and sums these products to compute the partial sum, i.e.,
$\bm{S}_{k}^{t}=\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{Y}_{m}^{t}=\sum_{m=1}^{K}\tau_{k,m}^{t}\cdot g_{k,m}\bm{Y}_{m}^{t}.$ (8)
If $\mathcal{U}_{k}^{t}=\mathcal{U}_{k}$, meaning that client $k$ successfully decodes the message from all its neighbors, then $\bm{S}_{k}^{t}$ is referred to as a complete partial sum, since the coefficients corresponding to each neighbor’s contribution in $\bm{S}_{k}^{t}$ fully match those in $\bm{G}$.
Then, client $k$ sends $\bm{S}_{k}^{t}$ to the server through link $\tau_{k}^{t}$, where
[…]
Let $\mathcal{U}^{t}$ denote the set of complete partial sums arrived at the server, i.e., $\mathcal{U}^{t}=\{k:\mathcal{U}_{k}^{t}=\mathcal{U}_{k}\}\cap\{k:\tau_{k}^{t}=1\}$, the server aggregation can be summarized below.
[…]
    Combinator: Let $\bm{c}_{f_{t}}$ denote the $f_{t}$-th row of the combination matrix $\bm{C}$. The combinator $\bm{c}_{f_{t}}$ corresponds to $\mathcal{U}^{t}$ should satisfy
    $\forall k:\;c_{f_{t},k}\neq 0\Rightarrow k\in\mathcal{U}^{t}.$ (10)
    Such $\bm{c}_{f_{t}}$ exists iff $\lvert\mathcal{U}^{t}\rvert\geq K-s$.
[…]
    Server Aggregation: If such $\bm{c}_{f_{t}}$ exists, the server can compute the global model by combining these complete partial sums according to $\bm{c}_{f_{t}}$, i.e.,
    $\bm{\Theta}^{t}$ $=\bm{\Theta}^{t-1}+\frac{1}{K}\sum_{k=1}^{K}c_{f_{t},k}\bm{S}_{k}^{t}.$ (11)
Any two: $\bm{\Theta}_{1}+\bm{\Theta}_{2}+\bm{\Theta}_{3}$ $\mathcal{D}_{1}$$\mathcal{D}_{2}$$\mathcal{D}_{3}$ $\frac{1}{2}\bm{\Theta}_{1}+\bm{\Theta}_{2}$ $\frac{1}{2}\bm{N}_{1}-\bm{N}_{2}$ $\bm{\Theta}_{2}-\bm{\Theta}_{3}$ $\bm{N}_{2}-\bm{N}_{3}$ $\frac{1}{2}\bm{\Theta}_{1}+\bm{\Theta}_{3}$ $\frac{1}{2}\bm{N}_{1}+\bm{N}_{3}$ $\bm{\Theta}_{1}+\bm{N}_{1}$$\bm{\Theta}_{2}+\bm{N}_{2}$$\bm{\Theta}_{3}+\bm{N}_{3}$
Figure 2: An illustrative overview of the proposed secure CoGC framework with simplified notations for clarity.
Core Idea: If the secret keys $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ are designed with a special structure such that
$\sum_{k=1}^{K}\bm{N}_{k}^{t}=0,$ (12)
it can be verified that the server aggregation in (11) gives exactly the global model pursued by (5) without introducing any additional ambiguity, as depicted in Fig. 2, i.e.,
$\frac{1}{K}\sum_{k=1}^{K}c_{f_{t},k}\bm{S}_{k}^{t}$ $=\frac{1}{K}\sum_{k=1}^{K}\left(\Delta\bm{\Theta}_{k}^{t,I}+\bm{N}_{k}^{t}\right)=\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t,I}.$ (13)
During transmission and aggregation, no third party can accurately infer any individual local model. In SecCoGC, if the server receives at least KsK-s complete partial sums, it can fully eliminate the privacy noise and accurately reconstruct the global model. Otherwise, clients proceed to the next round using their latest local models by setting $\bm{\Theta}_{k}^{t+1,0}=\bm{\Theta}_{k}^{t,I}$.
Training terminates when the total number of executed rounds reaches $T$, with the final round resulting in a successful reconstruction of the global model.
III-B Secret Key Generation
Let us define $L$ noise components $\bm{Z}_{l}^{t}\sim\mathcal{N}(\mathbf{0},\bm{I}_{D})$,  [^1] The noise components $\bm{Z}_{l}^{t}$ are not necessarily centered at 0. Here, we assume zero-mean for the convenience of theoretical derivations. where $l\in[L]$ and $\bm{I}_{D}$ is the identity matrix of size $D\times D$. On each client $k$, the secret key is constructed as a combination of these noise components, i.e., $\bm{N}_{k}^{t}=\sum_{l=1}^{L}\alpha_{k,l}\bm{Z}_{l}^{t}$. Let $\bm{A}$ (of size $K\times L$) denote the generator matrix comprising all $\alpha_{k,l}$. To ensure non-zero secret keys, any row in $\bm{A}$ should contain at least one non-zero entry, i.e., $\forall k$, $\exists l:\alpha_{k,l}\neq 0$. In matrix form, the secret keys on all clients are generated by
$\begin{bmatrix}\bm{N}_{1}^{t}\\ \vdots\\ \bm{N}_{K}^{t}\end{bmatrix}=\bm{A}\cdot\begin{bmatrix}\bm{Z}_{1}^{t}\\ \vdots\\ \bm{Z}_{L}^{t}\end{bmatrix}$. (14)
Next, we begin by reformulating the problem of secure aggregation to accommodate the real field, and then specify the corresponding constraints on the generator matrix $\bm{A}$.
Condition III.1 (Secure Aggregation in $\mathbb{R}$).
First, the secure aggregation scheme must ensure the correct computation of the global model, as specified in (12). This requirement further implies that the sum of the elements in each column of the generator matrix $\bm{A}$ must be zero, i.e.,
$[\mathrm{Correctness}]\;\;\;\;\;\;\forall l\in[L]:\sum_{k=1}^{K}\alpha_{k,l}=0.$ (15)
Second, any linear combination involving fewer than KK(an adjustable threshold) secret keys must be non-trivial, i.e., the zero vector must be excluded from the span of any such subset so that no combination of fewer than $K$ secret keys cancels out to zero. This requires
$[\mathrm{Security}]\;\;\;\;\;\;\;\;\;\mathrm{rank}\left(\bm{A}\right)=K-1.$ (16)
This condition guarantees that partial access to fewer than $K$ clients’ information does not undermine security. Note that K1K-1 is the largest possible rank of $\bm{A}$ to satisfy the correctness condition. The rationale behind this threshold $K$ is further discussed below.
Remark III.1 (Necessity of the Security Condition).
The security condition in (16) can be overly strict, as the GC decoding only relies on certain linear combinations of partial sums, while other combinations with incorrect coefficients cannot reveal meaningful information about the underlying model. However, under alternative GC decoding schemes, such as the one proposed in [31], condition (16) remains critical, as it offers privacy guarantees to each individual local model that hold regardless of the employed decoding strategy.
We note that the real-field addition operation cannot give perfect privacy, as the masked version $\Delta\bm{\Theta}_{k}^{t,I}+\bm{N}_{k}^{t}$ cannot be fully independent from the original message $\Delta\bm{\Theta}_{k}^{t,I}$. So for any distributions of $\Delta\bm{\Theta}_{k}^{t,I}$ and $\bm{N}_{k}^{t}$, the mutual information is non-zero, i.e., $\forall P_{\Delta\bm{\Theta}_{k}^{t}},P_{\bm{N}_{k}^{t}}:\mathrm{I}(\Delta\bm{\Theta}_{k}^{t,I};\Delta\bm{\Theta}_{k}^{t,I}+\bm{N}_{k}^{t})>0$. This property differs fundamentally from that of finite fields, where underlie most existing secure aggregation schemes. For this reason, we adopt $\mu$-local mutual information privacy ($\mu$-LMIP)[5] to measure the privacy level provided by a real-field secure aggregation scheme. The $\mu$-LMIP aligns well with the core objectives of secure aggregation.
Definition III.1 ($\mu$-LMIP).
For a distribution space $P_{\bm{\Theta}}$, a randomized mechanism $\mathcal{M}$ satisfies $\mu$-LMIP iff
$\mathrm{I}(\bm{\Theta};\mathcal{M}(\bm{\Theta}))\leq\mu\;\;\mathrm{bits}.$ (17)
Here, the mechanism $\mathcal{M}$ represents each procedure in SecCoGC, as will be detailed in Section V.
Next, we establish the necessary constraints such that $\bm{A}$ satisfying Condition III.1 exists. The matrix $\bm{A}$ contains $KL$ entries, and (15) imposes $L$ linear constraints. Thus, if $KL\geq L$, which holds for all $K$ and $L$, the system has non-zero solution. Additionally, (16) introduces an extra constraint, requiring $L\geq K-1$. These lead us to the following lemma.
Lemma III.1 (Existence Condition).
A generator matrix $\bm{A}\in\mathbb{R}^{K\times L}$ that gives distinct secret keys satisfying Condition III.1 exists iff
$L\geq K-1.$ (18)
Construction III.1.
Condition III.1 suggests a straightforward method for constructing the generator matrix $\bm{A}$, as outlined below.
[…]
    Step 1: For the first K1K-1 rows in $\bm{A}$, each entry $\alpha_{k,l},k\in\{1,\cdots,K-1\}$ is generated randomly in $\mathbb{R}$ for $\forall l$.
[…]
    Step 2: For the last row, set $\alpha_{K,l}=-\sum_{k=1}^{K-1}\alpha_{K,l}$ for $\forall l$.
The first K1K-1 secret keys are independently generated with full freedom, while the last key is fully dependent on them. Following this, the nature of random generation in $\mathbb{R}$ will ensure the rank condition in (16) with probability $1$.
Example III.1.
Suppose there are $K=5$ clients participating in the training, and $L=K$ independent noise components $\bm{Z}_{l}^{t}\sim\mathcal{N}(\mathbf{0},\bm{I}_{D})$ are used to generate the secret keys. Following Construction III.1, an example of $\bm{A}$ is given below.
$\bm{A}=\begin{bmatrix}1.41&-0.80&0.21&0.72&-0.08\\ 0.29&0.69&-1.16&2.58&-1.93\\ 0.19&0.83&-1.14&-0.66&-0.43\\ 1.58&-0.24&0.10&0.18&-1.79\\ -3.47&-0.48&1.99&-2.82&4.24\\ \end{bmatrix}$. (19)
The secret keys are further generated according to (14). It can be verified that the secret keys satisfy Condition III.1.
III-C Merits and Drawbacks
By inheriting the CoGC framework, SecCoGC maintains computational efficiency and ensures optimality in heterogeneous networks. Moreover, it provides global privacy across all levels of the learning protocol. Specifically, SecCoGC protects against: (i) peer inference attacks, (ii) relay interception, and (iii) server-side inference attacks. Importantly, SecCoGC enables the exact recovery of the global model under unreliable communication, which ensures the model’s utility.
However, Construction III.1 often leads to variations in secret key power. On one hand, signals with very low or high magnitudes are susceptible to distortion or clipping by the typically nonlinear power amplifier, resulting in inaccurate and noisy aggregation. On the other hand, the uneven randomness introduced into local models can lead to unequal levels of privacy preservation across clients. Since learning weights are determined by default (training data), unrelated to factors, e.g., communication and privacy, such disparities in the power of secret keys are unjustified and could raise fairness concerns. The following example concretely illustrates this issue.
Example III.2 (Fairness Issue).
Following Construction III.1, the last key is generated by $\bm{N}_{5}^{t}=-\sum_{k=1}^{4}\bm{N}_{k}^{t}$, so the power of $\bm{N}_{5}^{t}$ can be accumulated significantly. Following Example III.1, we have
$\bm{N}_{1}^{t}\sim\mathcal{N}(\mathbf{0},3.19\bm{I}_{D}),\;\bm{N}_{2}^{t}\sim\mathcal{N}(\mathbf{0},12.28\bm{I}_{D}),$
$\bm{N}_{3}^{t}\sim\mathcal{N}(\mathbf{0},2.64\bm{I}_{D}),\;\bm{N}_{4}^{t}\sim\mathcal{N}(\mathbf{0},5.80\bm{I}_{D}),$
$\bm{N}_{5}^{t}\sim\mathcal{N}(\mathbf{0},42.30\bm{I}_{D}).$
The last client receives privacy noise with power 3163-16 times higher than others, resulting in disproportionately stronger privacy protection.
IV Toward Fair and Secure Aggregation: Fair-SecCoGC
The condition for fair privacy preservation requires that all secret keys have identical power, i.e., $\forall k\in[K]:\bm{N}_{m}^{t}\sim\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})$. This can be alternatively formulated as follows.
Condition IV.1 (Fairness Condition).
The sum of squared $\ell_{2}$-norm of rows in $\bm{A}$ are identical, i.e.,
$[\mathrm{Fairness}]\;\;\;\;\;\;\forall k\in[K]:\lVert\underline{\bm{\alpha}}_{k}\rVert_{2}^{2}=\lambda^{2},$ (20)
where $\underline{\bm{\alpha}}_{k}$ is the $k$-th row in $\bm{A}$ and $\lambda>0$ is a constant.
The fairness condition in (20) imposes $K$ additional constraints. The solution of $\bm{A}$ satisfying (15) and (20) is non-trivial, if $KL\geq K+L$, which requires $K\geq 2$, $L\geq 2$. Combined with Lemma III.1, we have the following lemma.
Lemma IV.1 (Existence Condition of Fair $\bm{A}$).
A generator matrix $\bm{A}\in\mathbb{R}^{K\times L}$ that gives distinct secret keys and satisfies both the fairness condition (Condition IV.1) and the secure condition (Condition III.1) iff
$L\geq K-1.$ (21)
Furthermore, If $\bm{A}$ is of cyclic structure, the density of the non-zero elements per row is bounded by the following lemma.
Lemma IV.2 (Lower Bound on Density).
If $\bm{A}$ is of cyclic structure, the number of nonzero entries per row should satisfy
$\forall k:\;\lVert\bm{\alpha}_{k}\rVert_{0}\geq 2.$ (22)
Next, we proceed to key construction. To achieve fair secret key generation, the design must simultaneously satisfy the column-wise linear constraints required for security (Condition III.1), and the row-wise $\ell_{2}$-norm constraints required for fairness (Condition IV.1). This dual requirement poses a challenge to code construction, as it necessitates the joint design of all rows and columns, rather than treating them separately. Below, two construction methods are presented.
Construction IV.1 (General Construction).
With the adjustable parameter $\lambda^{2}$ controlling the privacy noise power and $L=K$, the following construction generates fair secret keys.
[…]
    Step 1: For $\forall k\in[K]$, $\forall l\neq k,l\neq(k+1)\mod L$, $\alpha_{k,l}$ is generated randomly with full freedom.
[…]
    Step 2: The remaining $2K$ elements $\alpha_{k,l}$, where $\forall k\in[K],l=k$ and $l=(k+1)\mod L$, are fully determined by solving the following $2K$ constraints imposed by the correctness condition and fairness condition, i.e.,
    $\forall l:\sum_{k=1}^{K}\alpha_{k,l}=0,\text{and}\;\forall k:\sum_{l=1}^{L}\alpha_{k,l}^{2}=\lambda^{2}.$ (23)
The (23) can be solved by e.g., the substitution method, the Gröbner Basis Method, etc. However, solving these equations is computationally complex in large systems. To overcome this limitation, we propose a simplified construction of the generator matrix with a cyclic structure, which produces fair and secure secret keys without solving any equations. The core idea is to set the off-diagonal elements identical first, and then compute the diagonal so that the $\ell_{2}$-norm across rows is enforced to be equal. In this case, adjustments of rows do not affect their linear relationships across columns. Under both Construction IV.1 and IV.2, the secret keys are correlated.
Construction IV.2 (Simplified Construction).
Let $L=K$ and each row contains $\gamma\geq 1$ (Lemma IV.2) off-diagonal non-zeros. The following construction of $\bm{A}$ yields fair secret keys.
[…]
    Step 1: For $\forall k\in[K]$, $\forall l=(k+l)\mod K$, $l\in\{1,\cdots,\gamma\}$, set the off-diagonal nonzero elements to
    $\alpha_{k,l}=\frac{\lambda}{\sqrt{\gamma^{2}+\gamma}}.$ (24)
[…]
    Step 2: For $\forall k\in[K]$, $\forall l=k$, the diagonal is set according to the sum of each column, i.e.,
    $\alpha_{k,k}=-\frac{\gamma\lambda}{\sqrt{\gamma^{2}+\gamma}}.$ (25)
    So far, the executed procedures have enforced equal $\ell_{2}$-norm across rows and ensured the secure condition. So for each client $k$, the privacy noise $\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{d})$.
Example IV.1.
With the same setting in Example III.1, set $\gamma=2$, and $\lambda^{2}=6$, Construction IV.2 gives the following generator matrix,
$\bm{A}=\begin{bmatrix}-2&1&1&0&0\\ 0&-2&1&1&0\\ 0&0&-2&1&1\\ 1&0&0&-2&1\\ 1&1&0&0&-2\\ \end{bmatrix}$. (26)
It can be verified that $\forall k:\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},6\bm{I}_{D})$ and that $\sum_{k=1}^{5}\bm{N}_{k}^{t}=\mathbf{0}$, ensuring fair privacy preservation while achieving exact recovery of the global model.
V Local Mutual Information Privacy (LMIP)
This section examines LMIP of each local model update in SecCoGC at all levels in the learning process. At the $t$-th round, both clients and the server are aware of the previous global model $\bm{\Theta}^{t}$, knowing $\Delta\bm{\Theta}_{k}^{t,I}$ is equivalent to knowing $\bm{\Theta}_{k}^{t,I}$. For simplicity, we omit the index $I$ in the notation below.
V-1 Peer-to-Peer LMIP
This section aims to quantify the maximum information leakage during the gradient-sharing phase, when the best secret keys are used. The clients are assumed to be honest but curious, they are honest because they faithfully perform their role in the system. But they can be interested in learning sensitive information about other clients. We assume that each secret key $\bm{N}_{k}^{t}\in\mathbb{R}^{D}$ satisfies a total power constraint, such that $\operatorname{Tr}(\operatorname{Cov}[\bm{N}_{k}^{t}])\leq D\lambda^{2}$. Similarly, we assume that each local update $\Delta\bm{\Theta}_{k}^{t}\in\mathbb{R}^{D}$ satisfies a total power constraint of $D\zeta^{2}$, such that $\operatorname{Tr}(\operatorname{Cov}[\Delta\bm{\Theta}_{k}^{t}])\leq D\zeta^{2}$. The local model updates $\{\Delta\bm{\Theta}_{k}^{t}\}_{k=1}^{K}$ are assumed to be mutually independent since local training is performed independently by each client.
Theorem V.1 (Peer-to-Peer LMIP).
In both SecCoGC and Fair-SecCoGC, the transmission from client $k$ to $m$ satisfy $\mu_{1}$-LMIP, where
$\mu_{1}=(1-p_{m,k})\frac{D}{2}\log\left(1+\frac{\zeta^{2}}{\lambda^{2}}\right).$ (27)
Proof.
The maximal information leakage provided by the best privacy noise can be formulated as follows,
$\adjustlimits{\sup}_{\Delta\bm{\Theta}_{k}^{t}}{\inf}_{\bm{N}_{k}^{t}}\mathrm{I}(\Delta\bm{\Theta}_{k}^{t};\tau_{m,k}(\Delta\bm{\Theta}_{k}^{t}+\bm{N}_{k}^{t})|\tau_{m,k})$ (28a)
$=$ $\adjustlimits{\sup}_{\Delta\bm{\Theta}_{k}^{t}}{\inf}_{\bm{N}_{k}^{t}}(1-p_{m,k})\mathrm{I}(\Delta\bm{\Theta}_{k}^{t};\tau_{m,k}(\Delta\bm{\Theta}_{k}^{t}+\bm{N}_{k}^{t})|\tau_{m,k}=1)$ (28b)
$=$ $\adjustlimits{\sup}_{\Delta\bm{\Theta}_{k}^{t}}{\inf}_{\bm{N}_{k}^{t}}(1-p_{m,k})\left(\mathrm{h}(\Delta\bm{\Theta}_{k}^{t}+\bm{N}_{k}^{t})-\mathrm{h}(\bm{N}_{k}^{t})\right)$ (28c)
$=$ $\sup_{\Delta\bm{\Theta}_{k}^{t}}(1-p_{m,k})\left(\mathrm{h}(\Delta\bm{\Theta}_{k}^{t}+\bm{N}_{k}^{t})-\mathrm{h}\left(\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})\right)\right)$ (28d)
$=$ $(1-p_{m,k})\left(\mathrm{h}\left(\mathcal{N}(\mathbf{0},(\lambda^{2}+\zeta^{2})\bm{I}_{D})\right)-\mathrm{h}\left(\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})\right)\right)$ (28e)
$=$ $(1-p_{m,k})\frac{D}{2}\log\left(1+\frac{\zeta^{2}}{\lambda^{2}}\right),$ (28f)
where (28b) owes to fact that if $\tau_{m,k}=0$, then client $m$ learns nothing about client $k$, i.e., $\mathrm{I}(\Delta\bm{\Theta}_{k}^{t};\tau_{m,k}(\Delta\bm{\Theta}_{k}^{t}+\bm{N}_{k}^{t})|\tau_{m,k}=0)=0$, (28c) is due to the independence between $\{\Delta\bm{\Theta}_{k}^{t}\}_{k=1}^{K}$ and $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$, (28d) and (28e) follows from the fact that, under a total power (trace) constraint, the maximum differential entropy is achieved by a Gaussian vector with a diagonal covariance matrix whose diagonal entries are equal. ∎
The above results lead to two key conclusions. First, Fair-SecCoGC not only ensure fair privacy protection among clients, but also produces optimal keys $\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})$ under a total power constraint of $D\lambda^{2}$, in the sense that it offers the smallest information leakage. Therefore, in the analysis below, we employ these optimal secret keys for all $k$. Second, any other input source than a Gaussian vector with a diagonal covariance matrix $\zeta^{2}\bm{I}_{D}$ results in lower information leakage.
V-2 LMIP at Relay
This section examines the information leakage in the partial sums when a client acts as a central relay. The secret keys $\left\{\bm{N}_{k}^{t}\right\}_{k=1}^{K}$ are correlated across $k$, due to the coordinated nature of secret keys in secure aggregation. The server is assumed to follow the protocol faithfully but may attempt to infer sensitive information about $\Delta\bm{\Theta}_{k}^{t}$. In the strictest case, where the server has access to the full connectivity realizations $\bm{\mathcal{T}}:=[\tau_{m,k}^{t}]_{K\times K}$ among clients, the information leakage for each local update, contained in the observed partial sums (complete or incomplete), is characterized below.
Theorem V.2 (LMIP in Masked Partial Sums).
In Fair-SecCoGC, the partial sums computed by client $k$, in protecting client $k_{1}$’s contribution, satisfy $\mu_{2}$-LMIP, where
$\mu_{2}=(1-p_{k,k_{1}})\frac{D}{2}\cdot$
$\left(\log\left((2\pi e)g_{k,k_{1}}^{2}\zeta^{2}\right)+\log\left(\frac{\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}}{\lVert\bm{\Lambda}_{k}^{t,-k_{1}}\bm{A}\rVert^{2}}\right)\right).$ (29)
Proof.
The server receives the partial sum $\bm{S}_{k}^{t}$ from client $k$ and wishes to infer sensitive information about client $k_{1}$. The information leakage can be expressed by
$\mathrm{I}\left(\Delta\bm{\Theta}_{k_{1}}^{t};\bm{S}_{k}^{t}|\bm{\mathcal{T}},\bm{G}\right)$
$=$ $\mathrm{I}\left(\Delta\bm{\Theta}_{k_{1}}^{t};\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\left(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t}\right)\bigg{|}\bm{\mathcal{T}},\bm{G}\right)$
$=$ $\mathrm{h}\left(\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\left(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t}\right)\bigg{|}\bm{\mathcal{T}},\bm{G}\right)$
$\hskip 22.76219pt-\mathrm{h}\left(\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\left(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t}\right)\bigg{|}\Delta\bm{\Theta}_{k_{1}}^{t},\bm{\mathcal{T}},\bm{G}\right)$
$=$ $\mathrm{h}\left(\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\left(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t}\right)\bigg{|}\bm{\mathcal{T}},\bm{G}\right)$
$\hskip 22.76219pt-\mathrm{h}\left(\sum_{m\neq k_{1}}\bm{\Lambda}_{k}^{t}(m)\left(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t}\right)\bigg{|}\bm{\mathcal{T}},\bm{G}\right),$ (30)
where $\bm{\Lambda}_{k}^{t}(m)=\tau_{k,m}g_{k,m}$ is the actually received coefficient by client $k$ from client $m\in[K]$. The (30) follows that $\bm{\Lambda}_{k}^{t}(k_{1})$ is fully determined by $\bm{\mathcal{T}}$, and $\bm{G}$ and that $\Delta\bm{\Theta}_{k_{1}}^{t}$ is independent of $\{\Delta\bm{\Theta}_{k}^{t}\}_{k\neq k_{1}}$, $\bm{\mathcal{T}}$, and $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$. Due to Guassian distribution and independence between $\{\Delta\bm{\Theta}_{k}^{t}\}_{k=1}^{K}$ and $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$, the differential entropy in (30) becomes separable,
$\mathrm{h}\left(\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\Delta\bm{\Theta}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\hskip-2.84526pt\right)\hskip-2.84526pt+\mathrm{h}\left(\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\hskip-2.84526pt\right)$
- $\mathrm{h}\left(\hskip-2.84526pt\sum_{m\neq k_{1}}\hskip-2.84526pt\bm{\Lambda}_{k}^{t}(m)\Delta\bm{\Theta}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\hskip-2.84526pt\right)\hskip-2.84526pt-\mathrm{h}\hskip-2.84526pt\left(\hskip-2.84526pt\sum_{m\neq k_{1}}\hskip-2.84526pt\bm{\Lambda}_{k}^{t}(m)\bm{N}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\hskip-2.84526pt\right).$ (31)
Due to Guassian distribution and independence among $\Delta\bm{\Theta}_{k}^{t}$ for all $k$ and independent communications, the sum can be taken out, (31) further becomes
$\sum_{m=1}^{K}\mathrm{h}\left(\bm{\Lambda}_{k}^{t}(m)\Delta\bm{\Theta}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\right)+\mathrm{h}\left(\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\right)$
$-\hskip-5.69054pt\sum_{m\neq k_{1}}\hskip-2.84526pt\mathrm{h}\left(\bm{\Lambda}_{k}^{t}(m)\Delta\bm{\Theta}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\hskip-2.84526pt\right)\hskip-2.84526pt-\mathrm{h}\hskip-2.84526pt\left(\sum_{m\neq k_{1}}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\right)$
$=\mathrm{h}\left(\bm{\Lambda}_{k}^{t}(k_{1})\Delta\bm{\Theta}_{k_{1}}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\right)+\mathrm{h}\left(\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\right)$
$-\mathrm{h}\left(\sum_{m\neq k_{1}}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\right).$ (32)
Let $\bm{\Lambda}_{k}^{t}=[\bm{\Lambda}_{k}^{t}(1),\cdots,\bm{\Lambda}_{k}^{t}(K)]$ stores all the actually received coefficients at client $k$, it further holds that
$\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{k}^{t}=\bm{\Lambda}_{k}^{t}\begin{bmatrix}\bm{N}_{1}^{t}\\ \vdots\\ \bm{N}_{K}^{t}\end{bmatrix}=\bm{\Lambda}_{k}^{t}\bm{A}\cdot\begin{bmatrix}\bm{Z}_{1}^{t}\\ \vdots\\ \bm{Z}_{L}^{t}\end{bmatrix}$. (33)
Thus, the variance of the aggregated correlated Gaussian privacy noise $\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{k}^{t}$ is given by
$\nu_{k}^{t}\triangleq\mathrm{var}\left(\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{m}^{t}\right)=\lVert\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}\bm{I}_{D}.$ (34)
Let $\bm{\Lambda}_{k}^{t,-k_{1}}$ denote $\bm{\Lambda}_{k}^{t}$ but with $\bm{\Lambda}_{k}^{t}(k_{1})=0$. Similarly, we get
$\nu_{k}^{t,-k_{1}}\triangleq\mathrm{var}\left(\sum_{m\neq k_{1}}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{m}^{t}\right)=\lVert\bm{\Lambda}_{k}^{t,-k_{1}}\bm{A}\rVert^{2}\bm{I}_{D}.$ (35)
Finally, the right side of(32) becomes
$\frac{D}{2}(1-p_{k,k_{1}})\log\left((2\pi e)g_{k,k_{1}}^{2}\zeta^{2}\right)\frac{D}{2}\log\left(\frac{\lVert\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}}{\lVert\bm{\Lambda}_{k}^{t,-k_{1}}\bm{A}\rVert^{2}}\right).$ (36)
[…]
V-3 LMIP at the Server
To better showcase the conclusion, we assume $\operatorname{Cov}[\Delta\bm{\Theta}_{k}^{t}])=\mathrm{Diag}(\bm{I}_{D}\zeta^{2})$.
Theorem V.3 (LMIP in Server Aggregation).
In secure aggregation, the server can reconstruct the exact global model $\sum_{m=1}^{K}w_{m}\Delta\bm{\Theta}_{m}^{t}$ without learning about each local update $\Delta\bm{\Theta}_{k}^{t}$. The global model satisfies $\mu_{3}$-LMIP, where
$\mu_{3}=\frac{D}{2}\log\left(1+\frac{w_{k}^{2}}{\sum_{m\neq k}w_{m}^{2}}\right).$ (37)
In the special case, where $w_{k}=\frac{1}{M}$, $\mu_{3}=\frac{D}{2}\log\left(1+\frac{1}{M-1}\right)$.
Proof.
The information leakage about $\Delta\bm{\Theta}_{k}^{t}$ in the exact global model is given by
$\mathrm{I}\left(\Delta\bm{\Theta}_{k}^{t};\sum_{m=1}^{K}w_{m}\Delta\bm{\Theta}_{m}^{t}\bigg{|}\{w_{m}\}_{m=1}^{K}\right)$
$=$ $\mathrm{h}\hskip-2.84526pt\left(\sum_{m=1}^{K}\hskip-2.84526ptw_{m}\Delta\bm{\Theta}_{m}^{t}\bigg{|}\{w_{m}\}_{m=1}^{K}\hskip-2.84526pt\right)\hskip-2.84526pt-\hskip-1.42262pt\mathrm{h}\hskip-2.84526pt\left(\hskip-2.84526pt\sum_{m\neq k}\hskip-2.84526ptw_{m}\Delta\bm{\Theta}_{m}^{t}\bigg{|}\{w_{m}\}_{m=1}^{K}\hskip-2.84526pt\hskip-1.42262pt\right)$
$=$ $\mathrm{h}\left(\mathcal{N}\left(\mathbf{0},\sum_{m=1}^{K}w_{m}^{2}\zeta^{2}\bm{I}_{D}\right)\right)-\mathrm{h}\left(\mathcal{N}\left(\mathbf{0},\sum_{m\neq k}w_{m}^{2}\zeta^{2}\bm{I}_{D}\right)\hskip-2.84526pt\right)$
$=$ $\frac{D}{2}\log\left(1+\frac{w_{k}^{2}}{\sum_{m\neq k}w_{m}^{2}}\right).$ (38)
Remark V.1.
Lemma V.3 provides a key insight into privacy behavior in secure aggregation. As the weight $w_{k}$ increases, the information leakage $\mu_{3}$ also increases. This indicates that, for model parameters of comparable scale, greater importance leads to weaker privacy guarantees the global model. Interestingly, when $w_{k}=\frac{1}{K}$, i.e., all clients are equally important, we observe that $\lim_{K\rightarrow\infty}\mu_{3}\rightarrow 0$. However, this vanishing leakage does not imply asymptotically perfect privacy in large-scale systems. On the contrary, it stems from the fact that $\lim_{K\rightarrow\infty}\mathrm{h}\left(\mathcal{N}\left(\mathbf{0},K\cdot\frac{1}{K^{2}}\zeta^{2}\bm{I}_{D}\right)\right)=\lim_{K\rightarrow\infty}\mathrm{h}\left(\mathcal{N}\left(\mathbf{0},(K-1)\cdot\frac{1}{K^{2}}\zeta^{2}\bm{I}_{D}\right)\right)=$-$\infty$, which implies no privacy.
VI Local Differential Privacy (LDP)
This section examines the model privacy by LDP across all levels in Fair-SecCoGC. We begin by presenting the relevant definitions, adapted to our context for clarity and readability.
Definition VI.1 ($(\epsilon,\delta)$-LDP).
The randomized mechanism $\mathcal{M}:\mathcal{O}\rightarrow\mathbb{R}^{D}$ is $(\epsilon,\delta)$ locally differentially private if for all dataset instances $\forall\bm{\Theta},\bm{\Theta}^{\prime}\in\mathcal{O}$, and any measurable set $\mathcal{S}\subseteq\mathrm{Range}(\mathcal{M})$, it holds that
$\mathrm{P_{r}}(\mathcal{M}(\bm{\Theta})\in\mathcal{S})\leq e^{\epsilon}\mathrm{P_{r}}(\mathcal{M}(\bm{\Theta}^{\prime})\in\mathcal{S})+\delta.$ (39)
Definition VI.2 (Guassian Mechanism[3]).
If a node wishes to release a function $f(\bm{\Theta})$ of an input $\bm{\Theta}$ subject to $(\epsilon,\delta)$-LDP. The Gaussian release mechanism is defined as
$\mathcal{M}(\bm{\Theta})\triangleq f(\bm{\Theta})+\mathcal{N}(0,\varepsilon^{2}\bm{I}).$ (40)
If the $\ell_{2}$-sensitivity of the function is bounded by $\Delta_{f}$, i.e., $\lVert f(\bm{\Theta})-f(\bm{\Theta}^{\prime})\rVert\leq\Delta_{f}$, then $\forall\bm{\Theta},\bm{\Theta}^{\prime}$, $\forall\delta\in(0,1]$, the Guassian mechanism satisfies $(\epsilon,\delta)$-LDP, where
$\epsilon=\frac{\Delta_{f}}{\varepsilon}\left[2\log\left(\frac{1.25}{\delta}\right)\right]^{\frac{1}{2}}.$ (41)
VI-1 Peer-to-Peer LDP
The local model updates are masked locally prior to transmission as in (6), i.e., $f(\Delta\bm{\Theta}_{k}^{t})=\Delta\bm{\Theta}_{k}^{t}$. Even if client $m$ can decode client $k$’s transmission, it should not confidently learn about $\Delta\bm{\Theta}_{k}^{t}$.
Theorem VI.1 (Peer-to-Peer LDP).
Given the secret key $\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})$, $\Delta\bm{\Theta}_{k}^{t},\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathbb{B}_{D}(R)$, and $\forall\delta_{m,k}^{(1)}\in(0,1]$, the masked transmission from client $k$ to client $m$ via unreliable link $\mathrm{Ber}(1-p_{m,k})$ is $(\epsilon_{m,k}^{(1)},(1-p_{m,k})\delta_{m,k}^{(1)})$-differentially private, where
$\epsilon_{m,k}^{(1)}=\frac{2R}{\lambda}\left[2\log\left(\frac{1.25}{\delta_{m,k}^{(1)}}\right)\right]^{\frac{1}{2}}.$ (42)
In other words, for any measurable set $\mathcal{S}$, it holds that
$\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t}\in\mathcal{O}_{k})\leq$ $e^{\epsilon_{m,k}^{(1)}}\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathcal{O}_{k})$
$\hskip 14.22636pt+(1-p_{m,k})\delta_{m,k}^{(1)}.$ (43)
Proof.
Proofs are provided in [32, Appendix A].  [^2] The supplementary file can be found at https://arxiv.org/abs/2507.07565.
VI-2 LDP at Relay
The client $k$ computes the partial sum $\bm{W}_{k}^{t}=\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t})$, as in (8), where $\mathcal{U}_{k}^{t}$ contains the received updates by client $k$ and its own update. Let $\bm{M}_{k}^{t}=\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{N}_{m}^{t}$ be the aggregated privacy noise. Following (34), we have $\nu_{k}^{t}=\mathrm{var}(\bm{M}_{k}^{t})$. When an eavesdropper (possibly the server) observes the masked partial sum $\bm{W}_{k}^{t}$, it should not be able to distinguish the identity of any participating client, i.e., it should not confidently infer whether client $j$’s contribution is included in $\bm{W}_{k}^{t}$. The privacy comes from two stochastic factors: (i) the random participation of the neighbors of client $k$, and (ii) the aggregated privacy noise, yet the privacy noises $\bm{N}_{m}^{t}$ are correlated across clients due to the coordinated nature of the secret keys. Let $\bm{W}_{k}^{t,-j}\triangleq\sum_{m\in\mathcal{U}_{k}^{t}\setminus\{j\}}g_{k,m}(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t})$ denote the partial sum excluding client $j$’s contribution. Similarly, define $\bm{M}_{k}^{t,-j}=\sum_{m\in\mathcal{U}_{k}^{t}\setminus\{j\}}g_{k,m}\bm{N}_{m}^{t}$, and its variance $\nu_{k}^{t,-j}=\mathrm{var}(\bm{M}_{k}^{t,-j})$. While the network topology renders both $\nu_{k}^{t}$ and $\nu_{k}^{t,-j}$ random variables, this challenge can be effectively addressed through an appropriate choice of the Bernstein parameter [7]. Moreover, the challenge of the presence of the unbounded r.v. $\bm{N}_{j}^{t}$ in $\ell_{2}$-sensitivity can be addressed using Chi-squared tail bound. The privacy of the presence of client $j$’s contributions in $k$-th partial sum is formalized below.
Theorem VI.2 (LDP for Each Individual Contribution).
Given secret keys $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ generated by Construction IV.2, then the variance of $\bm{M}_{k}^{t}$ is given by $\nu_{k}^{t,-j}\bm{I}_{D}=\lVert\bm{\Lambda}_{k}^{t,-j}\bm{A}\rVert^{2}\bm{I}_{D}$, and $\bar{\nu}_{k}^{-j}=\mathbb{E}[\nu_{k}^{t,-j}]=\left(\lambda^{2}\sum_{m=1}^{K}(1-p_{k,m})g_{m,k}^{2}+\bm{\Lambda}_{k}^{\mathrm{E},-j}\bm{A}\bm{A}^{\top}(\bm{\Lambda}_{k}^{\mathrm{E},-j})^{\top}\right)$, where $\bm{\Lambda}_{k}^{\mathrm{E},-j}=\mathbb{E}[\bm{\Lambda}_{k}^{t,-j}]$. For $\forall\delta_{0}$, $r_{1}>0$ and $\delta^{\prime}\in(0,1)$ such that $\mathrm{P_{r}}(\lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\geq r_{1})\leq\delta^{\prime}$, [^3] The choice of Bernstein parameter $r_{1}$ is discussed in [32, Appendix F]. the partial sum $\bm{W}_{k}^{t}$ is $(\epsilon_{k,j}^{(2)},(1-p_{k,j})(\delta^{\prime}+\delta_{j}^{(2)}))$-differentially private w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$ in protecting the identity of any participating node $j$, where
$\epsilon_{k,j}^{(2)}=\begin{cases}\left[2\log\left(\frac{1.25}{\delta_{j}^{(2)}}\right)\right]^{\frac{1}{2}}\frac{\lvert g_{k,j}\rvert\left(R+\lambda\sqrt{D(1+\delta_{0})}\right)}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}\;\;\;\mathrm{if}\;p_{k,j}\geq 0,\\ \hskip 71.13188pt0\;\;\;\;\;\;\;\;\;\;\;\;\;\;\hskip 42.67912pt\;\;\;\mathrm{if}\;p_{k,j}=0.\end{cases}$ (44)
In other words, for any measurable set $\mathcal{S}$, it holds that
$\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\in\mathcal{S})\leq e^{\epsilon_{k,j}^{(2)}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S})+(1-p_{k,j})(\delta^{\prime}+\delta_{j}^{(2)}),$ (45)
w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$.
Proof.
Proofs are provided in [32, Appendix B]. ∎
Apart from the participating identity, when the transmission from client $j$ is perturbed, i.e., when $\Delta\bm{\Theta}_{j}^{t}=\bm{u}$ is perturbed to $\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime}$, the aggregated partial sums should not fluctuate significantly. This privacy guarantee is formalized below.
Theorem VI.3 (LDP Against Perturbation of Local data).
Given secret keys $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ generated by Construction IV.2, then its variance is $\nu_{k}^{t}=\lVert\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}\bm{I}_{D}$. The $\bar{\nu}_{k}=\mathbb{E}[\nu_{k}^{t}]$ and $\bm{\Lambda}_{k}^{\mathrm{E}}=\mathbb{E}[\bm{\Lambda}_{k}^{t}]$ are specified in Theorem VI.2. For $r_{1}>0$ and $\delta^{\prime}\in(0,1)$ such that $\mathrm{P_{r}}(\lvert\nu_{k}-\bar{\nu}_{k}\rvert\geq r_{1})\leq\delta^{\prime}$, the partial sum $\bm{W}_{k}^{t}$ is $(\epsilon_{k,j}^{(3)},(1-p_{k,j})(\delta^{\prime}+\delta_{j}^{(3)})$-differentially private in protecting the local date of any participating node $j$ from perturbations, where
$\epsilon_{k,j}^{(3)}=\begin{cases}2\cdot\left[2\log\left(\frac{1.25}{\delta_{j}^{(3)}}\right)\right]^{\frac{1}{2}}\frac{\lvert g_{k,j}\rvert R}{\sqrt{\bar{\nu}_{k}-r_{1}}}\;\;\;\;\;\mathrm{if}\;p_{k,j}\geq 0,\\ \hskip 42.67912pt\;\;\;\;\;0\;\;\;\;\;\;\hskip 42.67912pt\;\;\;\mathrm{if}\;p_{k,j}=0.\end{cases}$ (46)
In other words, for any $\bm{u},\bm{u^{\prime}}\in\mathbb{B}_{D}(R)$, and any measurable set $\mathcal{S}$, it holds that
$\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})$ $\leq e^{\epsilon_{k,j}^{(3)}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime})$
$+(1-p_{k,j})(\delta^{\prime}+\delta_{j}^{(3)}).$ (47)
Proof.
Proofs are provided in [32, Appendix C]. ∎
VI-3 LDP at the Server
Suppose the server receives multiple partial sums from clients, denoted by $\mathcal{V}^{t}$, but fails to reconstruct the global model using the combinator $\bm{c}_{f^{\prime}}$ ($\bm{c}_{f^{\prime}}$-Failure). In that case, the resulting aggregation, $\bm{F}_{f^{\prime}}^{t}=\sum_{k\in\mathcal{V}^{t}}c_{f^{\prime},k}\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}(\Delta\bm{\Theta}_{m}^{t,I}+\bm{N}_{m}^{t})$, is noisy, and the effective noise in the aggregation is $\bm{N}_{F_{f^{\prime}}}^{t}=\sum_{k\in\mathcal{V}^{t}}c_{f^{\prime},k}\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{N}_{m}^{t}$. From $\bm{F}_{f^{\prime}}^{t}$, the server should not infer the participation of client $j$ at the relaying client $k$. The variance of $\bm{N}_{F_{f^{\prime}}}^{t}$ is $\nu_{F_{f^{\prime}}}^{t}\bm{I}_{D}=\mathrm{var}(\bm{N}_{F_{f^{\prime}}}^{t})\bm{I}_{D}$ (specified in Appendix I), and its expectation is $\bar{\nu}_{F_{f^{\prime}}}^{t}=\mathbb{E}[\mathrm{var}(\bm{N}_{F_{f^{\prime}}}^{t})]$. The stochasticity providing privacy for client $j$’s contribution are: (i) the local noise $\bm{N}_{j}^{t}$ added by client $j$; (ii) the intermittent nature of the links $\tau_{k,j}$ from client $j$ to its neighbors $k$; and (iii) the intermittent nature of the link $\tau_{k}$ from client $k$ to the server. Let $\bm{F}_{f^{\prime}}^{t,\{-j,k\}}$ be the aggregated results excluding client $j$’s contribution at relay $k$. Similarly, define $\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}$ and $\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}$. The privacy is formalized below.  [^4] Note that, due to the correlation among the secret keys, the standard composition theorem of DP [3] cannot apply.
Theorem VI.4 (LDP in $\bm{c}_{f^{\prime}}$-Failure).
Given secret keys $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ generated by Construction IV.2 and the combinator $\bm{c}_{f^{\prime}}$, then for any $r_{2}>0$ and $\delta^{\prime}\in(0,1)$ such that $\mathrm{P_{r}}(\lvert\nu_{f^{\prime}}^{t,\{k,-j\}}-\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}\rvert\geq r_{2})\leq\delta^{\prime}$, [^5] The choice of Bernstein parameter $r_{2}$ is discussed in [32, Appendix F]. then, in $\bm{c}_{f^{\prime}}$-Failure, $\bm{F}_{f^{\prime}}^{t}$ is $(\epsilon_{f^{\prime},\{k,j\}}^{(4)},(1-p_{k})(1-p_{k,j})(\delta_{f^{\prime},\{k,j\}}^{(4)}+\delta^{\prime})$ differentially private w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$ in protecting the identity of any participating client $j$ at relay $k$, where
$\epsilon_{f^{\prime},\{k,j\}}^{(4)}\hskip-2.84526pt=\hskip-2.84526pt\left[\hskip-1.42262pt2\log\left(\hskip-2.84526pt\frac{1.25}{\delta_{f^{\prime},\{k,j\}}^{(4)}}\hskip-2.84526pt\right)\hskip-4.2679pt\right]^{\frac{1}{2}}\hskip-5.69054pt\frac{\lvert c_{f^{\prime},k}g_{k,j}\rvert(R+\lambda\sqrt{D(1+\delta_{0})})}{\sqrt{\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}-r_{2}}}.$ (48)
In other words, for any measurable set $\mathcal{S}$, it holds that
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\in\mathcal{S})\leq$ $e^{\epsilon_{f^{\prime},\{k,j\}}^{(4)}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S})$
$\hskip-14.22636pt+(1-p_{k})(1-p_{k,j})(\delta_{f^{\prime},\{k,j\}}^{(4)}+\delta^{\prime}).$ (49)
Moreover, for any $\bm{u},\bm{u^{\prime}}\in\mathbb{B}_{D}(R)$, it holds that
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})$ $\leq e^{\epsilon_{k,j,f^{\prime}}^{(5)}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime})$
$+(1-\tilde{p}_{j,f^{\prime}})(\delta^{\prime}+\delta_{j}^{(5)}),$ (50)
w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$, where
$\epsilon_{k,j,f^{\prime}}^{(5)}=2\left[2\log\left(\frac{1.25}{\delta_{j}^{(5)}}\right)\right]^{\frac{1}{2}}\frac{\sum_{k=1}^{K}\lvert c_{f^{\prime},k}g_{k,j}\rvert R}{\sqrt{\bar{\nu}^{t,f^{\prime}}-r_{2}}},$ (51a)
$\tilde{p}_{j,f^{\prime}}=\hskip-2.84526pt\sum_{\mathcal{V}_{j,f^{\prime}}^{t}\in\{\mathcal{V}_{j}\cap\mathcal{V}_{f^{\prime}}\}}\prod_{k_{1}\in\mathcal{V}_{j}^{t}}(1-p_{k_{1}})\cdot p_{k_{1},j}\hskip-2.84526pt\hskip-8.53581pt\prod_{k_{2}\in\{\mathcal{V}_{j}\cap\mathcal{V}_{f^{\prime}}\}\setminus\mathcal{V}_{j,f^{\prime}}^{t}}\hskip-8.53581ptp_{k_{2}}.$ (51b)
The $\tilde{p}_{j,f^{\prime}}$ denotes the probability of client $j$’s contribution not being conveyed by those participating in $\bm{c}_{f^{\prime}}$-combination, and $\mathcal{V}_{f^{\prime}}\triangleq\{k:c_{f^{\prime},k}\neq 0\}$.
Proof.
Proofs are provided in [32, Appendix D]. ∎
With a slight abuse of the paper organization, we examine privacy via obfuscation by other local models in successful server aggregation. Unlike standard LDP, where noise is explicitly added to each input, privacy arises naturally from the randomness or uncertainty in the other local models. Nevertheless, we can leverage results from LDP theory to derive the following results. Specifically, the server should not see significant changes in the global model $\bm{M}^{t}=\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t}$ resulting from the presence or perturbation of $\Delta\bm{\Theta}_{j}^{t}$. To facilitate theoretical analysis, we assume $\operatorname{Cov}[\Delta\bm{\Theta}_{k}^{t}])=\mathrm{Diag}(\bm{I}_{D}\zeta^{2})$.
Theorem VI.5 (Privacy in Global Model).
Given $\{\Delta\bm{\Theta}_{k}^{t}\}_{k=1}^{K}$ and $\forall k:\Delta\bm{\Theta}_{k}^{t}\sim\mathcal{N}(\bm{0},\zeta^{2}\bm{I}_{D})$, for any $\delta_{0}>0$ such that $\mathrm{P_{r}}\left(\left\lVert\Delta\bm{\Theta}_{j}^{t}\right\rVert\leq\zeta\sqrt{D(1+\delta_{0})}\right)\geq 1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$, the global model $\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t}$ is $(\epsilon^{(6)},\delta^{(6)})$-differentially private w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$ in protecting the identity of any node $j$, where
$\epsilon^{(6)}=\sqrt{\frac{D(1+\delta_{0})}{K-1}}\left[2\log\left(\frac{1.25}{\delta^{(6)}}\right)\right]^{\frac{1}{2}}.$ (52)
In other words, for any measurable set $\mathcal{S}$, it holds that
$\mathrm{P_{r}}(\bm{M}^{t,-j}\in\mathcal{S})\leq e^{\epsilon^{(6)}}\mathrm{P_{r}}(\bm{M}^{t}\in\mathcal{S})+\delta^{(6)},$ (53)
w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$. Moreover, for any $\bm{u},\bm{u^{\prime}}\in\mathbb{B}_{D}(r_{3})$, it holds that
$\mathrm{P_{r}}(\bm{M}^{t,-j}$ $\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})$
$\leq e^{\epsilon^{(7)}}\mathrm{P_{r}}(\bm{M}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})+\delta^{(7)}.$ (54)
w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$, where
$\epsilon^{(7)}=2\sqrt{\frac{D(1+\delta_{0})}{K-1}}\left[2\log\left(\frac{1.25}{\delta^{(7)}}\right)\right]^{\frac{1}{2}}.$ (55)
Proof.
Proofs are provided in [32, Appendix E]. ∎
Remark VI.1.
When $K\rightarrow\infty$, we observe $\epsilon^{(6)}\rightarrow 0$ and $\epsilon^{(7)}\rightarrow 0$, which indicates that secure aggregation in IID settings is not differentially private for large systems when clients have equal importance.
VII Reliability
This section evaluates the server’s ability to recover the exact global model. To achieve exact recovery, the server must receive at least KsK-s complete partial sums. Otherwise, the server aggregation is not meaningful. This leads us to the following three disjoint failure scenarios.
(i) In the gradient-sharing phase, at least $s$ clients do not receive all $s$ masked gradients from its neighbors. Then no matter how good the client-to-server channels are, the server cannot perform a meaningful aggregation. The probability of this event is given by $P_{1}$ in [31, (10)].
(ii) In the gradient-sharing phase, all transmissions are successful, but there are at least $s$ client-to-server transmission failures. As a result, although all partial sums computed at clients are complete, fewer than KsK-s of them eventually arrive at the server. The probability of this event is given by $P_{2}$ in [31, (11)].
(iii) In the gradient-sharing phase, if $v$ (fewer than $s$) clients do not receive the masked gradients from all of its $s$ neighbors, and if more than svs-v transmissions from clients holding complete partial sums to the server fail, it is impossible to recover the global model. The probability of this event is given by $P_{3}$ in [31, (14)].
For all remaining scenarios beyond the three cases above, the global model can be reconstructed accurately. In other words, SecCoGC yields a binary recovery outcome for the global model, following $\mathrm{Ber}(1-P_{O})$, where
$P_{O}=P_{1}+P_{2}+P_{3}.$ (56)
One may wonder what $P_{O}$ can influence, especially given that the global model is perfect whenever it can be recovered. If the communication repeats until the server can successfully recover the global model, $P_{O}$ decides the communication overhead per round, detailed in [31, Remark 3]. In this case, the convergence is guaranteed [33]. If the local training is performed consecutively based on its own latest trained local model when the server fails to recover the global model, $P_{O}$ will statistically affect the length of local accumulation. In this case, the convergence rate depends on $P_{O}$, as will be discussed in the next section. In both designs, the cost-efficient choice of $s$ in SecCoGC is similar to [31, Section V]
VIII Convergence
This section provides a non-convex convergence rate analysis of SecCoGC under an arbitrary choice of local solvers. The analysis also applies to the general CoGC and GC. To begin with, we present several assumptions concerning LOFs that are used throughout the analysis.
Assumption VIII.1 (Smoothness).
For $\forall i\in[K]$, each local objective function is lower bounded by $\mathcal{L}_{k}(x)\geq\mathcal{L}^{\star}$ and is Lipschitz differentiable and its gradient $\nabla\mathcal{L}_{k}(x)$ is $G$-smooth, i.e., $\lVert\nabla\mathcal{L}_{k}(x)-\nabla\mathcal{L}_{k}(y)\rVert\leq G\lVert x-y\rVert$, $\forall x,y\in\mathbb{R}^{D}$.
Assumption VIII.2 (Unbiasedness and Bounded Variance).
For $\forall i\in[K]$, the local stochastic gradient is an unbiased estimator of the true local gradient, i.e., $\mathbb{E}_{\xi}[\nabla\mathcal{L}_{k}(x|\xi)]=\nabla\mathcal{L}_{k}(x)$, where $\nabla\mathcal{L}_{k}(x)$ has bounded data variance $\mathbb{E}_{\xi}[\lVert\nabla\mathcal{L}_{k}(x|\xi)-\nabla\mathcal{L}_{k}(x)\rVert^{2}]\leq\sigma^{2}$, $\sigma^{2}>0$.
Assumption VIII.3 (Bounded Dissimilarity).
For any set of $\forall\{w_{k}\}_{K=1}^{M}:\sum_{K=1}^{M}w_{k}=1$, the dissimilarity between the local objective functions $\nabla\mathcal{L}_{k}(x)$ and the global objective function $\nabla\mathcal{L}(x)$ is bounded by $\sum_{k=1}^{K}w_{k}\lVert\nabla\mathcal{L}_{k}(x)\rVert^{2}\leq\beta^{2}\lVert\nabla\mathcal{L}(x)\rVert^{2}+\kappa^{2}$, $\beta^{2}\geq 1,\kappa^{2}\geq 0$.
This combination of assumptions VIII.1VIII.3 is the most relaxed than those commonly adopted in the literature [15] and allows for broad applicability in real-world settings.
Lemma VIII.1 (Accumulation of Local Updates).
With arbitrary local solvers, the accumulated local update on each client after $R_{t_{s}}$ training rounds, each containing $I$ consecutive iterations, is given by
[…]
$\sum_{i_{r}=1}^{R_{t_{s}}I}a_{i_{r}}\mathbb{E}\left[\left\lVert\Delta\bm{\Theta}_{k}^{t,R_{t_{s}}I}-\bm{\Theta}^{t-1}\right\rVert^{2}\right]$
$\leq\frac{2\eta^{2}R_{t_{s}}^{3}\lVert\bm{a}\rVert_{1}^{3}}{1-2\eta^{2}G^{2}R_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}}\mathbb{E}\left[\lVert\nabla\mathcal{L}_{k}\left(\bm{\Theta}^{t-1}\right)\rVert^{2}\right],$ (58)
where $a_{i_{r}}=a_{i_{r}\bmod I},i_{r}=1,\cdots,R_{t_{s}}I$, and $a_{i_{r}\bmod I}$ corresponds to entries in $\bm{a}$, defined in (4).
Sketch Proof.
By replacing II with $R_{t_{s}}I$ and substituting (57) into [15, Lemma 3], we derive Lemma VIII.1. ∎
Theorem VIII.1.
Given an arbitrary local solver $\bm{a}$, for any algorithms that result in the binary recovery of the global model with failure probability $P_{O}$, if the learning rate is set as $\eta=\frac{1}{G}\sqrt{\frac{M}{T}}$ and the threshold of total training rounds $T$ is large enough, under Assumptions VIII.1VIII.3, the following convergence bound yields with probability $99.74\%$,
$\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}\mathbb{E}\left[\lVert\Delta\mathcal{L}(\bm{\Theta}^{t-1})\rVert^{2}\right]\leq\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}\mathbb{E}\left[\lVert\Delta\mathcal{L}(\bm{\Theta}^{t-1})\rVert^{2}\right]$
$\leq\frac{2G(\mathcal{L}(\bm{\Theta}^{T})-\mathcal{L}^{\star})}{\lVert\bm{a}\rVert_{1}(1-P_{O})\sqrt{KT}}+\frac{2\lVert\bm{a}\rVert_{1}(1+P_{O})}{\left(1-P_{O}\right)^{2}}\sqrt{\frac{K}{T}}(\sigma^{2}+2\kappa^{2}).$ (59)
where $T_{s}=\min\{n:\sum_{t=1}^{n}R_{t_{s}}\geq T\}$ denotes the total number of successful global update.
Proof.
Proofs are provided in [32, Appendix G]. ∎
Remark VIII.1 (Role of $P_{O}$).
A large value of $P_{O}$ not only reduces the participation frequency, thereby slowing the convergence rate by $1-P_{O}$, but also introduces a slowdown by $(1-P_{O})^{2}$, due to data heterogeneity and local data variance.
Remark VIII.2 (Role of Privacy).
The convergence analysis demonstrates that the convergence behavior of SecCoGC is independent of its privacy performance, offering a theoretical optimality guarantee under arbitrarily strong privacy levels.
IX Simulations
IX-A Experimental Setups
We evaluate the following methods on MNIST and CINIC-10, across varying levels of data heterogeneity, privacy preservation, and network conditions.
  1. (i)
    Fair-SecCoGC, as described in Section III and IV.
  2. (ii)
    Standard FL with perfect connectivity. This benchmark provides insights into the ideal performance for any FL system under the same data heterogeneity.
  3. (iii)
    Non-private standard FL under unreliable communication from clients to the server.
  4. (iv)
    Private standard FL under unreliable communication, where the Gaussian mechanism is employed.
  5. (v)
    Private diversity network codes (DNC)-based FL[24], where the Gaussian mechanism is employed.
The update rule in partial participation in methods (iii), (iv), (v) is non-blind and follows [15, (17)]. That is, the effective step size is not reduced by missing updates.
The MNIST [34] is an easy-level benchmark dataset with clean and standardized format. The CINIC-10 [35], formed by both CIFAR-10 and downsampled ImageNet, is more difficult to train due to greater diversity and real-world noise. To simulate varying levels of data heterogeneity, local datasets are assigned using the Dirichlet distribution. The concentration parameter $\Gamma$ controls the level of heterogeneity among clients, where smaller values of $\Gamma$ indicate a higher imbalance.
(a) Symmetric network, $\Gamma=0.1$
(b) Asymmetric network, $\Gamma=0.1$
(c) Symmetric network, $\Gamma=0.2$
(d) Asymmetric network, $\Gamma=0.2$
Figure 3: Performance comparison under varying privacy levels, network conditions, and data heterogeneity on the MNIST dataset.
(a) Symmetric network, $\Gamma=0.1$
(b) Asymmetric network, $\Gamma=0.1$
(c) Symmetric network, $\Gamma=0.2$
(d) Asymmetric network, $\Gamma=0.2$
Figure 4: Performance comparison under varying privacy levels, network conditions, and data heterogeneity on the CINIC dataset.
In the simulation, the number of clients is set to $K=10$. Each client is assigned an equal number of data samples. The total number of training rounds is set to $T=100$, and each client performs $I=5$ local training iterations per round. The coding parameter $s$ is set to $7$ without further specification. The inter-client connectivity is configured with $p_{m,k}=0.9,\forall m\neq k$. In symmetric network settings, the local connectivity is set uniformly as $p_{k}=0.7$ for all clients. In asymmetric networks, each client is assigned a connectivity value $p_{k}\in\{0.500,0.533,\cdots,0.8\}$, for $k=\{1,\cdots,K\}$. For MNIST, the privacy level $\lambda$ is set to $0.05$ and $0.1$; for CINIC-10, it is set to $0.03$ and $0.05$. Additional hyperparameters and neural network (NN) architectures are detailed in Table I.
TABLE I: NN architecture, hyperparameter specifications
Datasets MNIST CINIC-10
Neural network CNN CNN
Model architecture
C(1,10) – C(10,20) – DL(50) – L(10)
C(3,32) – RMC(32,32) – RMDL(512) – RDL(256) – RDL(10)
Loss function Negative log-likelihood loss (NLLL) Cross-entropy loss
Learning rate $\eta$ $\eta=0.002$ $\eta=0.02$
Batch size 1024
C(# in-channel, # out-channel): a 2D convolution layer (kernel size 3, stride 1, padding 1); R: ReLU activation function; M: a 2D max-pool layer (kernel size 2, stride 2); L: (# outputs): a fully-connected linear layer; D: a dropout layer (probability 0.2).
IX-B Results and Discussions
We report the average results of multiple runs. For a fair comparison, all results are truncated at the $100^{\mathrm{th}}$ training round. The simulation results are presented in Fig. 3 and Fig. 4. These results demonstrate the strong robustness of SecCoGC against frequent communication failures between clients and the server, while preserving arbitrarily strong privacy guarantees. Detailed analyses are provided below.
Fast Convergence Under Privacy Preservation
When $\lambda$ is set to $0.05$ for MNIST and $0.03$ for CINIC-10, i.e., moderate levels of privacy noise, the private benchmark methods (including both private standard FL and private DNC) can work, but with significantly degraded performance. In addition to slower and less stable convergence, the final accuracy is also substantially reduced. This is because noise increases the amplitude of oscillations during optimization, and the larger step size prevents the final convergence point from accurately approaching the true global optimum. For this reason, even though DNC-based FL nearly ensures full recovery of all masked local models, masking still causes a considerable performance gap between ideal FL with perfect connectivity and DNC-based FL. As for the private standard FL, the perturbation induced by privacy-preserving noise, compounded by unreliable communication, disrupts the optimization trajectory of the masked global model. This results in a more pronounced divergence from the true global optimum, thereby leading to inferior performance compared to private DNC. With an increased privacy level ($\lambda=0.1$ for MNIST and $\lambda=0.05$ for CINIC-10), the utility degradation introduced by privacy noise becomes unacceptable. In this high-privacy regime, private standard FL consistently diverges across all configurations. Despite private DNC’s near-complete removal of communication unreliability, the global model’s accuracy remains significantly degraded, highlighting the intrinsic trade-off between strong privacy guarantees and model utility. In contrast, our proposed SecCoGC consistently approaches ideal performance across all privacy levels. This is because SecCoGC reconstructs the global model exactly whenever recovery is possible. Nonetheless, we also observe slight fluctuations of SecCoGC compared to ideal FL, which is due to the increased variance resulting from the accumulation of local model updates, implied by Lemma VIII.1.
Straggler Mitigation Against Unreliable Communication
On both MNIST and CINIC-10, the performance drop solely due to unreliable communication becomes more pronounced with increasing data imbalance and network asymmetry, as evidenced by the widening gap between ideal FL and unprivate standard FL. In contrast, our proposed SecCoGC naturally ensures the optimality over heterogeneous networks with varying data heterogeneity, owing to its binary decoding mechanism. It is worth noting that SecCoGC effectively mitigates the impact of unreliable communication and achieves exact recovery of the global model through its structured coding scheme. This coding structure enforces the correct coefficient of each component in the aggregation, thereby ensuring that recovery is independent of any prior information. Under the same conditions, state-of-the-art methods [16, 19, 15, 17] cannot operate effectively in our setting. Furthermore, SecCoGC leverages this coefficient-enforcing property to enable secure aggregation, wherein user privacy is preserved and ambiguity is ultimately removed from the final aggregation, providing additional benefits that existing straggler-resilient schemes do not offer. Further details regarding its straggler resilience follow similarly to the analysis in [31].
X Conclusion
In this work, we introduce SecCoGC, a real-field secure aggregation scheme designed to enhance privacy and reliability under communication uncertainties. Building upon this, we propose Fair-SecCoGC, an extension that ensures fairness in user privacy protection while achieving optimal privacy preservation given a total power constraint of the privacy noise. We formalize the underlying problem, present key construction methods, and analyze their integration within FL frameworks. By ensuring the coordination structure of secret keys and the exact reconstruction of the global model, the proposed schemes support fast convergence and optimal performance under frequent communication disruptions, without requiring prior knowledge of the network topology. Through rigorous theoretical analysis and extensive empirical evaluation, we demonstrate that both SecCoGC and Fair-SecCoGC offer strong robustness to unreliable communication while providing arbitrarily strong privacy guarantees.
Appendices
Appendix A Proof of Theorem VI.1
The model of peer-to-peer LDP from client $k$ to client $m$ is similar to the standard Gaussian mechanism with noise $\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})$, but with imperfect connectivity $\tau_{m,k}$. The client $m$ should not confidently learn whether the local model update $\Delta\bm{\Theta}_{k}^{t}$ or $\Delta\bm{\Theta}_{k}^{t^{\prime}}$ is contained in the received masked message $\bm{Y}_{k}^{t}$. When client $m$ fails to decode message from client $k$, i.e., $\tau_{m,k}=0$, we have $\bm{Y}_{k}^{t}=0$ regardless of the values of $\Delta\bm{\Theta}_{k}^{t},\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathbb{B}_{D}(R)$, it yields perfect privacy, i.e.,
$\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t}\in\mathcal{O}_{k},\tau_{m,k}=0)$
$\hskip 56.9055pt=\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathcal{O}_{k},\tau_{m,k}=0),$ (60)
If $\tau_{m,k}=1$, client $m$ successfully decode $\bm{Y}_{k}^{t}$, the standard Gaussian mechanism applies, with the $\ell_{2}$-sensitivity given by
$\sup_{\Delta\bm{\Theta}_{k}^{t},\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathbb{B}_{D}(R)}\lVert\Delta\bm{\Theta}_{k}^{t}-\Delta\bm{\Theta}_{k}^{t^{\prime}}\rVert\leq 2R.$
Thus, it holds that
$\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t}\in\mathcal{O}_{k},\tau_{m,k}=1)$
$\leq e^{\epsilon_{m,k}^{(1)}}\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathcal{O}_{k},\tau_{m,k}=1)+\delta_{k}^{(1)},$ (61)
where
$\epsilon_{m,k}^{(1)}=\frac{2R}{\lambda}\left[2\log\left(\frac{1.25}{\delta_{m,k}^{(1)}}\right)\right]^{\frac{1}{2}}.$ (62)
Following similar steps in [7, (73)], we have
$\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t}\in\mathcal{O}_{k})$
$=p_{m,k}\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t}\in\mathcal{O}_{k},\tau_{m,k}=0)$
$\hskip 14.22636pt+(1-p_{m,k})\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t}\in\mathcal{O}_{k},\tau_{m,k}=1)$ (63a)
$\leq p_{m,k}e^{\epsilon_{m,k}^{(1)}}\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t}\in\mathcal{O}_{k},\tau_{m,k}=0)$
$\hskip 14.22636pt+(1-p_{m,k})e^{\epsilon_{m,k}^{(1)}}\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathcal{O}_{k},\tau_{m,k}=1)$
$\hskip 14.22636pt+(1-p_{m,k})\delta_{k}^{(1)}$ (63b)
$=e^{\epsilon_{m,k}^{(1)}}\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathcal{O}_{k})+(1-p_{m,k})\delta_{k}^{(1)},$ (63c)
where (63b) is due to (60) and that $e^{\epsilon_{m,k}^{(1)}}\geq 1$.
Appendix B Proof of Theorem VI.2
Now let us proceed to LDP at the relaying client $k$. The relaying client $k$ successfully collects the masked local model updates from clients in $\mathcal{U}_{k}^{t}$ (including itself), computes their partial sum $\bm{W}_{k}^{t}$, and transmits $\bm{W}_{k}^{t}$ to the server. During the transmission, an eavesdropper should not be able to distinguish the participation of client $j$ by observing $\bm{W}_{k}^{t}$.
If client $j\notin\mathcal{U}_{k}^{t}$, then $\bm{W}_{k}^{t,-j}=\bm{W}_{k}^{t}$. So for any measurable set $\mathcal{S}$, it holds that
$\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-2.84526pt\in\hskip-1.42262pt\mathcal{S})\hskip-1.42262pt=$ $\mathrm{P_{r}}(j\in\mathcal{U}_{k}^{t})\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t})$
$+\mathrm{P_{r}}(j\notin\mathcal{U}_{k}^{t})\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\hskip-2.84526pt\mathcal{S}|j\notin\mathcal{U}_{k}^{t}).$ (64)
Consider the network realizations reflected by $\mathcal{U}_{k}^{t}$, such that corresponding variance $\nu_{k}^{t,-j}$ lies within the event $\mathcal{E}\triangleq\{\lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r_{1}\}$, where $r$ is chosen such that $\mathrm{P_{r}}(\mathcal{E}^{\mathrm{c}})<\delta^{\prime}$, $\delta^{\prime}\in(0,1)$, then
$\hskip-1.42262pt\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t})=$ $\mathrm{P_{r}}(\mathcal{E})\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})$
$\hskip-34.1433pt+\mathrm{P_{r}}(\mathcal{E}^{\mathrm{c}})\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E}^{\mathrm{c}})$
$\hskip-42.67912pt\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E}).$ (65)
Moreover, by counting all network realizations $\mathcal{A}$ lying within the event $\mathcal{E}$, we have
$\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})=\hskip-19.91692pt\sum_{\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r_{1}\end{subarray}}\hskip-19.91692pt\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\mathrm{P_{r}}(\mathcal{A}).$ (66)
For each specific realization $\mathcal{A}$, the form of its privacy loss is given by
$\log\left(\frac{\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}=\bm{z}|i\in\mathcal{A})}{\mathrm{P_{r}}(\bm{W}_{k}^{t}=\bm{z}|i\in\mathcal{A})}\right)$
$=$ $\log\left(\frac{\exp{\left(-\frac{\lVert\bm{z}-\sum_{m\in\mathcal{A}\setminus\{j\}}(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t})\rVert^{2}}{2\nu_{k}^{t,-j}}\right)}}{\exp{\left(-\frac{\lVert\bm{z}-\sum_{m\in\mathcal{A}}(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t})\rVert^{2}}{2\nu_{k}^{t,-j}}\right)}}\right)$
$=$ $-\frac{1}{2\nu_{k}^{t,-j}}\Bigg{(}\left\lVert\bm{z}-\sum_{m\in\mathcal{A}\setminus\{j\}}(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t})\right\rVert^{2}$
$\hskip 85.35826pt-\left\lVert\bm{z}-\sum_{m\in\mathcal{A}}(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t})\right\rVert^{2}\Bigg{)}.$ (67)
The form of (67) is identical to the privacy loss of the Gaussian mechanism with Gaussian noise $\nu_{k}^{t,-j}$. The $\ell_{2}$-sensitivity is bounded by
$\sup_{\hskip-8.53581pt\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r_{1}\end{subarray}}\hskip-2.84526pt\left\lVert\sum_{m\in\mathcal{A}}\hskip-2.84526ptg_{k,m}(\Delta\bm{\Theta}_{m}^{t}\hskip-2.84526pt+\hskip-2.84526pt\bm{N}_{m}^{t})\hskip-1.42262pt-\hskip-14.22636pt\sum_{m\in\mathcal{A}\setminus\{j\}}\hskip-11.38109ptg_{k,m}(\Delta\bm{\Theta}_{m}^{t}\hskip-2.84526pt+\hskip-2.84526pt\bm{N}_{m}^{t})\right\rVert$
$=\left\lVert g_{k,j}(\Delta\bm{\Theta}_{j}^{t}+\bm{N}_{j}^{t})\right\rVert\leq\left\lVert g_{k,j}\Delta\bm{\Theta}_{j}^{t}\right\rVert+\left\lVert g_{k,j}\bm{N}_{j}^{t}\right\rVert$
$\leq\lvert g_{k,j}\rvert R+\left\lVert g_{k,j}\bm{N}_{j}^{t}\right\rVert.$ (68)
Since that $g_{k,j}\bm{N}_{j}^{t}\sim\mathcal{N}(\bm{0},\lambda^{2}\bm{I}_{D})$, the second term in (68) can be bounded by Chi-squared tail bound, i.e., $\forall\delta_{0}$,
$\mathrm{P_{r}}\left(\left\lVert g_{k,j}\bm{N}_{j}^{t}\right\rVert\hskip-2.84526pt\leq\hskip-2.84526pt\lvert g_{k,j}\rvert\lambda\sqrt{D(1+\delta_{0})}\right)\hskip-1.42262pt\geq\hskip-1.42262pt1\hskip-1.42262pt-\hskip-1.42262pte^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}.$ (69)
Substituting (96) into (68), we can bound the $\ell_{2}$-sensitivity by
$\lvert g_{k,j}\rvert R+\left\lVert g_{k,j}\bm{N}_{j}^{t}\right\rVert\leq\lvert g_{k,j}\rvert\left(R+\lambda\sqrt{D(1+\delta_{0})}\right),$ (70)
with probability (w.p.) at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$.
Consequently, each specific realization $\mathcal{A}$ is $(\epsilon_{k,j}^{\mathcal{A},2},\delta_{j}^{(2)})$-differentially private w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$, i.e.,
$\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\leq e^{\epsilon_{k,j}^{\mathcal{A},2}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)},$ (71)
where $\epsilon_{k,j}^{\mathcal{A},2}=\hskip-2.84526pt\left[2\log\left(\frac{1.25}{\delta_{j}^{(2)}}\right)\right]^{\frac{1}{2}}\hskip-2.84526pt\frac{\lvert g_{k,j}\rvert\left(R+\lambda\sqrt{D(1+\delta_{0})}\right)}{\sqrt{\nu_{k}^{t,-j}}}\hskip-2.84526pt\triangleq\frac{c_{1}}{\sqrt{\nu_{k}^{t,-j}}}$.
Substitute (71) into (66), we get
$\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})$ (72a)
$\leq\hskip-14.22636pt\sum_{\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r\end{subarray}}\hskip-14.22636pt\mathrm{P_{r}}(\mathcal{A})\left(e^{\frac{c_{1}}{\sqrt{\nu_{k}^{t,-j}}}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)}\right)$ (72b)
$\leq e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\cdot\hskip-14.22636pt\sum_{\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r\end{subarray}}\hskip-14.22636pt\mathrm{P_{r}}(\mathcal{A})\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)}$ (72c)
$\leq e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\cdot\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)},$ (72d)
where (72c) holds since $\nu_{k}^{t,-j}>\bar{\nu}_{k}^{-j}-r_{1}$, (72d) is similar to (66).
Substitute (72d) into (65), we obtain
$\hskip-1.42262pt\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t})$
$\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\left(e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\cdot\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)}\right)$
$\leq\delta^{\prime}+e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\cdot\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)}.$ (73)
Furthermore, it holds that
$\hskip-1.42262pt\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\notin\mathcal{U}_{k}^{t})\leq e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\notin\mathcal{U}_{k}^{t}),$ (74)
since that $e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\geq 1$.
Substitute both (73) and (74) into (64), we get
$\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-2.84526pt\in\hskip-1.42262pt\mathcal{S})\hskip-1.42262pt\leq\hskip-1.42262pte^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-2.84526pt\in\hskip-1.42262pt\mathcal{S})\hskip-1.42262pt+\hskip-1.42262pt(1\hskip-1.42262pt-\hskip-1.42262ptp_{k,j})(\delta^{\prime}\hskip-2.84526pt+\hskip-1.42262pt\delta_{j}^{(2)}).$ (75)
The proof is completed.
Appendix C Proof of Theorem VI.3
For $\forall\bm{u},\bm{u}^{\prime}\in\mathbb{B}_{D}(R)$, similar to (64), it holds that
$\mathrm{P_{r}}(\bm{W}_{k}^{t}$ $\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})$
$=$ $\mathrm{P_{r}}(j\in\mathcal{U}_{k}^{t})\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},j\in\mathcal{U}_{k}^{t})$
$+\mathrm{P_{r}}(j\notin\mathcal{U}_{k}^{t})\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}.j\notin\mathcal{U}_{k}^{t}).$ (76)
Moreover, similar to (65), it holds that
$\mathrm{P_{r}}($ $\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},j\in\mathcal{U}_{k}^{t})$
$\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},j\in\mathcal{U}_{k}^{t},\mathcal{E}),$ (77)
and similar to (71), it holds that
$\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},i\in\mathcal{A})$
$\leq e^{\epsilon_{k,j}^{\mathcal{A},3}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime},i\in\mathcal{A})+\hskip-1.42262pt\delta_{j}^{(3)}.$ (78)
The $\ell_{2}$-sensitivity is given by
$\hskip 14.22636pt\sup_{\hskip-14.22636pt\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r_{1}\end{subarray}}\hskip 0.0pt\left\lVert\bm{W}_{k}^{t,-j}+\bm{u}+\bm{N}_{j}^{t}-\bm{W}_{k}^{t,-j}-\bm{u}^{\prime}-\bm{N}_{j}^{t}\right\rVert$
$\hskip 28.45274pt\leq 2\lvert g_{k,j}\rvert R,$ (79)
which gives $\epsilon_{k,j}^{\mathcal{A},3}=\left[2\log\left(\frac{1.25}{\delta_{j}^{(3)}}\right)\right]^{\frac{1}{2}}\frac{2\lvert g_{k,j}\rvert R}{\sqrt{\nu_{k}^{t}}}\triangleq\frac{c_{2}}{\sqrt{\nu_{k}^{t}}}$.
Similar to (74), we have
$\hskip-1.42262pt\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},j\in\mathcal{U}_{k}^{t})$
$\leq\delta^{\prime}+e^{\frac{c_{2}}{\sqrt{\bar{\nu}_{k}-r_{1}}}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime},j\in\mathcal{U}_{k}^{t})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(3)}.$ (80)
By substituting (74) into (89), we obtain the final expression,
$\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})\leq$ $e^{\frac{c_{2}}{\sqrt{\bar{\nu}_{k}-r_{1}}}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime})$
$+(1-p_{k,j})(\delta^{\prime}+\delta_{j}^{(3)}),$ (81)
thereby completing the proof.
Appendix D Proof of Theorem VI.4
When the combinator $\bm{c}_{f^{\prime}}$ fails to reconstruct the global model, the combination result is noisy, and the server should not learn the participation of client $j$ at the relaying client $k\in\mathcal{V}_{j}$. That is, from the aggregated result according to $f^{\prime}$-th combinator, $\bm{F}_{f^{\prime}}=\sum_{k\in\mathcal{V}^{t}}c_{f^{\prime},k}\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}(\Delta\bm{\Theta}_{m}^{t,I}+\bm{N}_{m}^{t})$ should not precisely reveal the identity of the $j$-th participating client at relay $k$. Let $\bm{F}_{f^{\prime}}^{t,\{k,-j\}}$ denote the combined results excluding the $j$-th contribution via relay $k$. It holds that
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\hskip-2.84526pt\in\mathcal{S})=$
$\mathrm{P_{r}}(\hskip-1.42262pt\{j\in\hskip-1.42262pt\mathcal{U}_{k}^{t}\}\hskip-1.42262pt\cap\hskip-1.42262pt\{k\in\hskip-1.42262pt\mathcal{V}^{t}\}\hskip-1.42262pt)\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\hskip-4.2679pt\in\hskip-1.42262pt\mathcal{S}|\{j\in\hskip-1.42262pt\mathcal{U}_{k}^{t}\}\hskip-1.42262pt\cap\hskip-1.42262pt\{k\in\hskip-1.42262pt\mathcal{V}^{t}\}\hskip-1.42262pt)$
$+\hskip-1.42262pt\mathrm{P_{r}}(\hskip-1.42262pt\{j\notin\mathcal{U}_{k}^{t}\}\hskip-1.42262pt\cup\hskip-1.42262pt\{k\notin\mathcal{V}^{t}\}\hskip-1.42262pt)\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-1.42262pt\mathcal{S}|\{j\notin\mathcal{U}_{k}^{t}\}\hskip-1.42262pt\cup\hskip-1.42262pt\{k\notin\hskip-1.42262pt\mathcal{V}^{t}\}\hskip-1.42262pt),$ (82)
where (82) follows that the scenarios where client $j$ fails to convey its local model update $\Delta\bm{\Theta}_{j}^{t}$ via relaying client $k$ either results from the event $\{j\notin\mathcal{U}_{k}^{t}\}$, where $j$ fails to share with client $m$; or $\{k\notin\mathcal{V}^{t}\}$, where $k$ fails to connect with the server.
Consider the events such that resulting variance $\nu_{k}^{t,-j}$ lies within the event $\mathcal{E}\triangleq\{\lvert\nu_{f^{\prime}}^{t,\{k,-j\}}-\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}\rvert\leq r_{2}\}$, where $r_{2}$ is chosen such that $\mathrm{P_{r}}(\mathcal{E}^{\mathrm{c}})<\delta^{\prime}$, $\delta^{\prime}\in(0,1)$, then
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\})$
$=$ $\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\},\mathcal{E})\mathrm{P_{r}}(\mathcal{E})$
$+\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t},\mathcal{E}^{c}\})\mathrm{P_{r}}(\mathcal{E}^{c})$
$\leq$ $\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\hskip-2.84526pt\in\hskip-1.42262pt\mathcal{S}|\{j\hskip-2.84526pt\in\mathcal{U}_{k}^{t}\}\hskip-1.42262pt\cap\hskip-1.42262pt\{k\in\hskip-2.84526pt\mathcal{V}^{t}\},\mathcal{E})\mathrm{P_{r}}(\mathcal{E})+\delta^{\prime}.$ (83)
For each realization of $\mathcal{U}_{k}^{t}=\mathcal{A}$ and $\mathcal{V}^{t}=\mathcal{B}$, similar to (67), the privacy loss is identical to that of the Gaussian mechanism, with variance $\nu_{f^{\prime}}^{t,\{k,-j\}}$. The $\ell_{2}$-sensitivity is bounded by
$\hskip 11.38109pt\sup_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \mathcal{B}\in[K],k\in\mathcal{B}\\ \lvert\nu_{f^{\prime}}^{t,\{k,-j\}}-\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}\rvert\leq r_{2}\end{subarray}}\hskip 0.0pt\left\lVert c_{f^{\prime},k}g_{k,j}\Delta\bm{\Theta}_{j}^{t}-c_{f^{\prime},k}g_{k,j}\bm{N}_{j}^{t}\right\rVert$
$\hskip 42.67912pt\leq\lvert c_{f^{\prime},k}g_{k,j}\rvert(R+\lVert\bm{N}_{j}^{t}\rVert)$
$\hskip 42.67912pt\leq\lvert c_{f^{\prime},k}g_{k,j}\rvert(R+\lambda\sqrt{D(1+\delta_{0})}),$ (84)
with probability $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$, where Chi-Square tail bound is applied in (84). Consequently, for each realization of $\mathcal{U}_{k}^{t}=\mathcal{A}$ and $\mathcal{V}^{t}=\mathcal{B}$, $\bm{F}_{f^{\prime}}^{t,\{k,-j\}}$ is $(\epsilon_{f^{\prime},\{k,j\}}^{\mathcal{A},4},\delta_{f^{\prime},\{k,j\}}^{(4)})$-differentially private, where $\epsilon_{f^{\prime},\{k,j\}}^{\mathcal{A},4}=\hskip-2.84526pt\left[2\log\left(\frac{1.25}{\delta_{f^{\prime},\{k,j\}}^{(4)}}\right)\right]^{\frac{1}{2}}\hskip-2.84526pt\frac{\lvert c_{f^{\prime},k}g_{k,j}\rvert(R+\lambda\sqrt{D(1+\delta_{0})})}{\sqrt{\nu_{f^{\prime}}^{t,\{k,-j\}}}}\hskip-2.84526pt\triangleq\frac{c_{4}}{\sqrt{\nu_{f^{\prime}}^{t,\{k,-j\}}}}$. Thus, it holds that
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\},\mathcal{E})$ (85a)
$\leq\hskip-14.22636pt\sum_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \mathcal{B}\in[K],k\in\mathcal{B}\\ \lvert\nu_{f^{\prime}}^{t,\{k,-j\}}-\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}\rvert\leq r_{2}\end{subarray}}\hskip-14.22636pt\mathrm{P_{r}}(\{j\in\mathcal{A}\}\cap\{k\in\mathcal{B}\})\Bigg{(}e^{\frac{c_{1}}{\sqrt{\nu_{f^{\prime}}^{t,\{k,-j\}}}}}\cdot$
$\hskip 28.45274pt\vskip-8.53581pt\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\in\mathcal{A}\}\cap\{k\in\mathcal{B}\})\hskip-1.42262pt+\hskip-1.42262pt\delta_{f^{\prime},\{k,j\}}^{(4)}\Bigg{)}$ (85b)
$\leq e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}-r_{2}}}}\cdot\hskip-14.22636pt\sum_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \mathcal{B}\in[K],k\in\mathcal{B}\\ \lvert\nu_{f^{\prime}}^{t,\{k,-j\}}-\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}\rvert\leq r_{2}\end{subarray}}\hskip-14.22636pt\mathrm{P_{r}}(\{j\in\mathcal{A}\}\cap\{k\in\mathcal{B}\})\cdot$
$\hskip 28.45274pt\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\in\mathcal{A}\}\cap\{k\in\mathcal{B}\})\hskip-1.42262pt+\hskip-1.42262pt\delta_{f^{\prime},\{k,j\}}^{(4)}$ (85c)
$\leq e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\},\mathcal{E})\hskip-1.42262pt$
$\hskip 170.71652pt+\hskip-1.42262pt\delta_{f^{\prime},\{k,j\}}^{(4)},$ (85d)
where (85c) holds since $\nu_{f^{\prime}}^{t,\{k,-j\}}>\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}-r_{2}$, (85d) is similar to (66).
Furthermore, according to (83), we have
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\})$
$\leq$ $\mathrm{P_{r}}(\mathcal{E})\bigg{(}e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\},\mathcal{E})$
$\hskip 128.0374pt+\delta_{f^{\prime},\{k,j\}}^{(4)}\bigg{)}+\delta^{\prime}$ (86a)
$\leq$ $e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\})$
$\hskip 142.26378pt+\delta_{f^{\prime},\{k,j\}}^{(4)}+\delta^{\prime}.$ (86b)
In addition, we have
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\notin\mathcal{U}_{k}^{t}\}\cup\{k\notin\mathcal{V}^{t}\})$
$\leq e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\notin\mathcal{U}_{k}^{t}\}\cup\{k\notin\mathcal{V}^{t}\}).$ (87)
Substitute (86b) and (87) into (82), we get
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\in\mathcal{S})\leq e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S})$
$\hskip 28.45274pt+\mathrm{P_{r}}(\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\})\left(\delta_{f^{\prime},\{k,j\}}^{(4)}+\delta^{\prime}\right),$ (88)
where $\mathrm{P_{r}}(\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\})=(1-p_{k})(1-p_{k,j})$. The proof is thereby completed.
When $\Delta\bm{\Theta}_{j}^{t}=u$ is perturbed to $\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime}$. The $\bm{F}_{f^{\prime}}^{t}$ should not fluctuate significantly. Notably, any relaying neighbor of client $j$ (including client $j$), say client $k$, can convey the perturbation to $\bm{F}_{f^{\prime}}^{t}$ if $c_{f^{\prime}k}\neq 0$. The perturbation does not affect $\bm{F}_{f^{\prime}}^{t}$ only if all relaying neighbors of client $j$ that successfully receive from $j$ either fail to connect to the server or correspond to a zero coefficient in $\bm{c}_{f^{\prime}}$. Denote the event that the perturbation of $\Delta\bm{\Theta}_{j}^{t}$ can cause a perturbation of $\bm{F}_{f^{\prime}}^{t}$ as event $\mathcal{J}$, then $\mathcal{J}=\{\exists k:k\in\mathcal{V}_{f^{\prime}},k\in\mathcal{V}^{t},j\in\mathcal{U}_{k}^{t}\}$.
For $\forall\bm{u},\bm{u}^{\prime}\in\mathbb{B}_{D}(R)$, it holds that
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})$ $=\mathrm{P_{r}}(\mathcal{J})\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},\mathcal{J})$
$\hskip-42.67912pt+\mathrm{P_{r}}(\mathcal{J}^{c})\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime},\mathcal{J}^{c}).$ (89)
Consider the events such that resulting variance $\nu_{f^{\prime}}^{t}$ lies within the event $\mathcal{E}\triangleq\{\lvert\nu_{f^{\prime}}^{t}-\bar{\nu}_{f^{\prime}}^{t}\rvert\leq r_{2}\}$, where $r_{2}$ is chosen such that $\mathrm{P_{r}}(\mathcal{E}^{\mathrm{c}})<\delta^{\prime}$, $\delta^{\prime}\in(0,1)$, then
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u},\mathcal{J})$ $=\mathrm{P_{r}}(\mathcal{E})\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u},\mathcal{J},\mathcal{E})$
$\hskip-45.5244pt+\mathrm{P_{r}}(\mathcal{E}^{c})\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},\mathcal{J},\mathcal{E}^{c})$ (90a)
$\hskip-56.9055pt\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u},\mathcal{J},\mathcal{E})$ (90b)
For each realization of $\mathcal{U}_{k}^{t}=\mathcal{A}$ and $\mathcal{V}^{t}=\mathcal{B}$, similar to (67), the privacy loss is identical to that of the Gaussian mechanism, with variance $\nu_{f^{\prime}}^{t}$. The $\ell_{2}$-sensitivity is bounded by
$\hskip 11.38109pt\sup_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{B}\in[K],\mathcal{B}\cap\mathcal{V}_{f^{\prime}}=\emptyset,\\ k\in\mathcal{B},\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \lvert\nu_{f^{\prime}}^{t}-\bar{\nu}_{f^{\prime}}^{t}\rvert\leq r_{2}\end{subarray}}\hskip 0.0pt\left\lVert\sum_{k\in\mathcal{B}}c_{f^{\prime},k}g_{k,j}(\bm{u}-\bm{u}^{\prime})\right\rVert$
$\hskip 42.67912pt\leq\sum_{k=1}^{K}\lvert c_{f^{\prime},k}g_{k,j}\rvert\cdot 2R.$ (91)
Consequently, for each realization of $\mathcal{U}_{k}^{t}=\mathcal{A}$ and $\mathcal{V}^{t}=\mathcal{B}$, $\bm{F}_{f^{\prime}}^{t}$ is $(\epsilon_{f^{\prime},j}^{\mathcal{A},\mathcal{B},5},\delta_{f^{\prime},j}^{(5)})$-differentially private, where $\epsilon_{f^{\prime},j}^{\mathcal{A},\mathcal{B},5}=\hskip-2.84526pt\left[2\log\left(\frac{1.25}{\delta_{f^{\prime},j}^{(5)}}\right)\right]^{\frac{1}{2}}\hskip-2.84526pt\frac{2R\sum_{k=1}^{K}\lvert c_{f^{\prime},k}g_{k,j}\rvert}{\sqrt{\nu_{f^{\prime}}^{t}}}\hskip-2.84526pt\triangleq\frac{c_{5}}{\sqrt{\nu_{f^{\prime}}^{t}}}$.
Thus, it holds that
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u},\mathcal{J},\mathcal{E})$
$=\sum_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{B}\in[K],\mathcal{B}\cap\mathcal{V}_{f^{\prime}}=\emptyset,\\ k\in\mathcal{B},\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \lvert\nu_{f^{\prime}}^{t}-\bar{\nu}_{f^{\prime}}^{t}\rvert\leq r_{2}\end{subarray}}\mathrm{P_{r}}(\mathcal{A},\mathcal{B})\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u},\mathcal{J}(\mathcal{A},\mathcal{B}),\mathcal{E})$ (92a)
$\leq\hskip 0.0pt\sum_{\hskip-28.45274pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{B}\in[K],\mathcal{B}\cap\mathcal{V}_{f^{\prime}}=\emptyset,\\ k\in\mathcal{B},\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \lvert\nu_{f^{\prime}}^{t}-\bar{\nu}_{f^{\prime}}^{t}\rvert\leq r_{2}\end{subarray}}\mathrm{P_{r}}(\mathcal{A},\mathcal{B})\big{(}e^{\epsilon_{f^{\prime},j}^{\mathcal{A},\mathcal{B},5}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J}(\mathcal{A},\mathcal{B}),\mathcal{E})$
$\hskip 170.71652pt+\delta_{f^{\prime},j}^{(5)}\big{)}$ (92b)
$\leq e^{\frac{c_{5}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J},\mathcal{E})+\delta_{f^{\prime},j}^{(5)}.$ (92c)
Similar to (83) and (86b), it holds that
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u},\mathcal{J})$
$\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\left(e^{\frac{c_{5}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J},\mathcal{E})+\delta_{f^{\prime},j}^{(5)}\right)$ (93a)
$\leq e^{\frac{c_{5}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J})+(\delta_{f^{\prime},j}^{(5)}+\delta^{\prime})$ (93b)
Additionally, it holds that
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J}^{c})$
$\leq e^{\frac{c_{5}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J}^{c})$ (94)
As a result,
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u})$
$\leq e^{\frac{c_{5}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime})+\mathrm{P_{r}}(\mathcal{J})(\delta_{f^{\prime},j}^{(5)}+\delta^{\prime}),$ (95)
where $\mathrm{P_{r}}(\mathcal{J})=1-\tilde{p}_{j,f^{\prime}}$, and $\tilde{p}_{j,f^{\prime}}$ denotes the probability of client $j$’s contribution not being conveyed by those participating in $\bm{c}_{f^{\prime}}$-combination.
Appendix E Proof of Theorem VI.5
Assume $\operatorname{Cov}[\Delta\bm{\Theta}_{k}^{t}])=\mathrm{Diag}(\bm{I}_{D}\zeta^{2})$, now we investigate the privacy protection for the identity of the participating client $j$ provided by other local model updates, in the successfully reconstructed global model $\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t}$. According to the Chi-squared tail bound, $\forall\delta_{0}$, it holds that
$\mathrm{P_{r}}\left(\left\lVert\Delta\bm{\Theta}_{j}^{t}\right\rVert\hskip-1.42262pt\leq\hskip-1.42262pt\zeta\sqrt{D(1+\delta_{0})}\right)\hskip-1.42262pt\geq\hskip-1.42262pt1\hskip-1.42262pt-\hskip-1.42262pte^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}.$ (96)
Thus, the $\ell_{2}$-sensitivity can be bounded by
$\left\lVert\frac{1}{K}\Delta\bm{\Theta}_{j}^{t}\right\rVert\leq\frac{\zeta}{K}\sqrt{D(1+\delta_{0})}$ (97)
with probability $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$. The effective noise $\frac{1}{K}\sum_{k\neq j}\Delta\bm{\Theta}_{k}^{t}\sim\mathcal{N}(0,\frac{K-1}{K}\zeta^{2}\bm{I}_{D})$, therefore, the conclusion of Gaussian mechanism applies. Specifically, $\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t}$ is $(\epsilon^{(6)},\delta^{(6)})$-differentially private in protecting the identity of client $j$ with probability $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$, where
$\epsilon^{(6)}=\frac{\frac{\zeta}{K}\sqrt{D(1+\delta_{0})}}{\frac{\sqrt{K-1}}{K}\zeta}\left[2\log\left(\frac{1.25}{\delta^{(6)}}\right)\right]^{\frac{1}{2}},$
i.e.,
$\epsilon^{(6)}=\sqrt{\frac{D(1+\delta_{0})}{K-1}}\left[2\log\left(\frac{1.25}{\delta^{(6)}}\right)\right]^{\frac{1}{2}}.$
When $\Delta\bm{\Theta}_{j}^{t}$ is perturbed from $\bm{u}$ to $\bm{u}^{\prime}$, the the $\ell_{2}$-sensitivity can be bounded by
$\left\lVert\frac{1}{K}(\bm{u}-\bm{u}^{\prime})\right\rVert\leq\frac{2\zeta}{K}\sqrt{D(1+\delta_{0})},$ (98)
with probability $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$. So $\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t}$ is $(\epsilon^{(7)},\delta^{(7)})$-differentially private in protecting the local data of client $j$ with probability $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$, where
$\epsilon^{(7)}=2\sqrt{\frac{D(1+\delta_{0})}{K-1}}\left[2\log\left(\frac{1.25}{\delta^{(7)}}\right)\right]^{\frac{1}{2}}.$ (99)
Appendix F Choices of Bernstein Parameters
Lemma F.1 (Bernstein’s Inequality).
Let $X$ be a zero-mean real-valued random variable, bounded by $M$, almost surely. Then, for $\forall r\geq 0$, we have
$\mathrm{P_{r}}(X\geq r)\leq\exp\left(-\frac{r^{2}}{\mathbb{E}[X^{2}]+\frac{2}{3}Mr}\right),$ (100)
or
$\mathrm{P_{r}}(\lvert X\rvert\geq r)\leq 2\exp\left(-\frac{r^{2}}{\mathbb{E}[X^{2}]+\frac{2}{3}Mr}\right).$ (101)
By choosing $\delta^{\prime}=2\exp\left(-\frac{r^{2}}{\mathbb{E}[X^{2}]+\frac{2}{3}Mr}\right)$, this is a quadratic in $r$ and $r$ can be solved accordingly, i.e.,
$r=\ln\left(\frac{2}{\delta^{\prime}}\right)\cdot\left(\frac{1}{3}M+\sqrt{\frac{1}{9}M^{2}+\frac{\mathbb{E}[X^{2}]}{\ln\left(\frac{2}{\delta^{\prime}}\right)}}\right).$ (102)
F-A Choice of $r_{1}$
Let $X_{k,1}=\nu_{k}^{t}-\bar{\nu}_{k}$. Then, $\mathbb{E}[X_{k,1}]=0$, and $\max\{X_{k,1}\}\leq\max\{\nu_{k}^{t}\}\leq\max_{k}\{\nu_{k}^{t}\}\triangleq M_{1}$. Due to Bernoulli r.v.s, $M_{1}$ is computable, through e.g., brute-force search. Moreover,
$\mathbb{E}[X_{k,1}^{2}]=\mathbb{E}[(\nu_{k}^{t}-\bar{\nu}_{k})^{2}]=\mathbb{E}[(\nu_{k}^{t})^{2}]-\bar{\nu}_{k}^{2},$ (103)
where $\bar{\nu}_{k}$ and $\mathbb{E}[(\nu_{k}^{t})^{2}]$ are given in (122) and (125b), respectively. Thereby, $r_{1}$ is chosen as
$r_{1}=\ln\left(\frac{2}{\delta^{\prime}}\right)\cdot\left(\frac{1}{3}M_{1}+\sqrt{\frac{1}{9}M_{1}^{2}+\frac{\mathbb{E}[X_{k,1}^{2}]}{\ln\left(\frac{2}{\delta^{\prime}}\right)}}\right).$ (104)
F-B Choice of $r_{2}$
Let $X_{k,2}=\nu_{f^{\prime}}^{t}-\bar{\nu}_{f^{\prime}}$. Then, $\mathbb{E}[X_{k,2}]=0$, and $\max\{X_{k,2}\}\leq\max\{\nu_{f^{\prime}}^{t}\}\leq\max_{f^{\prime}}\{\nu_{f^{\prime}}^{t}\}\triangleq M_{2}$. Due to Bernoulli r.v.s, $M_{2}$ is computable, through e.g., brute-force search. Moreover,
$\mathbb{E}[X_{k,2}^{2}]=\mathbb{E}[(\nu_{f^{\prime}}^{t}-\bar{\nu}_{f^{\prime}})^{2}]=\mathbb{E}[(\nu_{f^{\prime}}^{t})^{2}]-\bar{\nu}_{f^{\prime}}^{2},$ (105)
where $\bar{\nu}_{f^{\prime}}$ and $\mathbb{E}[(\nu_{f^{\prime}}^{t})^{2}]$ are given in (129) and (125b), respectively. Thereby, $r_{2}$ is chosen as
$r_{2}=\ln\left(\frac{2}{\delta^{\prime}}\right)\cdot\left(\frac{1}{3}M_{2}+\sqrt{\frac{1}{9}M_{2}^{2}+\frac{\mathbb{E}[X_{k,2}^{2}]}{\ln\left(\frac{2}{\delta^{\prime}}\right)}}\right).$ (106)
F-C Choice of $r_{3}$
According to Appendix E, we have $r_{3}=\zeta\sqrt{D(1+\delta_{0})}$.
Appendix G Proof of Theorem VIII.1
Under Assumption VIII.1, it holds that
$\mathbb{E}\left[\mathcal{L}(\bm{\Theta}^{t_{s}})\right]-\mathbb{E}\left[\mathcal{L}(\bm{\Theta}^{t_{s}-1})\right]$
$\leq\mathbb{E}\hskip-1.42262pt\left[\left\langle\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1}),\bm{\Theta}^{t_{s}}\hskip-1.42262pt-\hskip-1.42262pt\bm{\Theta}^{t_{s}-1}\right\rangle\right]\hskip-2.84526pt+\hskip-2.84526pt\frac{G}{2}\mathbb{E}\hskip-1.42262pt\left[\lVert\bm{\Theta}^{t_{s}}\hskip-1.42262pt-\hskip-1.42262pt\bm{\Theta}^{t_{s}-1}\rVert^{2}\right].$ (107)
Next, we bound the two terms in (107), respectively.
G-A Bounding the First Term
For the first term, we have
$\mathbb{E}\left[\left\langle\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1}),\bm{\Theta}^{t_{s}}-\bm{\Theta}^{t_{s}-1}\right\rangle\right]$
$=\mathbb{E}\hskip-2.84526pt\left[\hskip-2.84526pt\left\langle\hskip-2.84526pt\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1}),-\eta\frac{1}{K}\sum_{k=1}^{K}\sum_{i=1}^{R_{t_{s}}I}a_{i}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1},\bm{\xi}_{k}^{t_{s},i})\hskip-2.84526pt\right\rangle\hskip-2.84526pt\right]$
$=-\eta\hskip-2.84526pt\sum_{i=1}^{R_{t_{s}}I}\hskip-2.84526pta_{i}\mathbb{E}\hskip-2.84526pt\left[\hskip-2.84526pt\left\langle\hskip-2.84526pt\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1}),\frac{1}{K}\hskip-2.84526pt\sum_{k=1}^{K}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1})\hskip-2.84526pt\right\rangle\hskip-2.84526pt\right]$ (108a)
$\leq-\frac{1}{2}\eta\sum_{i=1}^{R_{t_{s}}I}a_{i}\mathbb{E}\left[\left\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\right\lVert^{2}\right]$
$\hskip 14.22636pt+\frac{1}{2}\eta\hskip-2.84526pt\sum_{i=1}^{R_{t_{s}}I}a_{i}\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\hskip-2.84526pt\left[\left\lVert\nabla\mathcal{L}(\bm{\Theta}_{k}^{t_{s}-1})\hskip-2.84526pt-\hskip-2.84526pt\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s},i-1})\right\lVert^{2}\right]$
$\hskip 14.22636pt-\frac{1}{2}\eta\sum_{i=1}^{R_{t_{s}}I}a_{i}\mathbb{E}\left[\left\lVert\frac{1}{K}\sum_{k=1}^{K}\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s},i-1})\right\lVert^{2}\right]$ (108b)
$\leq-\frac{1}{2}\eta\sum_{i=1}^{R_{t_{s}}I}a_{i}\mathbb{E}\left[\left\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\right\lVert^{2}\right]$
$\hskip 14.22636pt+\frac{1}{2}\eta\frac{G^{2}}{K}\sum_{i=1}^{R_{t_{s}}I}a_{i}\sum_{k=1}^{K}\mathbb{E}\left[\left\lVert\bm{\Theta}_{k}^{t_{s}-1}-\bm{\Theta}^{t_{s},i-1}\right\lVert^{2}\right],$ (108c)
where (108a) follows Assumption VIII.2, (108b) is due to the convexity of $\ell_{2}$-norm, and (108c) is due to Assumption VIII.1.
G-B Bounding the Second Term
For the second term, it holds that
$\frac{G}{2}\mathbb{E}\left[\lVert\bm{\Theta}^{t_{s}}-\bm{\Theta}^{t_{s}-1}\rVert^{2}\right]$
$=\frac{G}{2}\mathbb{E}\left[\left\lVert-\eta\frac{1}{K}\sum_{k=1}^{K}\sum_{i=1}^{R_{t_{s}}I}a_{i}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1},\bm{\xi}_{k}^{t,i})\right\rVert^{2}\right]$
$\leq\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\left[\left\lVert\sum_{i=1}^{R_{t_{s}}I}a_{i}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1},\bm{\xi}_{k}^{t_{s},i})\right\rVert^{2}\right]$ (109a)
$\leq\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i=1}^{R_{t_{s}}I}a_{i}\mathbb{E}\hskip-1.42262pt\left[\left\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1},\bm{\xi}_{k}^{t_{s},i})\right\rVert^{2}\right]$ (109b)
$=\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i=1}^{R_{t_{s}}I}a_{i}\Bigg{(}\mathbb{E}\bigg{[}\bigg{\lVert}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1},\bm{\xi}_{k}^{t_{s},i})$
$\hskip 14.22636pt-\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1})\bigg{\rVert}^{2}\bigg{]}+\mathbb{E}\left[\left\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1})\right\rVert^{2}\right]\Bigg{)}$ (109c)
$\leq\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i=1}^{R_{t_{s}}I}a_{i}\Bigg{(}\sigma^{2}+\mathbb{E}\bigg{[}\bigg{\lVert}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1})$
$\hskip 14.22636pt-\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s}-1})\bigg{\rVert}^{2}\bigg{]}+\mathbb{E}\left[\left\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s}-1})\right\rVert^{2}\right]\Bigg{)}$ (109d)
$\leq\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\sigma^{2}+\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i=1}^{R_{t_{s}}I}a_{i}\cdot$
$\bigg{(}G^{2}\mathbb{E}\bigg{[}\bigg{\lVert}\bm{\Theta}_{k}^{t_{s},i-1}-\bm{\Theta}^{t_{s}-1}\bigg{\rVert}^{2}\bigg{]}+\mathbb{E}\left[\left\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s}-1})\right\rVert^{2}\right]\Bigg{)}$ (109e)
$\leq\frac{\eta^{2}G}{2}R_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\left(\sigma^{2}+\beta^{2}\mathbb{E}\left[\left\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\right\rVert^{2}\right]+\kappa^{2}\right)$
$+\frac{\eta^{2}G^{3}}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i=1}^{R_{t_{s}}I}a_{i}\mathbb{E}\bigg{[}\bigg{\lVert}\bm{\Theta}_{k}^{t_{s},i-1}-\bm{\Theta}^{t_{s}-1}\bigg{\rVert}^{2}\bigg{]},$ (109f)
where (109a) and (109b) use Jensen’s Inequality, (109c) is due to unbiased estimation, (109d) follows Assumption VIII.2, (109e) follows Assumption VIII.1, and (109f) follows Assumption VIII.3.
G-C Immediate Result
Substitute (108c) and (109f) into (107), then substitute Lemma VIII.1, we obtain the immediate result for one round,
$\mathbb{E}\left[\left\langle\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1}),\bm{\Theta}^{t_{s}}-\bm{\Theta}^{t_{s}-1}\right\rangle\right]$
$\leq\left(-\frac{1}{2}\eta R_{t_{s}}\lVert\bm{a}\rVert_{1}+\eta^{2}GR_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\beta^{2}\right)\mathbb{E}\left[\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]$
$\hskip 14.22636pt+\eta^{2}GR_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\sigma^{2}+\eta^{2}GR_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\kappa^{2}$
$\hskip 14.22636pt+\left(\eta^{2}G^{3}R_{t_{s}}\lVert\bm{a}\rVert_{1}+\frac{1}{2}\eta G^{2}\right)\frac{2\eta^{2}R_{t_{s}}^{3}\lVert\bm{a}\rVert_{1}^{3}}{1-2\eta^{2}G^{2}R_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}}\cdot$
$\hskip 14.22636pt\hskip 14.22636pt\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\left[\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right],$ (110)
Using Assumption VIII.3, this further becomes
$\mathbb{E}\left[\left\langle\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1}),\bm{\Theta}^{t_{s}}-\bm{\Theta}^{t_{s}-1}\right\rangle\right]$
$\leq\left(-\frac{1}{2}\eta R_{t_{s}}\lVert\bm{a}\rVert_{1}+\eta^{2}GR_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\beta^{2}\right)\mathbb{E}\left[\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]$
$\hskip 14.22636pt+\eta^{2}GR_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\sigma^{2}+\eta^{2}GR_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\kappa^{2}$
$\hskip 14.22636pt+\left(\eta^{2}G^{3}R_{t_{s}}\lVert\bm{a}\rVert_{1}+\frac{1}{2}\eta G^{2}\right)\frac{2\eta^{2}R_{t_{s}}^{3}\lVert\bm{a}\rVert_{1}^{3}}{1-2\eta^{2}G^{2}R_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}}\cdot$
$\hskip 14.22636pt\hskip 14.22636pt\left(\beta^{2}\mathbb{E}\left[\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]+\kappa^{2}\right),$ (111)
G-D Average over Successful Rounds
After minor rearrangement, and averaging the immediate result over all $T_{s}$ successful rounds, we get
$\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}\bigg{(}\frac{1}{2}R_{t_{s}}\lVert\bm{a}\rVert_{1}-\eta GR_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\beta^{2}$-
$\hskip 2.84526pt\frac{2\eta R_{t_{s}}^{3}\lVert\bm{a}\rVert_{1}^{3}(\eta^{2}G^{3}R_{t_{s}}\lVert\bm{a}\rVert_{1}+\frac{1}{2}\eta G^{2})\beta^{2}}{1-2\eta^{2}G^{2}R_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}}\bigg{)}\mathbb{E}\left[\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]$
$\leq\frac{1}{\eta T_{s}}(\mathcal{L}(\bm{\Theta}^{0})-\mathcal{L}^{\star})+\frac{1}{2}\eta G\lVert\bm{a}\rVert_{1}^{2}\sigma^{2}\cdot\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}+$
$\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}\hskip-2.84526pt\bigg{(}\hskip-2.84526pt\eta G\lVert\bm{a}\rVert_{1}^{2}R_{t_{s}}^{2}\hskip-2.84526pt+\hskip-2.84526pt\frac{2\eta R_{t_{s}}^{3}\lVert\bm{a}\rVert_{1}^{3}(\eta^{2}G^{3}R_{t_{s}}\lVert\bm{a}\rVert_{1}\hskip-2.84526pt+\hskip-2.84526pt\frac{1}{2}\eta G^{2})}{1-2\eta^{2}G^{2}R_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}}\hskip-2.84526pt\bigg{)}\kappa^{2}.$ (112)
If the learning rate is chosen as $\eta=\frac{1}{L}\sqrt{\frac{K}{T}}$, by neglecting the non-dominant terms in the coefficients, (112) becomes
$\bigg{(}\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}\bigg{)}\cdot\frac{1}{2}\lVert\bm{a}\rVert_{1}\cdot\mathbb{E}\left[\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]$
$\leq\frac{G}{\sqrt{\frac{K}{T}}\cdot T_{s}}(\mathcal{L}(\bm{\Theta}^{0})-\mathcal{L}^{\star})+\frac{1}{2}\sqrt{\frac{K}{T}}\lVert\bm{a}\rVert_{1}^{2}\sigma^{2}\bigg{(}\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}\bigg{)}$
$\hskip 14.22636pt+\sqrt{\frac{K}{T}}\lVert\bm{a}\rVert_{1}^{2}\kappa^{2}\bigg{(}\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}\bigg{)}.$ (113)
G-E Relaxation with Probability Guarantee
By (120), (113) can be further bounded as follows with probability $0.9974$,
$\bigg{(}\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}\bigg{)}\cdot\frac{1}{2}\lVert\bm{a}\rVert_{1}\cdot\mathbb{E}\left[\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]$
$\leq\frac{G(\mathcal{L}(\bm{\Theta}^{0})-\mathcal{L}^{\star})}{\sqrt{\frac{K}{T}}\left(\frac{T}{e_{R}}-3\sqrt{\frac{\nu_{R}T}{e_{R}^{3}}}\right)}$
$+\frac{\lVert\bm{a}\rVert_{1}^{2}}{2}\sqrt{\frac{K}{T}}\left(\hskip-2.84526pte_{R^{2}}+3\sqrt{\frac{\nu_{R^{2}}}{\frac{T}{e_{R}}-3\sqrt{\frac{\nu_{R}T}{e_{R}^{3}}}}}\right)\hskip-2.84526pt\left(\sigma^{2}+\kappa^{2}\right).$ (114)
Substitute $e_{R},\nu_{R},e_{R^{2}},\nu_{R^{2}}$, we get
$\frac{1}{T_{s}}\hskip-2.84526pt\sum_{t_{s}=1}^{T_{s}}\mathbb{E}\left[\lVert\Delta\mathcal{L}(\bm{\Theta}^{t-1})\rVert^{2}\right]\hskip-1.42262pt\leq\hskip-1.42262pt\frac{1}{T_{s}}\hskip-2.84526pt\sum_{t_{s}=1}^{T_{s}}\hskip-2.84526ptR_{t_{s}}\mathbb{E}\left[\lVert\Delta\mathcal{L}(\bm{\Theta}^{t-1})\rVert^{2}\right]$
$\leq\frac{2G(\mathcal{L}(\bm{\Theta}^{0})-\mathcal{L}^{\star})}{\lVert\bm{a}\rVert_{1}\left((1-P_{O})\sqrt{KT}-3\sqrt{P_{O}(1-P_{O})}\right)}$
$\hskip 14.22636pt+2\lVert\bm{a}\rVert_{1}\sqrt{\frac{K}{T}}(\sigma^{2}+2\kappa^{2})\cdot\Bigg{(}\frac{1-P_{O}}{P_{O}}\mathrm{Li}_{-2}(P_{O})$
$\hskip 14.22636pt+3\sqrt{\frac{\frac{1-P_{O}}{P_{O}}\mathrm{Li}_{-4}(P_{O})-\left(\frac{1-P_{O}}{P_{O}}\mathrm{Li}_{-2}(P_{O})\right)^{2}}{(1-P_{O})T-3\sqrt{P_{O}(1P_{O})T}}}\Bigg{)}.$ (115)
By retaining only the dominant terms and substituting $\mathrm{Li}_{-2}(P_{O})=\frac{P_{O}(1+P_{O})}{(1-P_{O})^{3}}$, we complete the proof.
Appendix H Renewal Theory
According to the algorithm’s stopping rule, training terminates at the first round in which the global model is successful and the cumulative number of training rounds reaches the threshold $T$. Hence, we have
$T_{s}=\min\left\{n:\sum_{t=1}^{n}R_{t_{s}}\geq T\right\},$ (116)
where $R_{t_{s}}$ is the number of consecutive training rounds between two adjacent successful global model reconstructions. The $R_{t_{s}}$ renews after each successful reconstruction of the global model, thus $R_{t_{s}}\overset{i.i.d.}{\sim}\mathrm{Geo}(1-P_{O})$ for $\forall t_{s}$. Consequently, we have $e_{R}=\mathbb{E}[R_{t_{s}}]=\frac{1}{1-P_{O}}$ and $\nu_{R}=\mathrm{var}(R_{t_{s}})=\frac{P_{O}}{(1-P_{O})^{2}}$. We can also compute the expectation and variance of $R_{t_{s}}^{2}$, i.e.,
$e_{R^{2}}=\mathbb{E}[R_{t}^{2}]=\hskip-5.69054pt\sum_{R_{t}=1}^{+\infty}\hskip-2.84526ptR_{t}^{2}P_{O}^{R_{t}-1}(1-P_{O})=\hskip-2.84526pt\frac{1-P_{O}}{P_{O}}\mathrm{Li}_{-2}(P_{O}),$ (117a)
$\nu_{R^{2}}=\mathrm{var}(R_{t}^{2})=\mathbb{E}[R_{t}^{4}]-e_{R^{2}}^{2}$
$\hskip 14.22636pt=\frac{1-P_{O}}{P_{O}}\mathrm{Li}_{-4}(P_{O})-\left(\frac{1-P_{O}}{P_{O}}\mathrm{Li}_{-2}(P_{O})\right)^{2},$ (117b)
The polylogarithm function $\mathrm{Li}_{-v}(z)=\sum_{k=1}^{\infty}k^{v}z^{k}$ can be computed by $\mathrm{Li}_{-v}(z)=\left(z\frac{\partial(-\ln{(1-z)})}{\partial z}\right)^{v}\frac{z}{1-z}$.
By examining the step-by-step proof of the asymptotic properties of the standard renewal process, we find that these properties also hold for the expression in (116). Therefore, $T_{s}\approx\frac{T}{e_{R}}$. Then, according to CLT,
$\frac{T_{s}-\frac{T}{e_{R}}}{\sqrt{T}}\overset{d}{\rightarrow}\mathcal{N}(0,\frac{\nu_{R}}{e_{R}^{3}}),$ (118)
where $\overset{d}{\rightarrow}$ indicates convergence in distribution.
According to the strong law of large numbers, the empirical average of any function $f(R)$ over the first $T_{s}$ renewals converges to its expected value almost surely, i.e., $\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}\rightarrow\mathbb{E}[R_{t}^{2}]$. According to CLT,
$\sqrt{T_{s}}\left(\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}-e_{R^{2}}\right)\overset{d}{\rightarrow}\mathcal{N}(0,\nu_{R^{2}}).$ (119)
By calculating $\mathrm{cov}\left(\frac{T_{s}-\frac{T}{e_{R}}}{\sqrt{T}},\sqrt{T_{s}}\left(\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}-e_{R^{2}}\right)\right)$, we find that its value converges to 0, and the variance of the product decays as $\mathcal{O}(\frac{1}{T})$. Thus, for large $T$, (118) and (119) are nearly uncorrelated, although they are both generated from $R_{t_{s}}$. Since they are jointly Gaussian, this also implies asymptotic independence. Let $\mathcal{C}_{1}\triangleq\left\{\frac{T_{s}-\frac{T}{e_{R}}}{\sqrt{T}}>-3\sqrt{\frac{\nu_{R}}{e_{R}^{3}}}\right\}$, $\mathcal{C}_{2}\triangleq\left\{\sqrt{T_{s}}\left(\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}-e_{R^{2}}\right)<3\sqrt{\nu_{R^{2}}}\right\}$. According to $3\sigma$-rule, we have
$\mathrm{P_{r}}\left(\mathcal{C}_{1},\mathcal{C}_{2}\right)\approx 0.9987^{2}\approx 0.9974$ (120)
Equivalently, $\mathcal{C}_{1}=\left\{T_{s}>\frac{T}{e_{R}}-3\sqrt{\frac{\nu_{R}T}{e_{R}^{3}}}\right\}$, and $\mathcal{C}_{2}\triangleq\left\{\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}<e_{R^{2}}+3\sqrt{\frac{\nu_{R^{2}}}{T_{s}}}\right\}$.
Appendix I Calculation of Privacy Noise Variance
I-A Calculation of $\bar{\nu}_{k}$ and $\mathbb{E}[(\nu_{k}^{t})^{2}]$
Let us begin by computing the variance of the aggregated secret keys $\bm{M}_{k}^{t}$ at relaying client $k$. It can be verified that $\bm{M}_{k}^{t}=\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{N}_{m}^{t}=\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{k}^{t}$, so $\mathrm{var}(\bm{M}_{k}^{t})=\nu_{k}^{t}\bm{I}_{D}=\lVert\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}\bm{I}_{D}$ in (34). Similarly, we define $\bm{M}_{k}^{t,-j}=\sum_{m\in\mathcal{U}_{k}^{t}\setminus\{j\}}g_{k,m}\bm{N}_{m}^{t}=\sum_{m\neq j}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{k}^{t}$, and we have $\mathrm{var}(\bm{M}_{k}^{t,-j})=\nu_{k}^{t,-j}\bm{I}_{D}=\lVert\bm{\Lambda}_{k}^{t,-j}\bm{A}\rVert^{2}\bm{I}_{D}$ in (LABEL:eq:_agg_noise_no_j). The $\bm{\Lambda}_{k}^{t}$ depends on $\{\tau_{k,m}\}_{m=1}^{K}$, so the variance $\nu_{k}^{t}$ of the effective noise $\bm{M}_{k}^{t}$ is also a random variable concerning the network topology. It should be noted that due to correlation, the variance of $\bm{M}_{k}^{t}$ does not necessarily increase monotonously with the number of received updates $|\mathcal{U}_{k}^{t}\rvert$ at client $k$. The expected variance $\bar{\nu}_{k}$ is given by
$\bar{\nu}_{k}=\mathbb{E}_{\mathcal{U}_{k}^{t}}[\nu_{k}^{t}]=\mathbb{E}\left[\lVert\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}\right].$ (121)
Furthermore, we have
$\mathbb{E}\left[\lVert\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}\right]=\mathbb{E}\left[\bm{\Lambda}_{k}^{t}\bm{A}\bm{A}^{\top}(\bm{\Lambda}_{k}^{t})^{\top}\right]$
$=\operatorname{Tr}\left(\operatorname{Cov}[\bm{\Lambda}_{k}^{t}]\bm{A}\bm{A}^{\top}\right)+\mathbb{E}\left[\bm{\Lambda}_{k}^{t}\bm{A}\bm{A}^{\top}(\bm{\Lambda}_{k}^{t})^{\top}\right]$
$=\lambda^{2}\sum_{m=1}^{K}(1-p_{k,m})g_{m,k}^{2}+\bm{\Lambda}_{k}^{\mathrm{E}}\bm{A}\bm{A}^{\top}(\bm{\Lambda}_{k}^{\mathrm{E}})^{\top},$ (122)
where $\bm{\Lambda}_{k}^{\mathrm{E}}=\mathbb{E}[\bm{\Lambda}_{k}^{t}]=[(1-p_{k,1})g_{k,1},\cdots,(1-p_{k,K})g_{k,K}]$.
Similarly, we also have
$\bar{\nu}_{k}^{-j}=\mathbb{E}_{\mathcal{U}_{k}^{t}}[\nu_{k}^{t,-j}]=\mathbb{E}\left[\lVert\bm{\Lambda}_{k}^{t,-j}\bm{A}\rVert^{2}\right]\bm{I}_{D},$ (123)
and
$\mathbb{E}\left[\lVert\bm{\Lambda}_{k}^{t,-j}\bm{A}\rVert^{2}\right]$
$=\lambda^{2}\sum_{m=1}^{K}(1-p_{k,m})g_{m,k}^{2}+\bm{\Lambda}_{k}^{\mathrm{E},-j}\bm{A}\bm{A}^{\top}(\bm{\Lambda}_{k}^{\mathrm{E},-j})^{\top},$ (124)
where $\bm{\Lambda}_{k}^{\mathrm{E},-j}$ is identical to $\bm{\Lambda}_{k}^{\mathrm{E}}$ except the $j$-th element is set to 0.
It can be verified that $\nu_{k}^{t}$ can also be written in the form $\nu_{k}^{t}=\sum_{k_{1},k_{2}\in\mathcal{U}_{k}}C^{(k)}_{k_{1},k_{2}}\tau^{t}_{k,k_{1}}\tau^{t}_{k,k_{2}}$, where $C^{(k)}_{k_{1},k_{2}}=g_{k,k_{1}}[\bm{A}\bm{A}^{\top}]_{k_{1}k_{1}}g_{k,k_{2}}$. Following this, we have
$\left(\nu_{k}^{t}\right)^{2}$ $=\hskip-5.69054pt\sum_{k_{1},k_{2},k_{3},k_{4}\in\mathcal{U}_{k}}\hskip-5.69054ptC^{(k)}_{k_{1},k_{2}}C^{(k)}_{k_{3},k_{4}}\tau^{t}_{k,k_{1}}\tau^{t}_{k,k_{2}}\tau^{t}_{k,k_{3}}\tau^{t}_{k,k_{4}},$ (125a)
$\mathbb{E}\left[\left(\nu_{k}^{t}\right)^{2}\right]$ $=\hskip-5.69054pt\sum_{k_{1},k_{2},k_{3},k_{4}\in\mathcal{U}_{k}}\hskip-5.69054ptC^{(k)}_{k_{1},k_{2}}C^{(k)}_{k_{3},k_{4}}\mathbb{E}\left[\tau^{t}_{k,k_{1}}\tau^{t}_{k,k_{2}}\tau^{t}_{k,k_{3}}\tau^{t}_{k,k_{4}}\right],$ (125b)
where $\mathbb{E}\left[\tau^{t}_{k,k_{1}}\tau^{t}_{k,k_{2}}\tau^{t}_{k,k_{3}}\tau^{t}_{k,k_{4}}\right]$ is given by
$\mathbb{E}\left[\tau^{t}_{k,k_{1}}\tau^{t}_{k,k_{2}}\tau^{t}_{k,k_{3}}\tau^{t}_{k,k_{4}}\right]=\prod_{u\in\{k_{1}\}\cup\{k_{2}\}\cup\{k_{3}\}\cup\{k_{4}\}}(1-p_{k,u}),$ (126)
since $\tau^{t}_{k,k_{1}}$ and $\tau^{t}_{k,k_{2}}$ are independent Bernoulli r.v., $(\tau^{t}_{k,k_{1}})^{2}=\tau^{t}_{k,k_{1}}$, and $\{k_{1}\}\cup\{k_{2}\}\cup\{k_{3}\}\cup\{k_{4}\}$ ensures the uniqueness of the links in the product.
I-B Calculation of $\bar{\nu}_{f^{\prime}}$ and $\mathbb{E}[(\nu_{f^{\prime}}^{t})^{2}]$
The aggregated secret key according to $\bm{c}_{f^{\prime}}$ is
$\bm{N}_{F_{f^{\prime}}}^{t}$ $=\sum_{k\in\mathcal{V}^{t}}c_{f^{\prime},k}\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{N}_{m}^{t}$ (127a)
$=\sum_{k=1}^{K}c_{f^{\prime},k}\tau_{k}^{t}\sum_{m=1}^{K}\tau_{k,m}^{t}g_{k,m}\bm{N}_{m}^{t}$ (127b)
$=\sum_{k=1}^{K}c_{f^{\prime},k}\tau_{k}^{t}\sum_{m=1}^{K}\tau_{k,m}^{t}g_{k,m}\sum_{l=1}^{L}\alpha_{m,l}\bm{Z}_{l}^{t}$ (127c)
$=\sum_{l=1}^{L}\sum_{k=1}^{K}c_{f^{\prime},k}\tau_{k}^{t}\sum_{m=1}^{K}\tau_{k,m}^{t}g_{k,m}\alpha_{m,l}\bm{Z}_{l}^{t}.$ (127d)
Thus, $\mathrm{var}(\bm{N}_{F_{f^{\prime}}}^{t})=\nu_{f^{\prime}}^{t}\bm{I}_{D}$, where
$\nu_{f^{\prime}}^{t}$ $=\sum_{l=1}^{L}\left(\sum_{k=1}^{K}c_{f^{\prime},k}\tau_{k}^{t}\sum_{m=1}^{K}\tau_{k,m}^{t}g_{k,m}\alpha_{m,l}\right)^{2}$ (128a)
$=\sum_{l=1}^{L}\left(\sum_{k=1}^{K}c_{f^{\prime},k}\tau_{k}^{t}\bm{\Lambda}_{k}^{t}\bm{\alpha}_{l}\right)^{2}$ (128b)
$\hskip-8.53581pt=\sum_{l=1}^{L}\bm{\alpha}_{l}^{\top}\left(\sum_{k_{1},k_{2}}c_{f^{\prime},k_{1}}c_{f^{\prime},k_{2}}\tau_{k_{1}}^{t}\tau_{k_{2}}^{t}\bm{\Lambda}_{k_{1}}^{t}\bm{\Lambda}_{k_{2}}^{t}\right)\bm{\alpha}_{l}.$ (128c)
Thus, we have
$\bar{\nu}_{f^{\prime}}=\mathbb{E}\left[\nu_{f^{\prime}}^{t}\right]=\sum_{l=1}^{L}\bm{\alpha}_{l}^{\top}\left(\sum_{k_{1},k_{2}}c_{f^{\prime},k_{1}}c_{f^{\prime},k_{2}}\tau_{k_{1}}^{t}\tau_{k_{2}}^{t}\Sigma_{k}\right)\bm{\alpha}_{l},$
where $\Sigma_{k}=\mathbb{E}\left[\bm{\Lambda}_{k_{1}}^{t}\bm{\Lambda}_{k_{2}}^{t}\right]$ and is given by
$[\Sigma_{k}]_{ij}=\begin{cases}g_{k,i}^{2}(1-p_{k,i}),\;\;\;\;\;\text{if}\;i=j\\ \left[(\bm{\Lambda}_{k}^{\mathrm{E}})^{\top}\bm{\Lambda}_{k}^{\mathrm{E}}\right]_{ij}\;\;\;\;\;\text{if}\;i\neq j\end{cases}.$ (130)
Next, let us calculate $\mathbb{E}\left[(\nu_{f^{\prime}}^{t})^{2}\right]$. The (128a) can be rewritten as
$\nu_{f^{\prime}}^{t}=\sum_{l=1}^{L}\left(\sum_{k_{1},m_{1}}c_{f^{\prime},k_{1}}g_{k_{1},m_{1}}\tau_{k_{1}}^{t}\tau_{k_{1},m_{1}}^{t}\alpha_{m_{1},l}\right)^{2}$ (131a)
$=\sum_{l=1}^{L}\sum_{k_{1},m_{1}}\sum_{k_{2},m_{2}}c_{f^{\prime},k_{1}}c_{f^{\prime},k_{2}}g_{k_{1},m_{1}}g_{k_{2},m_{2}}\alpha_{m_{1},l}\alpha_{m_{2},l}$
$\hskip 113.81102pt\cdot\tau_{k_{1}}^{t}\tau_{k_{1},m_{1}}^{t}\tau_{k_{2}}^{t}\tau_{k_{2},m_{2}}^{t}.$ (131b)
Let $A_{k_{1},m_{1}}^{f^{\prime}}=c_{f^{\prime},k_{1}}g_{k_{1},m_{1}}$, $t(k_{1},m_{1})=\tau_{k_{1}}^{t}\tau_{k_{1},m_{1}}^{t}$, and $H_{m_{1},m_{2}}=\sum_{l=1}^{L}\alpha_{m_{1},l}\alpha_{m_{2},l}$ (131b) can be simplified to
$\nu_{f^{\prime}}^{t}=\sum_{k_{1},m_{1}}\sum_{k_{2},m_{2}}A_{k_{1},m_{1}}^{f^{\prime}}A_{k_{2},m_{2}}^{f^{\prime}}H_{m_{1},m_{2}}t(k_{1},m_{1})t(k_{2},m_{2}).$ (132)
Thus,
$(\nu_{f^{\prime}}^{t})^{2}=\hskip-5.69054pt\sum_{k_{1},m_{1}}\sum_{k_{2},m_{2}}\sum_{k_{3},m_{3}}\sum_{k_{4},m_{4}}A_{k_{1},m_{1}}^{f^{\prime}}A_{k_{2},m_{2}}^{f^{\prime}}A_{k_{3},m_{3}}^{f^{\prime}}A_{k_{4},m_{4}}^{f^{\prime}}$
$\hskip 14.22636pt\cdot H_{m_{1},m_{2}}H_{m_{3},m_{4}}t(k_{1},m_{1})t(k_{2},m_{2})t(k_{3},m_{3})t(k_{4},m_{4}).$ (133)
Thus,
$\mathbb{E}\left[(\nu_{f^{\prime}}^{t})^{2}\right]\hskip-2.84526pt=\hskip-5.69054pt\sum_{k_{1},m_{1}}\hskip-1.42262pt\sum_{k_{2},m_{2}}\hskip-1.42262pt\sum_{k_{3},m_{3}}\hskip-1.42262pt\sum_{k_{4},m_{4}}A_{k_{1},m_{1}}^{f^{\prime}}A_{k_{2},m_{2}}^{f^{\prime}}A_{k_{3},m_{3}}^{f^{\prime}}A_{k_{4},m_{4}}^{f^{\prime}}$
$\hskip 5.69046pt\cdot H_{m_{1},m_{2}}H_{m_{3},m_{4}}\mathbb{E}\left[t(k_{1},m_{1})t(k_{2},m_{2})t(k_{3},m_{3})t(k_{4},m_{4})\right],$ (134)
where
$\mathbb{E}\left[t(k_{1},m_{1})t(k_{2},m_{2})t(k_{3},m_{3})t(k_{4},m_{4})\right]$
$=\prod_{\begin{subarray}{c}u\in\{k_{1}\}\cup\{k_{2}\}\\ \cup\{k_{3}\}\cup\{k_{4}\}\end{subarray}}(1-p_{u})\prod_{\begin{subarray}{c}(u,v)\in\{(k_{1},m_{1})\}\cup\{(k_{2},m_{2})\}\\ \cup\{(k_{3},m_{3})\}\cup\{(k_{4},m_{4})\}\end{subarray}}(1-p_{u,v}),$ (135)
since $\mathbb{E}\left[t(k_{1},m_{1})\right]=(1-p_{k_{1}})(1-p_{k_{1},m_{1}})$, $t(k_{1},m_{1})=t(k_{2},m_{2})$ for $k_{1}=k_{2}$, $m_{1}=m_{2}$, and $\cup$ ensures the uniqueness of the set elements.