### v4

# Coding-Enforced Robust Secure Aggregation for Federated Learning Under Unreliable Communication

Abstract
This work studies privacy-preserving federated learning (ppFL) under unreliable communication. In ppFL, zero-sum privacy noises enables privacy protection without sacrificing model accuracy, effectively overcoming the privacy-utility trade-off. However, in practice, unreliable communication can randomly disrupt the coordination of zero-sum noises, leading to aggregation errors and unpredictable partial participation, which severely harm the model accuracy and learning performance. To overcome these challenges, we propose a robust coding-enforced structured secure aggregation method, termed secure cooperative gradient coding (SecCoGC), which enables exact reconstruction of the global model under unreliable communication while allowing for arbitrarily strong privacy preservation. In this paper, a complete problem formulation and constructions of real-field zero-sum privacy noise are presented, and fairness is introduced as a privacy metric. Privacy across all protocol layers in SecCoGC is evaluated, accounting for the correlation among privacy noises and their linear combination under unreliable communication. Moreover, a distinct convergence analysis for the FL algorithm with a binary outcome for global model recovery is provided. Experimental results demonstrate that SecCoGC achieves strong resilience to unreliable communication while maintaining varying levels of privacy preservation, yielding test accuracy improvements of up to $20\%$-$70\%$ over existing benchmark methods.
I Introduction
Federated learning (FL) enables numerous clients to collaboratively train models based on their local datasets, enhancing user privacy and reducing the volume of the transmitted data [1, 2]. However, the adversary can still infer sensitive information that is captured by local model updates, making FL susceptible to various model-based attacks, e.g., inversion attacks (IA) and membership inference attacks (MIA), etc.. To defend against these and other attacks, a widely adopted strategy in privacy-preserving FL (ppFL) is to inject privacy noise into local models to obscure sensitive information, through, e.g., the well-known Gaussian mechanism [3]. A number of studies confirm the effectiveness of independent Gaussian privacy noises with provable convergence guarantees [4, 5]. However, a strong privacy guarantee requires substantial noise at each client, leading to significant noise accumulation at the server, which severely impairs model accuracy and convergence. This inherent tension is called the utility-privacy trade-off. Both privacy leakage and model accuracy are significant concerns in FL. To mitigate the conflict between these two metrics, some researchers investigate the use of correlated privacy noises instead [6, 7]. Such designs allow for partial cancellation of randomness and mitigate the amplification issue of the accumulated privacy noise at the server, yet they are still constrained by the privacy-utility trade-off in nature.
Secure aggregation, which enables the server to learn the accurate aggregated model without observing the local ones, provides a promising solution to overcome the privacy-utility trade-off. Most studies on secure aggregation are based on homomorphic encryption, Shamir’s Secret Sharing (SSS), multi-party computation, zero-sum discrete and analog noises, and hybrid combinations of these techniques [8, 9, 10, 11, 12, 13, 14, 15, 16]  [^1] Zero-sum noise is considered a secure aggregation method, as it permits the server to obtain an accurate global model without revealing the local models [13, 14, 16]. . In these methods, the correct reconstruction of the global model requires that the coordinated encryption structure is reliable, which is highly dependent on reliable and predictable communication. However, in practice, communication channels are often unreliable due to real-world imperfections, e.g., fading and shadowing. Such communication uncertainty will result in random and unpredictable variations in both the pattern of participating clients and their cardinality, making the problem of secure aggregation particularly challenging under unreliable communication. Recent advances in dropout-resilient secure aggregation primarily build upon a combination of SSS with gradient coding (GC) on coded datasets [17], pairwise masking [9], and multi-round strategic secure communication scheme [13]. Despite their significant contributions, [17] incurs high encoding and decoding complexity, [9] allows the server to learn partial sums of the local models beyond the intended global model, [9, 13] suffers from the negative impact of partial participation, as will be discussed shortly.
Another critical challenge in FL posed by unreliable communication is the persistent, strict sub-optimality arising from random partial participation. The data generated at the network edge is inherently uncontrollable and typically follows non-independent and identically distributed (non-IID) patterns across clients. For this reason, the local models tend to deviate largely from the global model, referred to as client drift[18]. Unreliable communication further intensifies this issue by introducing statistically uneven local contributions in the global aggregation, which guides the global model towards a stationary point far from the true optimum, referred to as objective inconsistency [19, 20].
In ppFL under unreliable communication conditions, the aforementioned phenomena become even more pronounced due to the mutual coupling between communication and privacy noise. First, the injected privacy noise will randomly perturb both the convergence direction and the effective step size, leading to an unstable optimization trajectory. When such instability interacts with unreliable communication, the resulting optimization trajectory becomes highly unpredictable and heterogeneous. Second, under unreliable communication, the variance of the aggregated noise becomes highly random and unstable. With the correlated privacy noise design, the aggregated noise at the server can grow substantially, as the correlated components fail to cancel effectively as intended. The combined effect of these facts can severely degrade the global model accuracy and impair training performance, and in extreme cases, may even cause divergence of FL.
Existing methods to mitigate the adverse effects of unreliable communication include heterogeneity-aware client sampling [21], adaptive resource allocation [20], heterogeneity-aware aggregation [22, 23], collaborative relaying [24, 25], etc. But they all rely on precise prior information to ensure their effectiveness. In contrast, [26, 27, 28] eliminates this need by appropriate accumulation mechanisms. However, they are prone to divergence, as the accumulated gradient norm can grow excessively after several consecutive transmission failures. Another line of work employs coding strategies to mitigate stragglers in FL without requiring prior information, including the usage of coded networks to reduce straggler occurrence [29], the combination of dataset-replication-based GC and coded datasets [17], and gradient-sharing-based cooperative gradient coding (CoGC) [30]. While extensive efforts have focused on enhancing robustness against unreliable communication, these methods generally leave privacy preservation unaddressed. As a result, a substantial gap remains between straggler-resilient design and robust privacy protection, highlighting a critical challenge that must be addressed.
I-A Our Contributions
To preserve both global model utility and user privacy under unreliable communication, this paper proposes a coding-enforced structured aggregation method, which simultaneously ensures the accurate computation of the global model and the complete cancellation of the zero-sum privacy noises, thereby significantly improving learning performance while allowing for arbitrarily strong privacy. Apart from this, the binary outcome inherent in the GC decoding mechanism provides an additional strong privacy guarantee, ensuring that the server aggregation is either intact or not meaningful. Both rigorous theoretical analysis and extensive experimental validations are applied to support our findings. The key contributions are summarized below.
(C1) A complete problem formulation  [^2] Notably, [16] presents an essential covariance-based construction for zero-sum noise in real field, which is advantageous in preset covariance and noise power. But a complete problem formulation is still lacking. of real-field zero-sum noise is present. To the best of our knowledge, this work first discusses the design principle, conditions, and constructions of real-field zero-sum noises. Moreover, it first introduces fairness  [^3] In this work, fairness is defined as allocation of privacy preservation levels in proportion to the clients’ learning weights [^4] The fairness concern can be further explored through algorithmic and system-level approaches, such as fairness-aware privacy budget allocation, adaptive client scheduling, and resource management strategies in FL system designs. These topics are left open for future research. as a performance metric in user privacy preservation. .
(C2) We propose a novel robust coding-enforced secure aggregation method, termed SecCoGC, which leverages coding techniques to enforce a determined aggregation structure over local models. SecCoGC remains resilient to unreliable communication in the following aspects:
  • The privacy noise is enforced to fully cancel out through the predetermined coded aggregation structure, regardless of its privacy level. This ensures the accuracy of the global model under unreliable communication, while allowing for arbitrarily strong privacy.
  • The global model is guaranteed to be formed correctly based on all local models, ensuring the optimality of FL algorithms and avoiding the negative effect arising from the unpredictable partial participation patterns of clients.
To the best of our knowledge, no prior work has applied coding schemes to ensure robust, secure aggregation in FL under unreliable communication  [^5] In [31], the authors introduce a secure computation scheme for linearly separable functions. While it shares some conceptual similarity with our approach, the focus and distributed learning setting differ significantly..
(C3)Multi-view and in-depth theoretical privacy analyses are provided for SecCoGC across all protocol layers under local differential privacy (LDP), covering both successful and failed global model reconstruction. Our analysis tackles four distinctive challenges:
  • In cooperative networks, a local update can be delivered to the server both directly and indirectly through neighboring nodes. This multi-path delivery introduces a substantial conceptual challenge in analyzing local differential privacy.
  • The analysis captures the correlation structure among privacy noises and their linear aggregation over intermittent networks, which poses significant challenges to LDP analysis, as the variance of the aggregated correlated noise becomes a random variable determined by network intermittency.
  • The correlation among secret keys enables inference based on a client’s secret key, and this leakage is properly accounted for in our analysis.
  • Unreliable communication leads to the simultaneous absence of the local model updates and the privacy noise, causing the sensitivity function to be unbounded.
To the best of our knowledge, no prior work has ever considered or tackled the above challenges. Our analysis offers valuable guidance for future privacy studies involving unreliable communication or cooperative networks.
(C4) A unique non-convex convergence analysis with an arbitrary choice of local solver is provided for SecCoGC and FL algorithms that result in binary recovery of the global model. Distinctively, both the number of consecutive training rounds between two adjacent successful reconstructions of the global model, and the total number of training rounds, are dependent random variables, which significantly complicates the convergence analysis. Our theoretical results yield a standard convergence rate of $\mathcal{O}\left(\frac{1}{\sqrt{T}}\right)$ with probability $99.74\%$, holding under arbitrary local solvers and arbitrary privacy levels.
I-B Paper Organization
The remainder of this paper is organized as follows. Section II presents the system model and preliminaries. Section IV presents a complete formulation of real-field zero-sum noise. Section III introduces the proposed SecCoGC. Section V evaluates LDP across all layers in SecCoGC. Section VI analyses the convergence behavior of FL algorithms with binary global construction outcomes. Section VII provides extensive performance comparisons of the proposed method over multiple benchmarks. Conclusions are drawn in Section VIII, and supporting proofs are provided in the appendices of the supplementary file [32] [^6] The supplementary file can be found at https://arxiv.org/abs/2507.07565..
II System Model and Preliminaries
II-A Federated Learning (FL)
Consider a FL system consisting of $K$ clients, each holding local datasets $\mathcal{D}_{k}$ on it, coordinated by a central server. The goal is to learn a global model $\bm{\Theta}\in\mathbb{R}^{D}$ that minimizes the global objective function (GOF) defined in (1):
$\min_{\bm{\Theta}}\mathcal{L}(\bm{\Theta})\triangleq\sum_{k=1}^{K}w_{k}\mathcal{L}_{k}(\bm{\Theta},\mathcal{D}_{k}),$ (1)
through iterative training and communication. The learning weight $w_{k}$ measures the importance of clients, satisfying $\sum_{k=1}^{K}w_{k}=1$. W.L.O.G., this paper assumes $w_{k}=\frac{1}{K}$. The local objective function (LOF) that measures the accuracy of the model $\bm{\Theta}$ on local datasets $\mathcal{D}_{k}$ is defined by the average loss on all training samples in $\mathcal{D}_{k}$,
$\mathcal{L}_{k}(\bm{\Theta},\mathcal{D}_{k})\triangleq\frac{1}{\lvert\mathcal{D}_{k}\rvert}\sum_{\xi\in\mathcal{D}_{k}}l(\bm{\Theta},\xi),$ (2)
where $l(\bm{\Theta},\xi)$ is the sample loss of the model $\bm{\Theta}$ evaluated on a single training sample $\xi$.
In the $t$-th training round, where $t\in[T]$, all clients are initialized with the latest global model, i.e., $\bm{\Theta}_{k}^{t,0}=\bm{\Theta}^{t-1}$. The local training on client $k$ at the $i$-th iteration is given by
$\bm{\Theta}_{k}^{t,i}=\bm{\Theta}_{k}^{t,i-1}-\eta a_{i}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t,i-1},\bm{\xi}_{k}^{t,i}),i\in[I]$ (3)
where $\eta$ is the learning rate, $I$ is the number of local iterations, $a_{i}$ characterizes the type of local solver, $\bm{\xi}_{k}^{t,i}$ is the data patch containing $n_{p}$ data samples extracted from $\mathcal{D}_{k}$ with $\lvert\bm{\xi}_{k}^{t,i}\rvert=n_{p}$. The gradient $\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t,i-1},\bm{\xi}_{k}^{t,i})$ computed on the data patch $\bm{\xi}_{k}^{t,i}$ is defined by the averaged gradients over all data samples $\xi\in\bm{\xi}_{k}^{t,i}$. In matrix form, the local model update $\Delta\bm{\Theta}_{k}^{t,I}=\bm{\Theta}_{k}^{t,I}-\bm{\Theta}^{t}$ can be expressed by
[…]
With full participation of clients, the server updates $\bm{\Theta}^{t}$ as
$\bm{\Theta}^{t}=\bm{\Theta}^{t-1}+\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t,I}.$ (5)
II-B Threat Model
Our threat model follows a standard assumption adopted in multi-party computation (MPC) [9, 33, 4]. The server and clients are assumed honest-but-curious: they faithfully execute the protocol yet may attempt to infer other local models based on publicly observable information and their local knowledge. To be more specific, the clients may try to learn the local model or the participation of a client by exploiting observable data, e.g., the received messages from its neighbors or the computed partial sums, and correlations among secret keys. The server may try to infer the local model, the participation of a client at another relaying client, or the participation of a client based on the combined partial sum using any predefined combinator.
II-C Communication Model
The inter-user and uplink communication channels are assumed to be orthogonal and modeled as independent Bernoulli r.v.s. The link $\tau_{k}$, from client $k$ to the sever, is modeled by $\tau_{k}\sim\mathrm{Ber}(1-p_{k})$, the link $\tau_{k,m}$, from client $m$ to client $k$, is modeled by $\tau_{k,m}\sim\mathrm{Ber}(1-p_{k,m})$, where $p_{k}$ and $p_{k,m}$ are the probabilities when the channel is in outage. The $\tau_{k}=1$ and $\tau_{k,m}=1$ implies a successful communication and perfect recovery, $\tau_{k}=0$ and $\tau_{k,m}=0$ implies a complete failure. For $\forall k$, $p_{k,k}=0$, $\tau_{k,k}=1$, since client $k$ has access to its own local model and requires no transmission. For $k\neq m$, if no transmission occurs from client $m$ to client $k$, then $p_{k,m}=1$, $\tau_{k,m}=0$, as clients do not have direct access to each other’s local models. Moreover, any link is assumed to be statistically independent from each other, i.e., $\mathrm{(i)}\;\forall k_{1}\neq k_{2}:\tau_{k_{1}}\perp\tau_{k_{2}}$; $\mathrm{(ii)}\;\forall(k_{1},m_{1})\neq(k_{2},m_{2}):\tau_{k_{1},m_{1}}\perp\tau_{k_{2},m_{2}}$; and $\mathrm{(iii)}\;\forall k_{1},k_{2},m:\tau_{k_{1}}\perp\tau_{k_{2},m}$.
II-D Cyclic Gradient Codes
The cyclic gradient code is formed by pair-wise cyclic matrices $\bm{G}$ and $\bm{C}$ that are devised such that $\bm{C}\bm{G}=\mathbf{1}_{f\times K}$, where $f={K\choose s}$. In particular, $\bm{G}$ is the allocation matrix of size $K\times K$ in a cyclic pattern with $s$ nonzero entries per row, and $\bm{C}$ is the combination matrix of size $f\times K$, where each row represents a unique pattern with exactly $s$ zeros per row, encompassing all such possible patterns. The structure ensures that, even if up to $s$ rows are missing from $\bm{G}$, its resulting product with $\bm{C}$ still includes at least one all-one row. The construction for GC codes can be found in [34, Algo. II].
III The Proposed Method: Secure Cooperative Gradient Coding (SecCoGC)
[…]
           Any two: $\bm{\Theta}_{1}+\bm{\Theta}_{2}+\bm{\Theta}_{3}$ $\mathcal{D}_{1}$$\mathcal{D}_{2}$$\mathcal{D}_{3}$ $\frac{1}{2}\bm{\Theta}_{1}+\bm{\Theta}_{2}$ $\frac{1}{2}\bm{N}_{1}-\bm{N}_{2}$ $\bm{\Theta}_{2}-\bm{\Theta}_{3}$ $\bm{N}_{2}-\bm{N}_{3}$ $\frac{1}{2}\bm{\Theta}_{1}+\bm{\Theta}_{3}$ $\frac{1}{2}\bm{N}_{1}+\bm{N}_{3}$ $\bm{\Theta}_{1}+\bm{N}_{1}$$\bm{\Theta}_{2}+\bm{N}_{2}$$\bm{\Theta}_{3}+\bm{N}_{3}\endlxSVG@picture$
Figure 1: An overview of the proposed SecCoGC method with simplified notations, when $M=3$, $s=1$.
Each client $k\in\{1,\cdots,K\}\triangleq[K]$ holds a raw local update $\Delta\bm{\Theta}_{k}^{t,I}$ and a secret key $\bm{N}_{k}^{t}$, depicted in Fig. 1.  [^7] In this paper, secret key and privacy noise are used interchangeably. [^8] The secret keys are generated and distributed by a key management server, ensuring that each client has access only to its own keys, without knowledge of their correlations or of any key held by other clients. The secret keys $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ are designed to sum to zero,
$\sum_{k=1}^{K}\bm{N}_{k}^{t}=0.$ (6)
To protect data during transmission, each client applies additive masking to its raw local model update,
$\bm{Y}_{k}^{t}=\Delta\bm{\Theta}_{k}^{t,I}+\bm{N}_{k}^{t}.$ (7)
Then, each client $k$ transmits $\bm{Y}_{k}^{t}$ to its neighbors $m\in\mathcal{V}_{k}$, where $\mathcal{V}_{k}\triangleq\{m:g_{m,k}\neq 0\}$ (including client $k$ itself), through unreliable link $\tau_{m,k}^{t}$.
In the meanwhile, client $k$ also acts as a relay, listening to its neighboring client $m\in\mathcal{U}_{k}$, where $\mathcal{U}_{k}\triangleq\{m:g_{k,m}\neq 0\}$ (including client $k$ itself). Client $k$ may or may not be able to receive information from client $m$, depending on the realization of link $\tau_{k,m}^{t}$, where
[…]
Let $\mathcal{U}_{k}^{t}$ denote the set of updates that client $k$ can successfully learn (including its own updates) at the $t$-th round, i.e., $\mathcal{U}_{k}^{t}=\{m:\tau_{k,m}^{t}=1\}$. Then, client $k$ computes the product $g_{k,m}\bm{Y}_{m}^{t}$ for each $m\in\mathcal{U}_{k}^{t}$, and sums these products to compute the partial sum, i.e.,
$\bm{S}_{k}^{t}=\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{Y}_{m}^{t}=\sum_{m=1}^{K}\tau_{k,m}^{t}\cdot g_{k,m}\bm{Y}_{m}^{t}.$ (9)
If $\mathcal{U}_{k}^{t}=\mathcal{U}_{k}$, client $k$ can successfully decode the message from all of its neighbors, then $\bm{S}_{k}^{t}$ is referred to as a complete partial sum, since the coefficients in the partial sum $\bm{S}_{k}^{t}$ fully match those in $\bm{G}$.
Then, client $k$ sends $\bm{S}_{k}^{t}$ (not necessarily complete) to the server through link $\tau_{k}^{t}$, where
[…]
Let $\mathcal{U}^{t}$ denote the set of complete partial sums arrived at the server, i.e., $\mathcal{U}^{t}=\{k:\mathcal{U}_{k}^{t}=\mathcal{U}_{k}\}\cap\{k:\tau_{k}^{t}=1\}$, the server aggregation can be summarized below.
[…]
    Combinator: Let $\bm{c}_{f_{t}}$ denote the $f_{t}$-th row in the combination matrix $\bm{C}$. The combinator $\bm{c}_{f_{t}}$ corresponds to $\mathcal{U}^{t}$ should satisfy
    $\forall k:\;c_{f_{t},k}\neq 0\Rightarrow k\in\mathcal{U}^{t}.$ (11)
    Such $\bm{c}_{f_{t}}$ exists iff $\lvert\mathcal{U}^{t}\rvert\geq K-s$.
[…]
    Server Aggregation: If such $\bm{c}_{f_{t}}$ exists, the server can compute the global model by combining these complete partial sums according to $\bm{c}_{f_{t}}$,
    $\bm{\Theta}^{t}$ $=\bm{\Theta}^{t-1}+\frac{1}{K}\sum_{k=1}^{K}c_{f_{t},k}\bm{S}_{k}^{t}.$ (12)
It can be verified that the server aggregation in (12) gives exactly the global model pursued by (5) without introducing any additional ambiguity, i.e.,
$\frac{1}{K}\sum_{k=1}^{K}c_{f_{t},k}\bm{S}_{k}^{t}$ $=\frac{1}{K}\sum_{k=1}^{K}\left(\Delta\bm{\Theta}_{k}^{t,I}+\bm{N}_{k}^{t}\right)=\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t,I}.$ (13)
If the server receives at least KsK-s complete partial sums, it can fully eliminate the privacy noise and accurately reconstruct the global model. Otherwise, clients proceed to the next round using their latest local models by setting $\bm{\Theta}_{k}^{t+1,0}=\bm{\Theta}_{k}^{t,I}$ in (3).
Stopping Rule: The training terminates when the total number of executed rounds reaches $T$, with the final round resulting in a successful reconstruction of the global model. The complete training procedure is detailed in Algorithm 1.
The proposed SecCoGC ensures the correct computation of the global model and reliable cancellation of secret keys in heterogeneous networks through a fixed aggregation structure predetermined by the coding scheme. It not only ensures the optimality of FL algorithms but also protects the user’s privacy at all protocol layers.
Remark III.1 (Encoding and Decoding Complexity).
During client-to-client communication, each client needs to encode its own local updates for transmission and decode $s$ messages from its $s$ neighbors. During client-to-server transmission, at most $M$ clients encode their complete partial sums for transmission, and the server decodes from at most $M$ clients. In total, each client performs $2$ encoding operations and $s$ decoding operations, the server performs at most $M$ decoding operations.
Algorithm 1 Pseudocode of the SecCoGC Protocol
1:Input: $T,K,\eta$, $\bm{\Theta}_{k}^{0}$, $t\leftarrow 0$, $T_{s}$, $R_{t_{s}}\leftarrow 1$, $t_{s}\leftarrow 0$
2:while $\sum_{j=1}^{t_{s}}\leq T$ do
3:  $t\leftarrow t+1$
4:  Client side
5:  for client $k\in[K]$ do
6:   if $\iota=1$ then
7:     $\bm{\Theta}_{k}^{t,0}=\bm{\Theta}^{t-1}$
8:   else
9:     $\bm{\Theta}_{k}^{t+1,0}=\bm{\Theta}_{k}^{t,I}$
10:   end if
11:   for interation $i\in[I]$ do
12:     $\bm{\Theta}_{k}^{t,i}\leftarrow\bm{\Theta}_{k}^{t,i-1}-\eta a_{i}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t,i-1},\bm{\xi}_{k}^{t,i})$
13:   end for
14:   Update $\Delta\bm{\Theta}_{k}^{t,I}=\bm{\Theta}_{k}^{t,I}-\bm{\Theta}_{k}^{t,0}$
15:   Mask $\bm{Y}_{k}^{t}=\Delta\bm{\Theta}_{k}^{t,I}+\bm{N}_{k}^{t}$
16:   Send $\bm{Y}_{k}^{t}$ to neighbors in $k\in\mathcal{V}_{k}$
17:   Hear $\bm{Y}_{m}^{t}$ from neighbors in $m\in\mathcal{U}_{k}$
18:   Compute $\bm{S}_{k}^{t}=\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{Y}_{m}^{t}$
19:   Send $\bm{S}_{k}^{t}$ to the server
20:  end for
21:  Server side
22:  Detect the combinator $\bm{c}_{f_{t}}$
23:  if $\bm{c}_{f_{t}}$ exists then
24:   $\iota=1$ $\triangleright$ Indicator for GC decoding
25:   $t_{s}\leftarrow t_{s}+1$ $\triangleright$ Index for successful recovery
26:   Update $\bm{\Theta}^{t}=\bm{\Theta}^{t-1}+\frac{1}{K}\sum_{k=1}^{K}c_{f_{t},k}\bm{S}_{k}^{t}$
27:   $R_{t_{s}}\leftarrow 1$
28:  else
29:   $\iota=0$
30:   $R_{t_{s}}\leftarrow R_{t_{s}}+1$ $\triangleright$ Consecutive training rounds
31:  end if
32:  $r\leftarrow r+1$
33:end while
34:Update $T_{s}\leftarrow t_{s}$ $\triangleright$ Total number of global updates
IV Zero-Sum Noise in Real Field
This section begins by discussing the general real-field zero-sum noises. Then, it introduces fairness as a performance metric to ensure equitable privacy protection among clients.
IV-A General Construction
Define $L$ independent randomness components $\bm{Z}_{l}^{t}\sim\mathcal{N}(\mathbf{0},\bm{I}_{D})$,  [^9] The randomness $\bm{Z}_{l}^{t}$ are not necessarily zero-mean Gaussian vectors. It is assumed so for the ease of theoretical analysis. where $l\in[L]$ and $\bm{I}_{D}$ is the identity matrix of size $D\times D$. On each client $k$, the secret key is constructed as a combination of these randomness components, i.e., $\bm{N}_{k}^{t}=\sum_{l=1}^{L}\alpha_{k,l}\bm{Z}_{l}^{t}$. Let $\bm{A}$ (of size $K\times L$) denote the generator matrix comprising all $\alpha_{k,l}$. In matrix form, the secret keys on all clients are generated by
$\begin{bmatrix}\bm{N}_{1}^{t}\\ \vdots\\ \bm{N}_{K}^{t}\end{bmatrix}=\bm{A}\cdot\begin{bmatrix}\bm{Z}_{1}^{t}\\ \vdots\\ \bm{Z}_{L}^{t}\end{bmatrix}$. (14)
Below, the problem of secure aggregation is reformulated to accommodate the real field.
Condition IV.1 (Real-Field Zero-Sum Secure Aggregation).
Firstly, the secure aggregation scheme must ensure the correct computation of the global model in (6). This implies that each column of the generator matrix $\bm{A}$ must sum to zero, i.e.,
$[\mathrm{Correctness}]\;\;\;\;\;\;\forall l\in[L]:\sum_{k=1}^{K}\alpha_{k,l}=0.$ (15)
Secondly, any linear combination involving fewer than $K$ (an adjustable threshold) secret keys must be non-trivial, i.e., the zero vector must be excluded from the span of any such subset so that no combination of fewer than $K$ secret keys cancels out to zero. This requires
$[\mathrm{Security}]\;\;\;\;\;\;\;\;\;\mathrm{rank}\left(\bm{A}\right)=K-1.$ (16)
Thirdly, to avoid zero secret keys, any row in $\bm{A}$ should contain at least one non-zero entry, i.e., $\forall k$, $\exists l:\alpha_{k,l}\neq 0$.
Remark IV.1 (Necessity of the Security Condition in (16)).
Rank K1K-1 is the largest possible rank of $\bm{A}$ while ensuring correctness. It ensures that partial participation does not undermine security, since secret keys cannot be fully canceled. One may argue that the security condition in (16) is unnecessarily strong, considering the fact that GC decoding only leads to either the exact global model reconstruction with full participation or meaningless results. However, considering the client-to-client outage may lead to full rank of the coefficient matrix $\bm{G}$ and allowing for other decoding techniques [35], (16) remains essential under unreliable communication, as it offers privacy guarantees to each local model that hold regardless of the employed decoding strategy.
Lemma IV.1 (Existence Condition).
A generator matrix $\bm{A}\in\mathbb{R}^{K\times L}$ that gives distinct secret keys satisfying Condition IV.1 exists iff
$L\geq K-1.$ (17)
Proof.
The matrix $\bm{A}$ contains KLKL entries, and (15) imposes $L$ linear constraints. Thus, if $KL\geq L$, which holds for all $K$ and $L$, the system has non-zero solutions. Additionally, (16) introduces an extra constraint, requiring $L\geq K-1$. ∎
Construction IV.1 (General Construction).
Condition IV.1 suggests a straightforward construction for $\bm{A}$.
[…]
    Step 1: For the first K1K-1 rows in $\bm{A}$, each entry $\alpha_{k,l},k\in\{1,\cdots,K-1\}$ is generated randomly in $\mathbb{R}$ for $\forall l$.
[…]
    Step 2: For the last row, set $\alpha_{K,l}=-\sum_{k=1}^{K-1}\alpha_{K,l}$ for $\forall l$.
The first K1K-1 secret keys are independently generated with full freedom, while the last key is fully dependent on them. The random generation of the real-field numbers ensures (16) w.p. $1$.
IV-B Fair Zero-Sum Noise Construction
A flexible construction, e.g., Construction IV.1, often leads to variations in privacy noise power. However, signals with low or high magnitudes are susceptible to distortion or clipping by the typically nonlinear power amplifier, resulting in inaccurate and noisy aggregation. In addition, the uneven amount of randomness injected into local models implies unequal levels of privacy preservation across clients. Since the learning weights are predetermined and independent of factors such as communication or privacy, such disparities in noise power are unjustified and may raise fairness concerns.
A concrete example is given to illustrate this issue.
Example IV.1 (Fairness Issue).
Following Construction IV.1 with $K=5$ clients, the last key is generated by $\bm{N}_{5}^{t}=-\sum_{k=1}^{4}\bm{N}_{k}^{t}$, the power of $\bm{N}_{5}^{t}$ can be accumulated significantly. One example can be
$\bm{N}_{1}^{t}\sim\mathcal{N}(\mathbf{0},3.19\bm{I}_{D}),\;\bm{N}_{2}^{t}\sim\mathcal{N}(\mathbf{0},12.28\bm{I}_{D}),$
$\bm{N}_{3}^{t}\sim\mathcal{N}(\mathbf{0},2.64\bm{I}_{D}),\;\bm{N}_{4}^{t}\sim\mathcal{N}(\mathbf{0},5.80\bm{I}_{D}),$
$\bm{N}_{5}^{t}\sim\mathcal{N}(\mathbf{0},42.30\bm{I}_{D}).$
The last client receives privacy noise with power 3163-16 times higher than that of the other clients, leading to disproportionately stronger privacy protection despite all clients being assumed equally important.
The condition for fair privacy preservation requires that all privacy noises have identical power  [^10] This paper assumes equal client importance and therefore focuses on uniform privacy preservation. Secret key constructions that provide privacy levels according to heterogeneous learning weights are left for future work., i.e., $\forall k\in[K]:\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})$. This leads to the alternative condition below.
Condition IV.2 (Fairness Condition).
The sum of squared $\ell_{2}$-norm of rows in $\bm{A}$ are identical, i.e.,
$[\mathrm{Fairness}]\;\;\;\;\;\;\forall k\in[K]:\lVert\underline{\bm{\alpha}}_{k}\rVert_{2}^{2}=\lambda^{2},$ (18)
where $\underline{\bm{\alpha}}_{k}$ is the $k$-th row in $\bm{A}$ and $\lambda>0$ is a constant.
Lemma IV.2 (Existence Condition of Fair $\bm{A}$).
A generator matrix $\bm{A}\in\mathbb{R}^{K\times L}$ that gives distinct secret keys and satisfies both the fairness condition (Condition IV.2) and the secure condition (Condition IV.1) iff
$L\geq K-1.$ (19)
Proof.
The fairness condition in (18) imposes $K$ additional constraints. A non-trivial solution for $\bm{A}$ satisfying both (15) and (18) exists if $KL\geq K+L$, i.e., $K\geq 2$, $L\geq 2$. Combining this with Lemma IV.1 completes the proof. ∎
Lemma IV.3 (Lower Bound on Density).
If $\bm{A}$ is of cyclic structure, the number of nonzeros per row should satisfy
$\forall k:\;\lVert\bm{\alpha}_{k}\rVert_{0}\geq 2.$ (20)
To achieve fair zero-sum privacy noise generation, the design must simultaneously satisfy the column-wise linear constraints required for security, and the row-wise $\ell_{2}$-norm constraints required for fairness.
Construction IV.2 (General Fair Construction).
With the adjustable parameter $\lambda^{2}$ controlling the privacy noise power and $L=K$, the following construction generates fair secret keys.
[…]
    Step 1: For $\forall k\in[K]$, $\forall l\neq k,l\neq(k+1)\mod L$, $\alpha_{k,l}$ is generated randomly with full freedom.
[…]
    Step 2: The remaining $2K$ elements $\alpha_{k,l}$, where $\forall k\in[K],l=k$ and $l=(k+1)\mod L$, are fully determined by solving the following $2K$ constraints imposed by the correctness condition and fairness condition, i.e.,
    $\forall l:\sum_{k=1}^{K}\alpha_{k,l}=0,\text{and}\;\forall k:\sum_{l=1}^{L}\alpha_{k,l}^{2}=\lambda^{2}.$ (21)
The (21) can be solved by, e.g., the substitution method, the Gröbner Basis Method, etc. However, solving these equations is computationally complex in large systems. Meanwhile, the dual requirements on rows and columns pose a challenge to code construction, as it necessitates the joint design of all rows and columns, rather than treating them separately.
To overcome these limitations, we introduce a simplified generator-matrix construction with a cyclic structure that yields fair and secure secret keys without requiring solving any equations. The core idea is to first set the off-diagonal elements equal, and then determine the diagonal entry so that the $\ell_{2}$-norm across rows is enforced to be equal. In such a design, adjustments to rows do not affect their linear relationships across columns.
Construction IV.3 (Simplified Construction).
 [^11] As a comparison, the covariance-based construction in [16] is tailored to generating privacy noise with a positive-definite correlation matrix using i.i.d. Gaussian randomness. In contrast, the constructions in this paper require neither the positive-definiteness requirement nor the Gaussian assumption.
Let $L=K$ and each row contains $\gamma\geq 1$ (Lemma IV.3) off-diagonal non-zeros. The following construction of $\bm{A}$ yields fair secret keys.
[…]
    Step 1: For $\forall k\in[K]$, $\forall l=(k+l)\mod K$, $l\in\{1,\cdots,\gamma\}$, set the off-diagonal nonzero elements to
    $\alpha_{k,l}=\frac{\lambda}{\sqrt{\gamma^{2}+\gamma}}.$ (22)
[…]
    Step 2: For $\forall k\in[K]$, $\forall l=k$, the diagonal is set according to the sum of each column, i.e.,
    $\alpha_{k,k}=-\frac{\gamma\lambda}{\sqrt{\gamma^{2}+\gamma}}.$ (23)
    So far, the executed procedures have enforced equal $\ell_{2}$-norm across rows and ensured the secure condition. So for each client $k$, the privacy noise $\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{d})$.
Example IV.2.
Set $\gamma=2$, and $\lambda^{2}=6$, Construction IV.3 gives the following generator matrix,
$\bm{A}=\begin{bmatrix}-2&1&1&0&0\\ 0&-2&1&1&0\\ 0&0&-2&1&1\\ 1&0&0&-2&1\\ 1&1&0&0&-2\\ \end{bmatrix}$. (24)
In this example, $\forall k:\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},6\bm{I}_{D})$ and $\sum_{k=1}^{5}\bm{N}_{k}^{t}=\mathbf{0}$.
In the remainder of this work, the fair zero-sum noises are employed, obtained from Construction IV.3.
Lemma IV.4 (Correlation Among Privacy Noises).
From Construction IV.3, $\forall k\in[K]:\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})$. In the $d$-th dimension, where $d\in\{1,\cdots,D\}$, the covariance matrix $\mathrm{Cov}(\bm{N}_{1}^{t}(d),\cdots,\bm{N}_{K}^{t}(d))=\bm{A}\bm{A}^{\top}\triangleq\bm{R}$. Consequently, for a pair of secret keys, $\forall k,m\in[K]$,
$\mathrm{Cov}\left(\bm{N}_{k}^{t}(d),\bm{N}_{m}^{t}(d)\right)=\begin{pmatrix}\lambda^{2}&\bm{R}_{k,m}\\ \bm{R}_{m,k}&\lambda^{2}\end{pmatrix}$, (25)
where $\bm{R}_{k,m}$ is the $k,m$-th element of $\bm{R}$. The conditional distribution of $\bm{N}_{k}^{t}$ given $\bm{N}_{m}^{t}=\bm{n}_{m}$ is Gaussian: $\bm{N}_{k}^{t}\mid\bm{N}_{m}^{t}=\bm{n}_{m}\sim\mathcal{N}\!\left(\frac{\bm{n}_{m}}{\lambda^{2}}\bm{R}_{m,k},\;(\lambda^{2}-\frac{1}{\lambda^{2}}\bm{R}_{m,k}\bm{R}_{k,m})\cdot\bm{I}_{D}\right)$.
V Local Differential Privacy (LDP)
This section examines the model privacy by LDP across all levels in SecCoGC. The relevant definitions, adapted to our context, are presented in Appendix A.
V-A Peer-to-Peer LDP
The local model update is masked by (7) before transmission. Even if client $m$ can decode client $k$’s transmission, it should not confidently learn about $\Delta\bm{\Theta}_{k}^{t}$. However, unlike [16, 4], client $m$ has access to $\bm{N}_{m}^{t}=\bm{n}_{m}$ as auxiliary information when inferring the $k$-th local model, and this auxiliary information is correlated with $\bm{N}_{k}^{t}$. According to Lemma IV.4, this equivalent to adding independent privacy noise $\mathcal{N}\left(\frac{\bm{n}_{m}}{\lambda^{2}}\bm{R}_{m,k},\;(\lambda^{2}-\frac{1}{\lambda^{2}}\bm{R}_{m,k}\bm{R}_{k,m})\cdot\bm{I}_{D}\right)$. This leads us to the following Theorem.
Theorem V.1 (Peer-to-Peer LDP).
Given the secret key $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ and their covariance matrix $\bm{R}\triangleq\bm{A}\bm{A}^{\top}$ in each dimension, $\Delta\bm{\Theta}_{k}^{t},\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathbb{B}_{D}(R)$, and $\forall\delta_{m,k}^{(1)}\in(0,1]$, the masked transmission from client $k$ to client $m$ via unreliable link $\mathrm{Ber}(1-p_{m,k})$ is $(\epsilon_{m,k}^{(1)},(1-p_{m,k})\delta_{m,k}^{(1)})$-differentially private, where
$\epsilon_{m,k}^{(1)}=\frac{2R}{\sqrt{\lambda^{2}-\frac{1}{\lambda^{2}}\bm{R}_{m,k}\bm{R}_{k,m}}}\left[2\log\left(\frac{1.25}{\delta_{m,k}^{(1)}}\right)\right]^{\frac{1}{2}}.$ (26)
In other words, for any measurable set $\mathcal{S}$, it holds that
$\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t}\in\mathcal{O}_{k})\leq$ $e^{\epsilon_{m,k}^{(1)}}\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathcal{O}_{k})$
$\hskip 14.22636pt+(1-p_{m,k})\delta_{m,k}^{(1)}.$ (27)
Proof.
Proofs are provided in Appendix A. ∎
V-B LDP at the Relaying Client
After communication among clients, the relaying client $k$ computes the partial sum $\bm{S}_{k}^{t}$ in (9), based on the received masked local updates in $\mathcal{U}_{k}^{t}$. Let $\bm{S}_{k}^{t,-j}\triangleq\sum_{m\in\mathcal{U}_{k}^{t}\setminus\{j\}}g_{k,m}(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t})$ denote the partial sum excluding client $j$’s contribution. Let $\bm{M}_{k}^{t}=\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{N}_{m}^{t}$ be the aggregated privacy noises and $\nu_{k}^{t}\bm{I}_{D}=\mathrm{var}(\bm{M}_{k}^{t})$ be its variance. Similarly, define $\bm{M}_{k}^{t,-j}=\sum_{m\in\mathcal{U}_{k}^{t}\setminus\{j\}}g_{k,m}\bm{N}_{m}^{t}$, and its variance $\nu_{k}^{t,-j}\bm{I}_{D}=\mathrm{var}(\bm{M}_{k}^{t,-j})$. Define axillary matrix $\bm{\Lambda}_{k}^{t}=[\bm{\Lambda}_{k}^{t}(1),\cdots,\bm{\Lambda}_{k}^{t}(K)]$ with $\bm{\Lambda}_{k}^{t}(m)=\tau_{k,m}g_{k,m}$, and $\bm{\Lambda}_{k}^{t,-k_{1}}$ denote $\bm{\Lambda}_{k}^{t}$ with its $k_{1}$-th entry set to zero.
When an eavesdropper (possibly the server) observes the partial sum $\bm{S}_{k}^{t}$, it should not be able to identify any participating client $j$, i.e., it should not confidently infer whether client $j$’s contribution $g_{k,j}(\Delta\bm{\Theta}_{j}^{t}+\bm{N}_{j}^{t})$ is included in $\bm{S}_{k}^{t}$. In this case, privacy stems from: (i) the random participation of client $j$ at client $k$, and (ii) the aggregated privacy noise $\bm{M}_{k}^{t,-j}$. The challenges are threefold: (i) The privacy noises $\bm{N}_{m}^{t}$ are correlated, $\nu_{k}^{t}$ and $\nu_{k}^{t,-j}$ do not vary monotonically with the number of participating devices; (ii) The $\nu_{k}^{t}$ and $\nu_{k}^{t,-j}$ are random variables with respect to the network topology; (iii) The message to be protected, i.e., the $j$-th contribution $g_{k,j}(\Delta\bm{\Theta}_{j}^{t}+\bm{N}_{j}^{t})$, is unbounded. These challenges are addressed through the choice of the Bernstein parameter [4] and the Chi-squared tail bound.
Theorem V.2 (LDP in Protecting Identity in Partial Sums).
Given secret keys $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ generated by Construction IV.3, then the variance of $\bm{M}_{k}^{t}$ is given by $\nu_{k}^{t,-j}\bm{I}_{D}=\lVert\bm{\Lambda}_{k}^{t,-j}\bm{A}\rVert^{2}\bm{I}_{D}$, and $\bar{\nu}_{k}^{-j}=\mathbb{E}[\nu_{k}^{t,-j}]=\left(\lambda^{2}\sum_{m=1}^{K}(1-p_{k,m})g_{m,k}^{2}+\bm{\Lambda}_{k}^{\mathrm{E},-j}\bm{A}\bm{A}^{\top}(\bm{\Lambda}_{k}^{\mathrm{E},-j})^{\top}\right)$, where $\bm{\Lambda}_{k}^{\mathrm{E},-j}=\mathbb{E}[\bm{\Lambda}_{k}^{t,-j}]$. For $\forall\delta_{0}$, $r_{1}>0$ and $\delta^{\prime}\in(0,1)$ such that $\mathrm{P_{r}}(\lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\geq r_{1})\leq\delta^{\prime}$, [^12] The choice of Bernstein parameter $r_{1}$$r_{2}$ $r_{3}$$r_{4}$ is discussed in Appendix F. the partial sum $\bm{S}_{k}^{t}$ is $(\epsilon_{k,j}^{(2)},(1-p_{k,j})(\delta^{\prime}+\delta_{j}^{(2)}))$-differentially private w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$ in protecting the identity of any participating node $j$, where
$\epsilon_{k,j}^{(2)}=\begin{cases}\left[2\log\left(\frac{1.25}{\delta_{j}^{(2)}}\right)\right]^{\frac{1}{2}}\frac{\lvert g_{k,j}\rvert\left(R+\lambda\sqrt{D(1+\delta_{0})}\right)}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}\;\;\;\mathrm{if}\;p_{k,j}\geq 0,\\ \hskip 71.13188pt0\;\;\;\;\;\;\;\;\;\;\;\;\;\;\hskip 42.67912pt\;\;\;\mathrm{if}\;p_{k,j}=0.\end{cases}$ (28)
In other words, for any measurable set $\mathcal{S}$, it holds that
$\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\in\mathcal{S})\leq e^{\epsilon_{k,j}^{(2)}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S})+(1-p_{k,j})(\delta^{\prime}+\delta_{j}^{(2)}),$ (29)
w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$.
Proof.
Proofs are provided in [32, Appendix B]. ∎
Apart from the participating identity, when the transmission from client $j$ is perturbed from $\Delta\bm{\Theta}_{j}^{t}=\bm{u}$ to $\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime}$, the aggregated partial sums should not fluctuate significantly. Different from the previous case, the noise offering privacy protection here is $\bm{M}_{k}^{t}$, which includes the $j$-th secret key.
Theorem V.3 (LDP Against Perturbation of Local Data in Partial Sums).
Given secret keys $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ generated by Construction IV.3, then its variance is $\nu_{k}^{t}\bm{I}_{D}=\lVert\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}\bm{I}_{D}$. The $\bar{\nu}_{k}=\mathbb{E}[\nu_{k}^{t}]$ and $\bm{\Lambda}_{k}^{\mathrm{E}}=\mathbb{E}[\bm{\Lambda}_{k}^{t}]$ are specified in Theorem V.2. For $r_{2}>0$ and $\delta^{\prime}\in(0,1)$ such that $\mathrm{P_{r}}(\lvert\nu_{k}-\bar{\nu}_{k}\rvert\geq r_{2})\leq\delta^{\prime}$, the partial sum $\bm{S}_{k}^{t}$ is $(\epsilon_{k,j}^{(3)},(1-p_{k,j})(\delta^{\prime}+\delta_{j}^{(3)})$-differentially private in protecting the local date of any participating node $j$ from perturbations, where
$\epsilon_{k,j}^{(3)}=\begin{cases}2\cdot\left[2\log\left(\frac{1.25}{\delta_{j}^{(3)}}\right)\right]^{\frac{1}{2}}\frac{\lvert g_{k,j}\rvert R}{\sqrt{\bar{\nu}_{k}-r_{2}}}\;\;\;\;\;\mathrm{if}\;p_{k,j}\geq 0,\\ \hskip 42.67912pt\;\;\;\;\;0\;\;\;\;\;\;\hskip 42.67912pt\;\;\;\mathrm{if}\;p_{k,j}=0.\end{cases}$ (30)
In other words, for any $\bm{u},\bm{u^{\prime}}\in\mathbb{B}_{D}(R)$, and any measurable set $\mathcal{S}$, it holds that
$\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})$ $\leq e^{\epsilon_{k,j}^{(3)}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime})$
$+(1-p_{k,j})(\delta^{\prime}+\delta_{j}^{(3)}).$ (31)
Proof.
Proofs are provided in [32, Appendix C]. ∎
V-C LDP at the Server
Assume that the server receives partial sums from clients in $\mathcal{V}^{t}\triangleq\{k:\tau_{k}^{t}=1\}$, but fails to reconstruct the global model using the combinator $\bm{c}_{f^{\prime}}$ (referred to as a $\bm{c}_{f^{\prime}}$-Failure). The resulting aggregation result $\bm{F}_{f^{\prime}}^{t}=\sum_{k\in\mathcal{V}^{t}}c_{f^{\prime},k}\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}(\Delta\bm{\Theta}_{m}^{t,I}+\bm{N}_{m}^{t})$ still contains residual privacy noise. The effective noise in this aggregated result is given by $\bm{N}_{F_{f^{\prime}}}^{t}=\sum_{k\in\mathcal{V}^{t}}c_{f^{\prime},k}\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{N}_{m}^{t}$. The variance of $\bm{N}_{F_{f^{\prime}}}^{t}$ is $\nu_{F_{f^{\prime}}}^{t}\bm{I}_{D}=\mathrm{var}(\bm{N}_{F_{f^{\prime}}}^{t})$, and its expectation is $\bar{\nu}_{F_{f^{\prime}}}^{t}=\mathbb{E}[\mathrm{var}(\bm{N}_{F_{f^{\prime}}}^{t})]$. Similarly, define $\bm{F}_{f^{\prime}}^{t,\{-j,k\}}=\sum_{k\in\mathcal{V}^{t}}c_{f^{\prime},k}\sum_{m\in\mathcal{U}_{k}^{t}\setminus\{j\}}g_{k,m}(\Delta\bm{\Theta}_{m}^{t,I}+\bm{N}_{m}^{t})$ as the aggregation excluding client $j$’s contribution at relaying client $k$, and $\bm{N}_{F_{f^{\prime}}}^{t,\{k,-j\}}=\sum_{k\in\mathcal{V}^{t}}c_{f^{\prime},k}\sum_{m\in\mathcal{U}_{k}^{t}\setminus\{j\}}g_{k,m}\bm{N}_{m}^{t}$. Moreover, define $\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}=\mathrm{var}(\bm{N}_{F_{f^{\prime}}}^{t,\{k,-j\}})$ and $\bar{\nu}_{F_{f^{\prime}}}^{\{k,-j\}}=\mathbb{E}[\mathrm{var}(\bm{N}_{F_{f^{\prime}}}^{t,\{k,-j\}})]$ . The $\nu_{F_{f^{\prime}}}^{t}$, $\bar{\nu}_{F_{f^{\prime}}}^{t}$, $\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}$ and $\bar{\nu}_{F_{f^{\prime}}}^{\{k,-j\}}$ are further specified in Appendix G. Let $\mathcal{V}_{f^{\prime}}\triangleq\{k:c_{f^{\prime},k}\neq 0\}$, and $\mathcal{V}_{f^{\prime}}^{t}\triangleq\mathcal{V}_{f^{\prime}}\cap\mathcal{V}^{t}$.
From $\bm{F}_{f^{\prime}}^{t}$, the server should not infer the participation of client $j$ at the relaying client $k$. This privacy is offered by: (i) the privacy noise $\bm{N}_{F_{f^{\prime}}}^{t,\{k,-j\}}$, which consists of: (a) the privacy noises sent from other clients (excluding client $j$), $\mathcal{U}_{k}\setminus\{j\}$, to client $k$; and (b) the privacy noises sent from $\mathcal{U}_{m}$ to client $m\in\mathcal{V}_{f^{\prime}}\setminus\{k\}$, corresponding to non-zero combination coefficients $c_{f^{\prime},m}$; (ii) the link intermittency $\tau_{k,j}$ from client $j$ to client $k$; and (iii) the link intermittency $\tau_{k}$ from client $k$ to the server.
If $\Delta\bm{\Theta}_{j}^{t}=\bm{u}$ is perturbed to $\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime}$, the server should not observe a significant fluctuation in the resulting $\bm{F}_{f^{\prime}}^{t}$. The analytical difficulty arises as the perturbation propagates through multiple relaying nodes. The server can observe the perturbation unless every relaying path fails. The perturbation of $j$-th local data is protected by (i) the privacy noise $\bm{N}_{F_{f^{\prime}}}^{t}$, which consists of: (a) the privacy noises sent from clients (including client $j$), $\mathcal{U}_{k}$, to client $k$; and (b) the privacy noises sent from $\mathcal{U}_{m}$ to client $m\in\mathcal{V}_{f^{\prime}}\setminus\{k\}$, corresponding to non-zero combination coefficients $c_{f^{\prime},m}$; (ii) the link intermittency $\tau_{m,j}$ from client $j$ to all clients $m\in\mathcal{V}_{j}$ that relay the $j$-th message; and (iii) the link intermittency $\tau_{m}$ from client $m\in\mathcal{V}_{j}$ to the server.
Lemma V.1.
The server cannot observe the perturbation of the local data on client $j$ through $\bm{c}_{f^{\prime}}$-combination, only if all relaying clients that successfully hear from client $j$ and corresponds to a non-zero combination coefficient fails to communicate with the server, i.e., for $\forall m\in\mathcal{V}_{j,f^{\prime}}^{t}\triangleq\mathcal{V}_{j}^{t}\cap\mathcal{V}_{f^{\prime}}$, $\tau_{m}=0$, where $\mathcal{V}_{j}^{t}\triangleq\{m:m\in\mathcal{V}_{j}\}\cap\{m:\tau_{m,j}\neq 0\}$ denotes the set of neighboring clients who successfully hears from client $j$, and $\mathcal{V}_{f^{\prime}}\triangleq\{k:c_{f^{\prime},k}\neq 0\}$ denotes the clients who correspond to a non-zero combination coefficients in $\bm{c}_{f^{\prime}}$. As a result, the probability of this event is given by
$\tilde{p}_{j,f^{\prime}}=\hskip-2.84526pt\sum_{\mathcal{V}_{j,f^{\prime}}^{t}\in\{\mathcal{V}_{j}\cap\mathcal{V}_{f^{\prime}}\}}\prod_{k_{1}\in\mathcal{V}_{j,f^{\prime}}^{t}}(1-p_{k_{1}})\cdot p_{k_{1},j}\hskip-2.84526pt\hskip-8.53581pt\prod_{k_{2}\in\{\mathcal{V}_{j}\cap\mathcal{V}_{f^{\prime}}\}\setminus\mathcal{V}_{j,f^{\prime}}^{t}}\hskip-8.53581ptp_{k_{2}}.$ (32)
Theorem V.4 (LDP in $\bm{c}_{f^{\prime}}$-Failure).
Given secret keys $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ generated by Construction IV.3 and the combinator $\bm{c}_{f^{\prime}}$, then for any $r_{3}>0$ and $\delta^{\prime}\in(0,1)$ such that $\mathrm{P_{r}}(\lvert\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}-\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}\rvert\geq r_{3})\leq\delta^{\prime}$, then, in $\bm{c}_{f^{\prime}}$-Failure, $\bm{F}_{f^{\prime}}^{t}$ is $(\epsilon_{f^{\prime},\{k,j\}}^{(4)},(1-p_{k})(1-p_{k,j})(\delta_{f^{\prime},\{k,j\}}^{(4)}+\delta^{\prime})$ differentially private w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$ in protecting the identity of any participating client $j$ at relay $k$, where
$\epsilon_{f^{\prime},\{k,j\}}^{(4)}\hskip-2.84526pt=\hskip-2.84526pt\left[\hskip-1.42262pt2\log\left(\hskip-2.84526pt\frac{1.25}{\delta_{f^{\prime},\{k,j\}}^{(4)}}\hskip-2.84526pt\right)\hskip-4.2679pt\right]^{\frac{1}{2}}\hskip-5.69054pt\frac{\lvert c_{f^{\prime},k}g_{k,j}\rvert(R+\lambda\sqrt{D(1+\delta_{0})})}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}-r_{3}}}.$ (33)
In other words, for any measurable set $\mathcal{S}$, it holds that
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\in\mathcal{S})\leq$ $e^{\epsilon_{f^{\prime},\{k,j\}}^{(4)}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S})$
$\hskip-14.22636pt+(1-p_{k})(1-p_{k,j})(\delta_{f^{\prime},\{k,j\}}^{(4)}+\delta^{\prime}).$ (34)
Moreover, for any $r_{4}>0$ and $\delta^{\prime}\in(0,1)$ such that $\mathrm{P_{r}}(\lvert\nu_{F_{f^{\prime}}}^{t}-\bar{\nu}_{F_{f^{\prime}}}^{t}\rvert\geq r_{4})\leq\delta^{\prime}$, for any $\bm{u},\bm{u^{\prime}}\in\mathbb{B}_{D}(R)$, it holds that
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})$ $\leq e^{\epsilon_{k,j,f^{\prime}}^{(5)}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime})$
$+(1-\tilde{p}_{j,f^{\prime}})(\delta^{\prime}+\delta_{j}^{(5)}),$ (35)
w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$, where
$\epsilon_{k,j,f^{\prime}}^{(5)}=2\left[2\log\left(\frac{1.25}{\delta_{j}^{(5)}}\right)\right]^{\frac{1}{2}}\frac{\sum_{k=1}^{K}\lvert c_{f^{\prime},k}g_{k,j}\rvert R}{\sqrt{\bar{\nu}^{t,f^{\prime}}-r_{4}}}.$ (36)
Proof.
Proofs are provided in [32, Appendix D]. ∎
With a slight abuse of the paper organization, we examine privacy via obfuscation by other local models in successful server aggregation, considering the stochasticity of the local models. Unlike the previous LDP analysis, where noise is explicitly added to each input, privacy arises naturally from the randomness or uncertainty in the other local models. For this reason, the server cannot see significant changes in the global model $\bm{M}^{t}=\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t}$ resulting from the presence or perturbation of $\Delta\bm{\Theta}_{j}^{t}$. To facilitate theoretical analysis, it is assumed $\operatorname{Cov}[\Delta\bm{\Theta}_{k}^{t}])=\mathrm{Diag}(\bm{I}_{D}\zeta^{2})$ in this analysis.
Theorem V.5 (LDP in Global Model).
Given $\{\Delta\bm{\Theta}_{k}^{t}\}_{k=1}^{K}$ and $\forall k:\Delta\bm{\Theta}_{k}^{t}\sim\mathcal{N}(\bm{0},\zeta^{2}\bm{I}_{D})$, for any $\delta_{0}>0$ such that $\mathrm{P_{r}}\left(\left\lVert\Delta\bm{\Theta}_{j}^{t}\right\rVert\leq\zeta\sqrt{D(1+\delta_{0})}\right)\geq 1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$, the global model $\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t}$ is $(\epsilon^{(6)},\delta^{(6)})$-differentially private w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$ in protecting the identity of any node $j$, where
$\epsilon^{(6)}=\sqrt{\frac{D(1+\delta_{0})}{K-1}}\left[2\log\left(\frac{1.25}{\delta^{(6)}}\right)\right]^{\frac{1}{2}}.$ (37)
In other words, for any measurable set $\mathcal{S}$, it holds that
$\mathrm{P_{r}}(\bm{M}^{t,-j}\in\mathcal{S})\leq e^{\epsilon^{(6)}}\mathrm{P_{r}}(\bm{M}^{t}\in\mathcal{S})+\delta^{(6)},$ (38)
w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$. Moreover, for any $\bm{u},\bm{u^{\prime}}$, it holds that
$\mathrm{P_{r}}(\bm{M}^{t,-j}$ $\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})$
$\leq e^{\epsilon^{(7)}}\mathrm{P_{r}}(\bm{M}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime})+\delta^{(7)}.$ (39)
w.p. at least $\left(1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}\right)^{2}$, where
$\epsilon^{(7)}=2\sqrt{\frac{D(1+\delta_{0})}{K-1}}\left[2\log\left(\frac{1.25}{\delta^{(7)}}\right)\right]^{\frac{1}{2}}.$ (40)
Proof.
Proofs are provided in [32, Appendix E]. ∎
Remark V.1.
If $K\rightarrow\infty$, then $\epsilon^{(6)}\rightarrow 0$ and $\epsilon^{(7)}\rightarrow 0$.
VI Convergence
A non-convex convergence rate analysis of SecCoGC under an arbitrary choice of local solver is conducted. First, the assumptions concerning LOFs are present. The combination of assumptions VI.1VI.3 is the most relaxed in the non-convex FL analysis [21] and allows for broad applicability in real-world settings involving deep neural networks (DNNs).
Assumption VI.1 (Smoothness).
For $\forall i\in[K]$, each local objective function is lower bounded by $\mathcal{L}_{k}(x)\geq\mathcal{L}^{\star}$ and is Lipschitz differentiable and its gradient $\nabla\mathcal{L}_{k}(x)$ is $G$-smooth, i.e., $\lVert\nabla\mathcal{L}_{k}(x)-\nabla\mathcal{L}_{k}(y)\rVert\leq G\lVert x-y\rVert$, $\forall x,y\in\mathbb{R}^{D}$.
Assumption VI.2 (Unbiasedness and Bounded Variance).
For $\forall i\in[K]$, the local stochastic gradient is an unbiased estimator of the true local gradient, i.e., $\mathbb{E}_{\xi}[\nabla\mathcal{L}_{k}(x|\xi)]=\nabla\mathcal{L}_{k}(x)$, where $\nabla\mathcal{L}_{k}(x)$ has bounded data variance $\mathbb{E}_{\xi}[\lVert\nabla\mathcal{L}_{k}(x|\xi)-\nabla\mathcal{L}_{k}(x)\rVert^{2}]\leq\sigma^{2}$, $\sigma^{2}>0$.
Assumption VI.3 (Bounded Dissimilarity).
For any set of $\forall\{w_{k}\}_{K=1}^{M}:\sum_{K=1}^{M}w_{k}=1$, the dissimilarity between the local objective functions $\nabla\mathcal{L}_{k}(x)$ and the global objective function $\nabla\mathcal{L}(x)$ is bounded by $\sum_{k=1}^{K}w_{k}\lVert\nabla\mathcal{L}_{k}(x)\rVert^{2}\leq\beta^{2}\lVert\nabla\mathcal{L}(x)\rVert^{2}+\kappa^{2}$, $\beta^{2}\geq 1,\kappa^{2}\geq 0$.
A distinct feature of this convergence analysis is that both the number of consecutive training rounds between two adjacent successful recoveries of the global model and the total number of training rounds are random variables. Moreover, the latter is explicitly dependent on the realizations of the former. Let $t_{s}$ index the $t_{s}$-th successful global model construction, let $R_{t_{s}}$ denote the number of consecutive training rounds the FL system performs before achieving the $t_{s}$-th successful global model construction, and let $T_{s}$ be the total number of successful global updates under the stopping rule, their relationship is drawn in Fig. 2. Given the failure probability of the global model construction $P_{O}$  [^13] A detailed analysis of $P_{O}$ in SecCoGC is identical to the derivations in [35], to which the reader is referred for full detail., the following lemma characterizes their joint behavior.
Figure 2: Illustration of the number of training rounds $T$, the number of consecutive training rounds between two successful global recoveries $R_{t_{s}}$, and the total number of successful global updates $T_{s}$.
Lemma VI.1.
Define $\mathcal{C}_{1}=\left\{T_{s}>\frac{T}{e_{R}}-3\sqrt{\frac{\nu_{R}T}{e_{R}^{3}}}\right\}$, and $\mathcal{C}_{2}\triangleq\left\{\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}<e_{R^{2}}+3\sqrt{\frac{\nu_{R^{2}}}{T_{s}}}\right\}$, we have
$\mathrm{P_{r}}\left(\mathcal{C}_{1},\mathcal{C}_{2}\right)\approx 0.9987^{2}\approx 0.9974$ (41)
Proof.
The proof is provided in Appendix I. ∎
Another lemma concerning the accumulation of local updates for an arbitrary local solver is stated below.
Lemma VI.2.
With an arbitrary choice of local solver, the accumulated local update on each client after $R_{t_{s}}$ consecutive training rounds, each containing $I$ consecutive iterations, is
[…]
$\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\mathbb{E}\left[\left\lVert\Delta\bm{\Theta}_{k}^{t,R_{t_{s}}I}-\bm{\Theta}^{t-1}\right\rVert^{2}\right]$
$\leq\frac{2\eta^{2}R_{t_{s}}^{3}\lVert\bm{a}\rVert_{1}^{3}}{1-2\eta^{2}G^{2}R_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}}\mathbb{E}\left[\lVert\nabla\mathcal{L}_{k}\left(\bm{\Theta}^{t-1}\right)\rVert^{2}\right],$ (43)
where $a_{i_{s}}=a_{i_{s}\bmod I},i_{s}=1,\cdots,R_{t_{s}}I$, and $a_{i_{s}\bmod I}$ corresponds to entries in $\bm{a}$, defined in (4).
Sketch Proof.
By replacing II with $R_{t_{s}}I$ and substituting (42) into [21, Lemma 3], Lemma VI.2 is derived. ∎
Theorem VI.1.
Given an arbitrary local solver $\bm{a}$, for any algorithm that results in a binary recovery of the global model with failure probability $P_{O}$, if the learning rate is set as $\eta=\frac{1}{G}\sqrt{\frac{M}{T}}$ and the threshold of total training rounds $T$ is large enough, under Assumptions VI.1VI.3, the following convergence bound yields with probability $99.74\%$,
$\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}\mathbb{E}\left[\lVert\Delta\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]\leq\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}\mathbb{E}\left[\lVert\Delta\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]$
$\leq\frac{2G(\mathcal{L}(\bm{\Theta}^{0})-\mathcal{L}^{\star})}{\lVert\bm{a}\rVert_{1}(1-P_{O})\sqrt{KT}}+\frac{2\lVert\bm{a}\rVert_{1}(1+P_{O})}{\left(1-P_{O}\right)^{2}}\sqrt{\frac{K}{T}}(\sigma^{2}+2\kappa^{2}).$ (44)
where $T_{s}=\min\{n:\sum_{t_{s}=1}^{n}R_{t_{s}}\geq T\}$ denotes the total number of successful global update.
Proof.
Proofs are provided in [32, Appendix H]. ∎
Remark VI.1 (Role of $P_{O}$).
A large $P_{O}$ not only reduces the update frequency and slows the convergence rate by a factor of $1-P_{O}$, but also introduces an additional slowdown by $(1-P_{O})^{2}$ due to data heterogeneity and local data variance.
Remark VI.2 (Role of Privacy).
The convergence behavior of SecCoGC is independent of the privacy parameters, ensuring the optimality regardless of privacy levels.
VII Simulations
VII-A Experimental Setups
We evaluate the following methods on MNIST and CINIC-10, across varying levels of data heterogeneity, privacy preservation, and network conditions.
  1. (i)
    SecCoGC, as described in Section III.
  2. (ii)
    Standard FL with perfect connectivity. This benchmark provides insights into the ideal performance for any FL system under the same data heterogeneity.
  3. (iii)
    Non-private standard FL under unreliable communication from clients to the server.
  4. (iv)
    Private standard FL under unreliable communication, which employs the Gaussian mechanism.
  5. (v)
    Private diversity network codes (DNC)-based FL[29], which employs the Gaussian mechanism and enables full client-to-client communication among $K$ clients.
The update rule under partial participation (iii), (iv), (v) is set non-blind and follows [29, (17)]. In such a design, the effective step size is not reduced by the number of missing updates.
The MNIST [36] is an easy-level benchmark dataset with a clean and standardized format. The CINIC-10 [37], formed by both CIFAR-10 and downsampled ImageNet, is more difficult to train due to greater diversity and real-world noise. To simulate varying levels of data heterogeneity, local datasets are assigned using the Dirichlet distribution. The concentration parameter $\Gamma$ controls the level of heterogeneity among clients, where smaller values of $\Gamma$ indicate a higher imbalance.
(a) Symmetric network, $\Gamma=0.1$
(b) Asymmetric network, $\Gamma=0.1$
(c) Symmetric network, $\Gamma=0.2$
(d) Asymmetric network, $\Gamma=0.2$
Figure 3: Performance comparison under varying privacy levels, network conditions, and data heterogeneity on the MNIST dataset.
(a) Symmetric network, $\Gamma=0.1$
(b) Asymmetric network, $\Gamma=0.1$
(c) Symmetric network, $\Gamma=0.2$
(d) Asymmetric network, $\Gamma=0.2$
Figure 4: Performance comparison under varying privacy levels, network conditions, and data heterogeneity on the CINIC dataset.
In the simulation, the number of clients is set to $K=10$. Each client is assigned an equal number of data samples. The total number of training rounds is set to $T=100$, and each client performs $I=5$ local training iterations per round. The coding parameter $s$ is set to $7$ without further specification. The inter-client connectivity is configured with $p_{m,k}=0.9,\forall m\neq k$. In symmetric network settings, the local connectivity is set uniformly as $p_{k}=0.7$ for all clients. In asymmetric networks, each client is assigned a connectivity value $p_{k}\in\{0.500,0.533,\cdots,0.8\}$, for $k=\{1,\cdots,K\}$. For MNIST, the privacy level $\lambda$ is set to $0.05$ and $0.1$; for CINIC-10, it is set to $0.03$ and $0.05$. Additional hyperparameters and neural network (NN) architectures are detailed in Table I.
TABLE I: NN architecture, hyperparameter specifications
Datasets MNIST CINIC-10
Neural network CNN CNN
Model architecture
C(1,10) – C(10,20) – DL(50) – L(10)
C(3,32) – RMC(32,32) – RMDL(512) – RDL(256) – RDL(10)
Loss function Negative log-likelihood loss (NLLL) Cross-entropy loss
Learning rate $\eta$ $\eta=0.002$ $\eta=0.02$
Batch size 1024
C(# in-channel, # out-channel): a 2D convolution layer (kernel size 3, stride 1, padding 1); R: ReLU activation function; M: a 2D max-pool layer (kernel size 2, stride 2); L: (# outputs): a fully-connected linear layer; D: a dropout layer (probability 0.2).
VII-B Results and Discussions
We report the average results of multiple runs. For a fair comparison, all results are truncated at the $100^{\mathrm{th}}$ training round. The simulation results are presented in Fig. 3 and Fig. 4. These results demonstrate the strong robustness of SecCoGC against frequent communication failures between clients and the server, while preserving arbitrarily strong privacy guarantees. Detailed analyses are provided below.
VII-B1 Fast Convergence Under Privacy Preservation
When $\lambda$ is set to $0.05$ for MNIST and $0.03$ for CINIC-10, i.e., moderate levels of privacy noise, the private benchmark methods (including both private standard FL and private DNC) can work, but with significantly degraded performance. In addition to slower and less stable convergence, the final accuracy is also substantially reduced. This is because noise increases the amplitude of oscillations during optimization, and the larger step size prevents the final convergence point from accurately approaching the true global optimum. For this reason, even though DNC-based FL nearly ensures full recovery of all masked local models, masking still causes a considerable performance gap between ideal FL with perfect connectivity and DNC-based FL. As for the private standard FL, the perturbation induced by privacy-preserving noise, compounded by unreliable communication, disrupts the optimization trajectory of the masked global model. This results in a more pronounced divergence from the true global optimum, thereby leading to inferior performance compared to private DNC. With an increased privacy level ($\lambda=0.1$ for MNIST and $\lambda=0.05$ for CINIC-10), the utility degradation introduced by privacy noise becomes unacceptable. In this high-privacy regime, private standard FL consistently diverges across all configurations. Despite private DNC’s near-complete removal of communication unreliability, the global model’s accuracy remains significantly degraded, highlighting the intrinsic trade-off between strong privacy guarantees and model utility. In contrast, our proposed SecCoGC consistently approaches ideal performance across all privacy levels. This is because SecCoGC reconstructs the global model exactly whenever recovery is possible. Nonetheless, we also observe slight fluctuations of SecCoGC compared to ideal FL, which is due to the increased variance resulting from the accumulation of local model updates, implied by Lemma VI.2.
VII-B2 Straggler Mitigation Against Unreliable Communication
On both MNIST and CINIC-10, the performance drop solely due to unreliable communication becomes more pronounced with increasing data imbalance and network asymmetry, as evidenced by the widening gap between ideal FL and unprivate standard FL. In contrast, our proposed SecCoGC naturally ensures the optimality over heterogeneous networks with varying data heterogeneity, owing to its binary decoding mechanism. It is worth noting that SecCoGC effectively mitigates the impact of unreliable communication and achieves exact recovery of the global model through its structured coding scheme. This coding structure enforces the correct coefficient of each component in the aggregation, thereby ensuring that recovery is independent of any prior information. Under the same conditions, state-of-the-art methods [20, 24, 21, 22] cannot operate effectively in our setting. Furthermore, SecCoGC leverages this coefficient-enforcing property to enable secure aggregation, wherein user privacy is preserved and ambiguity is ultimately removed from the final aggregation, providing additional benefits that existing straggler-resilient schemes do not offer. Further details regarding its straggler resilience follow similarly to the analysis in [35].
VIII Conclusion
In this work, we introduce SecCoGC, a coding-enforced robust secure aggregation method. SecCoGC improves the reliability of zero-sum noise cancellation under unreliable communication, supports arbitrary privacy levels, and preserves the utility of the resulting global model. To conclude, a complete problem formulation of the real-field zero-sum noises is provided, along with the construction conditions and methods. A thorough performance analysis of the proposed SecCoGC in terms of privacy and convergence is provided. Finally, the superiority of the proposed method is validated through numerical comparisons with state-of-the-art benchmarks over multiple datasets.
Appendices
Due to the page limit, all supporting proofs are detailed in the supplementary file [32], which can be found at https://arxiv.org/abs/2507.07565.
In particular, the appendices are structured as follows.
  • Proofs of Theorem V.1 are in [32, Appendix A].
  • Proofs of Theorem V.2 are in [32, Appendix B].
  • Proofs of Theorem V.3 are in [32, Appendix C].
  • Proofs of Theorem V.4 are in [32, Appendix D].
  • Proofs of Theorem V.5 are in [32, Appendix E].
  • Choices of Bernstein Parameters $r_{1},r_{2},r_{3},r_{4}$ are discussed in [32, Appendix F and G].
  • Proofs of Theorem VI.1 are in [32, Appendix H and I].
Appendix A Proof of Theorem V.1
Definition A.1 ($(\epsilon,\delta)$-LDP).
The randomized mechanism $\mathcal{M}:\mathcal{O}\rightarrow\mathbb{R}^{D}$ is $(\epsilon,\delta)$ locally differentially private if for all dataset instances $\forall\bm{\Theta},\bm{\Theta}^{\prime}\in\mathcal{O}$, and any measurable set $\mathcal{S}\subseteq\mathrm{Range}(\mathcal{M})$, it holds that
$\mathrm{P_{r}}(\mathcal{M}(\bm{\Theta})\in\mathcal{S})\leq e^{\epsilon}\mathrm{P_{r}}(\mathcal{M}(\bm{\Theta}^{\prime})\in\mathcal{S})+\delta.$ (45)
Definition A.2 (Guassian Mechanism[3]).
If a node wishes to release a function $f(\bm{\Theta})$ of an input $\bm{\Theta}$ subject to $(\epsilon,\delta)$-LDP. The Gaussian release mechanism is defined as
$\mathcal{M}(\bm{\Theta})\triangleq f(\bm{\Theta})+\mathcal{N}(0,\varepsilon^{2}\bm{I}).$ (46)
If the $\ell_{2}$-sensitivity of the function is bounded by $\Delta_{f}$, i.e., $\lVert f(\bm{\Theta})-f(\bm{\Theta}^{\prime})\rVert\leq\Delta_{f}$, then $\forall\bm{\Theta},\bm{\Theta}^{\prime}$, $\forall\delta\in(0,1]$, the Guassian mechanism satisfies $(\epsilon,\delta)$-LDP, where
$\epsilon=\frac{\Delta_{f}}{\varepsilon}\left[2\log\left(\frac{1.25}{\delta}\right)\right]^{\frac{1}{2}}.$ (47)
Despite the standard Gaussian mechanism $\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})$ is applied to the peer-to-peer transmission from client $k$ to client $m$, the correlation between the secret key $\bm{N}_{m}^{t}$ at client $m$ and $\bm{N}_{k}^{t}$ can reveal some information about the unknown secret key. This correlation will weaken privacy to some extent. Lemma IV.4 quantitatively shows the uncertainty reduction reflected by a variance decrease from $\lambda^{2}$ to $\lambda^{2}-\frac{1}{\lambda^{2}}\bm{R}_{m,k}\bm{R}_{k,m}$. By utilizing the correlation between $\bm{N}_{m}^{t}$ and $\bm{N}_{k}^{t}$, the transmission is equivalent to applying the Gaussian mechanism $\mathcal{N}\!\left(\frac{\bm{n}_{m}}{\lambda^{2}}\bm{R}_{m,k},\;(\lambda^{2}-\frac{1}{\lambda^{2}}\bm{R}_{m,k}\bm{R}_{k,m})\cdot\bm{I}_{D}\right)$. The expected value $\frac{\bm{n}_{m}}{\lambda^{2}}\bm{R}_{m,k}$ does not affect the privacy loss, hence it is further equivalent to applying the Gaussian mechanism $\mathcal{N}\!\left(\mathbf{0},\;(\lambda^{2}-\frac{1}{\lambda^{2}}\bm{R}_{m,k}\bm{R}_{k,m})\cdot\bm{I}_{D}\right)$.
When client $m$ fails to decode message from client $k$, i.e., $\tau_{m,k}=0$, we have $\bm{Y}_{k}^{t}=0$ regardless of the values of $\Delta\bm{\Theta}_{k}^{t},\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathbb{B}_{D}(R)$, it yields perfect privacy, i.e.,
$\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t}\in\mathcal{O}_{k},\tau_{m,k}=0)$
$\hskip 56.9055pt=\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathcal{O}_{k},\tau_{m,k}=0),$ (48)
If $\tau_{m,k}=1$, client $m$ successfully decode $\bm{Y}_{k}^{t}$, the standard Gaussian mechanism applies, with the $\ell_{2}$-sensitivity given by
$\sup_{\Delta\bm{\Theta}_{k}^{t},\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathbb{B}_{D}(R)}\lVert\Delta\bm{\Theta}_{k}^{t}-\Delta\bm{\Theta}_{k}^{t^{\prime}}\rVert\leq 2R.$
Thus, it holds that
$\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t}\in\mathcal{O}_{k},\tau_{m,k}=1)$
$\leq e^{\epsilon_{m,k}^{(1)}}\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathcal{O}_{k},\tau_{m,k}=1)+\delta_{k}^{(1)},$ (49)
where
$\epsilon_{m,k}^{(1)}=\frac{2R}{\sqrt{\lambda^{2}-\frac{1}{\lambda^{2}}\bm{R}_{m,k}\bm{R}_{k,m}}}\left[2\log\left(\frac{1.25}{\delta_{m,k}^{(1)}}\right)\right]^{\frac{1}{2}}.$ (50)
Following similar steps in [4, (73)], we have
$\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t}\in\mathcal{O}_{k})$
$=p_{m,k}\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t}\in\mathcal{O}_{k},\tau_{m,k}=0)$
$\hskip 14.22636pt+(1-p_{m,k})\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t}\in\mathcal{O}_{k},\tau_{m,k}=1)$ (51a)
$\leq p_{m,k}e^{\epsilon_{m,k}^{(1)}}\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t}\in\mathcal{O}_{k},\tau_{m,k}=0)$
$\hskip 14.22636pt+(1-p_{m,k})e^{\epsilon_{m,k}^{(1)}}\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathcal{O}_{k},\tau_{m,k}=1)$
$\hskip 14.22636pt+(1-p_{m,k})\delta_{k}^{(1)}$ (51b)
$=e^{\epsilon_{m,k}^{(1)}}\mathrm{P_{r}}(\bm{Y}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathcal{O}_{k})+(1-p_{m,k})\delta_{k}^{(1)},$ (51c)
where (51b) is due to (48) and that $e^{\epsilon_{m,k}^{(1)}}\geq 1$.
Appendix B Proof of Theorem V.2
Now let us proceed to LDP at the relaying client $k$. The relaying client $k$ successfully collects the masked local model updates from clients in $\mathcal{U}_{k}^{t}$ (including itself), computes their partial sum $\bm{S}_{k}^{t}$, and transmits $\bm{S}_{k}^{t}$ to the server. During the transmission, an eavesdropper should not be able to distinguish the participation of client $j$ by observing $\bm{S}_{k}^{t}$.
If client $j\notin\mathcal{U}_{k}^{t}$, then $\bm{S}_{k}^{t,-j}=\bm{S}_{k}^{t}$. So for any measurable set $\mathcal{S}$, it holds that
$\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-2.84526pt\in\hskip-1.42262pt\mathcal{S})\hskip-1.42262pt=$ $\mathrm{P_{r}}(j\in\mathcal{U}_{k}^{t})\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t})$
$+\mathrm{P_{r}}(j\notin\mathcal{U}_{k}^{t})\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\hskip-2.84526pt\mathcal{S}|j\notin\mathcal{U}_{k}^{t}).$ (52)
Consider the network realizations reflected by $\mathcal{U}_{k}^{t}$, such that corresponding variance $\nu_{k}^{t,-j}$ lies within the event $\mathcal{E}\triangleq\{\lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r_{1}\}$, where $r_{1}$ is chosen such that $\mathrm{P_{r}}(\mathcal{E}^{\mathrm{c}})<\delta^{\prime}$, $\delta^{\prime}\in(0,1)$, then
$\hskip-1.42262pt\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t})=$ $\mathrm{P_{r}}(\mathcal{E})\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})$
$\hskip-34.1433pt+\mathrm{P_{r}}(\mathcal{E}^{\mathrm{c}})\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E}^{\mathrm{c}})$
$\hskip-42.67912pt\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E}).$ (53)
Moreover, by counting all network realizations $\mathcal{A}$ lying within the event $\mathcal{E}$, we have
$\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})=\hskip-19.91692pt\sum_{\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r_{1}\end{subarray}}\hskip-19.91692pt\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\mathrm{P_{r}}(\mathcal{A}).$ (54)
For each specific realization $\mathcal{A}$, the form of its privacy loss is given by
$\log\left(\frac{\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}=\bm{z}|i\in\mathcal{A})}{\mathrm{P_{r}}(\bm{S}_{k}^{t}=\bm{z}|i\in\mathcal{A})}\right)$
$=$ $\log\left(\frac{\exp{\left(-\frac{\lVert\bm{z}-\sum_{m\in\mathcal{A}\setminus\{j\}}(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t})\rVert^{2}}{2\nu_{k}^{t,-j}}\right)}}{\exp{\left(-\frac{\lVert\bm{z}-\sum_{m\in\mathcal{A}}(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t})\rVert^{2}}{2\nu_{k}^{t,-j}}\right)}}\right)$
$=$ $-\frac{1}{2\nu_{k}^{t,-j}}\Bigg(\left\lVert\bm{z}-\sum_{m\in\mathcal{A}\setminus\{j\}}(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t})\right\rVert^{2}$
$\hskip 85.35826pt-\left\lVert\bm{z}-\sum_{m\in\mathcal{A}}(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t})\right\rVert^{2}\Bigg).$ (55)
The form of (55) is identical to the privacy loss of the Gaussian mechanism with Gaussian noise $\nu_{k}^{t,-j}$. The $\ell_{2}$-sensitivity is bounded by
$\sup_{\hskip-8.53581pt\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r_{1}\end{subarray}}\hskip-2.84526pt\left\lVert\sum_{m\in\mathcal{A}}\hskip-2.84526ptg_{k,m}(\Delta\bm{\Theta}_{m}^{t}\hskip-2.84526pt+\hskip-2.84526pt\bm{N}_{m}^{t})\hskip-1.42262pt-\hskip-14.22636pt\sum_{m\in\mathcal{A}\setminus\{j\}}\hskip-11.38109ptg_{k,m}(\Delta\bm{\Theta}_{m}^{t}\hskip-2.84526pt+\hskip-2.84526pt\bm{N}_{m}^{t})\right\rVert$
$=\left\lVert g_{k,j}(\Delta\bm{\Theta}_{j}^{t}+\bm{N}_{j}^{t})\right\rVert\leq\left\lVert g_{k,j}\Delta\bm{\Theta}_{j}^{t}\right\rVert+\left\lVert g_{k,j}\bm{N}_{j}^{t}\right\rVert$
$\leq\lvert g_{k,j}\rvert R+\left\lVert g_{k,j}\bm{N}_{j}^{t}\right\rVert.$ (56)
Since that $g_{k,j}\bm{N}_{j}^{t}\sim\mathcal{N}(\bm{0},\lambda^{2}\bm{I}_{D})$, the second term in (56) can be bounded by Chi-squared tail bound, i.e., $\forall\delta_{0}$,
$\mathrm{P_{r}}\left(\left\lVert g_{k,j}\bm{N}_{j}^{t}\right\rVert\hskip-2.84526pt\leq\hskip-2.84526pt\lvert g_{k,j}\rvert\lambda\sqrt{D(1+\delta_{0})}\right)\hskip-1.42262pt\geq\hskip-1.42262pt1\hskip-1.42262pt-\hskip-1.42262pte^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}.$ (57)
Substituting (84) into (56), we can bound the $\ell_{2}$-sensitivity by
$\lvert g_{k,j}\rvert R+\left\lVert g_{k,j}\bm{N}_{j}^{t}\right\rVert\leq\lvert g_{k,j}\rvert\left(R+\lambda\sqrt{D(1+\delta_{0})}\right),$ (58)
with probability (w.p.) at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$.
Consequently, each specific realization $\mathcal{A}$ is $(\epsilon_{k,j}^{\mathcal{A},2},\delta_{j}^{(2)})$-differentially private w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$, i.e.,
$\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\leq e^{\epsilon_{k,j}^{\mathcal{A},2}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)},$ (59)
where $\epsilon_{k,j}^{\mathcal{A},2}=\hskip-2.84526pt\left[2\log\left(\frac{1.25}{\delta_{j}^{(2)}}\right)\right]^{\frac{1}{2}}\hskip-2.84526pt\frac{\lvert g_{k,j}\rvert\left(R+\lambda\sqrt{D(1+\delta_{0})}\right)}{\sqrt{\nu_{k}^{t,-j}}}\hskip-2.84526pt\triangleq\frac{c_{1}}{\sqrt{\nu_{k}^{t,-j}}}$.
Substitute (59) into (54), we get
$\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})$ (60a)
$\leq\hskip-14.22636pt\sum_{\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r\end{subarray}}\hskip-14.22636pt\mathrm{P_{r}}(\mathcal{A})\left(e^{\frac{c_{1}}{\sqrt{\nu_{k}^{t,-j}}}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)}\right)$ (60b)
$\leq e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\cdot\hskip-14.22636pt\sum_{\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r\end{subarray}}\hskip-14.22636pt\mathrm{P_{r}}(\mathcal{A})\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)}$ (60c)
$\leq e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\cdot\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)},$ (60d)
where (60c) holds since $\nu_{k}^{t,-j}>\bar{\nu}_{k}^{-j}-r_{1}$, (60d) is similar to (54).
Substitute (60d) into (53), we obtain
$\hskip-1.42262pt\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t})$
$\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\left(e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\cdot\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)}\right)$
$\leq\delta^{\prime}+e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\cdot\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)}.$ (61)
Furthermore, it holds that
$\hskip-1.42262pt\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\notin\mathcal{U}_{k}^{t})\leq e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\notin\mathcal{U}_{k}^{t}),$ (62)
since that $e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\geq 1$.
Substitute both (61) and (62) into (52), we get
$\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-2.84526pt\in\hskip-1.42262pt\mathcal{S})\hskip-1.42262pt\leq\hskip-1.42262pte^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-2.84526pt\in\hskip-1.42262pt\mathcal{S})\hskip-1.42262pt+\hskip-1.42262pt(1\hskip-1.42262pt-\hskip-1.42262ptp_{k,j})(\delta^{\prime}\hskip-2.84526pt+\hskip-1.42262pt\delta_{j}^{(2)}).$ (63)
The proof is completed.
Appendix C Proof of Theorem V.3
For $\forall\bm{u},\bm{u}^{\prime}\in\mathbb{B}_{D}(R)$, similar to (52), it holds that
$\mathrm{P_{r}}(\bm{S}_{k}^{t}$ $\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})$
$=$ $\mathrm{P_{r}}(j\in\mathcal{U}_{k}^{t})\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},j\in\mathcal{U}_{k}^{t})$
$+\mathrm{P_{r}}(j\notin\mathcal{U}_{k}^{t})\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}.j\notin\mathcal{U}_{k}^{t}).$ (64)
Moreover, similar to (53), it holds that
$\mathrm{P_{r}}($ $\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},j\in\mathcal{U}_{k}^{t})$
$\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},j\in\mathcal{U}_{k}^{t},\mathcal{E}),$ (65)
and similar to (59), it holds that
$\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},i\in\mathcal{A})$
$\leq e^{\epsilon_{k,j}^{\mathcal{A},3}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime},i\in\mathcal{A})+\hskip-1.42262pt\delta_{j}^{(3)}.$ (66)
The $\ell_{2}$-sensitivity is given by
$\hskip 14.22636pt\sup_{\hskip-14.22636pt\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r_{2}\end{subarray}}\hskip 0.0pt\left\lVert\bm{S}_{k}^{t,-j}+\bm{u}+\bm{N}_{j}^{t}-\bm{S}_{k}^{t,-j}-\bm{u}^{\prime}-\bm{N}_{j}^{t}\right\rVert$
$\hskip 28.45274pt\leq 2\lvert g_{k,j}\rvert R,$ (67)
which gives $\epsilon_{k,j}^{\mathcal{A},3}=\left[2\log\left(\frac{1.25}{\delta_{j}^{(3)}}\right)\right]^{\frac{1}{2}}\frac{2\lvert g_{k,j}\rvert R}{\sqrt{\nu_{k}^{t}}}\triangleq\frac{c_{2}}{\sqrt{\nu_{k}^{t}}}$.
Similar to (62), we have
$\hskip-1.42262pt\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},j\in\mathcal{U}_{k}^{t})$
$\leq\delta^{\prime}+e^{\frac{c_{2}}{\sqrt{\bar{\nu}_{k}-r_{2}}}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime},j\in\mathcal{U}_{k}^{t})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(3)}.$ (68)
By substituting (62) into (77), we obtain the final expression,
$\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})\leq$ $e^{\frac{c_{2}}{\sqrt{\bar{\nu}_{k}-r_{2}}}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime})$
$+(1-p_{k,j})(\delta^{\prime}+\delta_{j}^{(3)}),$ (69)
thereby completing the proof.
Appendix D Proof of Theorem V.4
When the combinator $\bm{c}_{f^{\prime}}$ fails to reconstruct the global model, the combination result is noisy, and the server should not learn the participation of client $j$ at the relaying client $k\in\mathcal{V}_{j}$. That is, from the aggregated result according to $f^{\prime}$-th combinator, $\bm{F}_{f^{\prime}}=\sum_{k\in\mathcal{V}^{t}}c_{f^{\prime},k}\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}(\Delta\bm{\Theta}_{m}^{t,I}+\bm{N}_{m}^{t})$ should not precisely reveal the identity of the $j$-th participating client at relay $k$. Let $\bm{F}_{f^{\prime}}^{t,\{k,-j\}}$ denote the combined results excluding the $j$-th contribution via relay $k$. It holds that
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\hskip-2.84526pt\in\mathcal{S})=$
$\mathrm{P_{r}}(\hskip-1.42262pt\{j\in\hskip-1.42262pt\mathcal{U}_{k}^{t}\}\hskip-1.42262pt\cap\hskip-1.42262pt\{k\in\hskip-1.42262pt\mathcal{V}^{t}\}\hskip-1.42262pt)\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\hskip-4.2679pt\in\hskip-1.42262pt\mathcal{S}|\{j\in\hskip-1.42262pt\mathcal{U}_{k}^{t}\}\hskip-1.42262pt\cap\hskip-1.42262pt\{k\in\hskip-1.42262pt\mathcal{V}^{t}\}\hskip-1.42262pt)$
$+\hskip-1.42262pt\mathrm{P_{r}}(\hskip-1.42262pt\{j\notin\mathcal{U}_{k}^{t}\}\hskip-1.42262pt\cup\hskip-1.42262pt\{k\notin\mathcal{V}^{t}\}\hskip-1.42262pt)\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-1.42262pt\mathcal{S}|\{j\notin\mathcal{U}_{k}^{t}\}\hskip-1.42262pt\cup\hskip-1.42262pt\{k\notin\hskip-1.42262pt\mathcal{V}^{t}\}\hskip-1.42262pt),$ (70)
where (70) follows that the scenarios where client $j$ fails to convey its local model update $\Delta\bm{\Theta}_{j}^{t}$ via relaying client $k$ either results from the event $\{j\notin\mathcal{U}_{k}^{t}\}$, where $j$ fails to share with client $m$; or $\{k\notin\mathcal{V}^{t}\}$, where $k$ fails to connect with the server.
Consider the events such that resulting variance $\nu_{k}^{t,-j}$ lies within the event $\mathcal{E}\triangleq\{\lvert\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}-\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}\rvert\leq r_{3}\}$, where $r_{3}$ is chosen such that $\mathrm{P_{r}}(\mathcal{E}^{\mathrm{c}})<\delta^{\prime}$, $\delta^{\prime}\in(0,1)$, then
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\})$
$=$ $\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\},\mathcal{E})\mathrm{P_{r}}(\mathcal{E})$
$+\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t},\mathcal{E}^{c}\})\mathrm{P_{r}}(\mathcal{E}^{c})$
$\leq$ $\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\hskip-2.84526pt\in\hskip-1.42262pt\mathcal{S}|\{j\hskip-2.84526pt\in\mathcal{U}_{k}^{t}\}\hskip-1.42262pt\cap\hskip-1.42262pt\{k\in\hskip-2.84526pt\mathcal{V}^{t}\},\mathcal{E})\mathrm{P_{r}}(\mathcal{E})+\delta^{\prime}.$ (71)
For each realization of $\mathcal{U}_{k}^{t}=\mathcal{A}$ and $\mathcal{V}^{t}=\mathcal{B}$, similar to (55), the privacy loss is identical to that of the Gaussian mechanism, with variance $\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}$. The $\ell_{2}$-sensitivity is bounded by
$\hskip 11.38109pt\sup_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \mathcal{B}\in[K],k\in\mathcal{B}\\ \lvert\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}-\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}\rvert\leq r_{3}\end{subarray}}\hskip 0.0pt\left\lVert c_{f^{\prime},k}g_{k,j}\Delta\bm{\Theta}_{j}^{t}-c_{f^{\prime},k}g_{k,j}\bm{N}_{j}^{t}\right\rVert$
$\hskip 42.67912pt\leq\lvert c_{f^{\prime},k}g_{k,j}\rvert(R+\lVert\bm{N}_{j}^{t}\rVert)$
$\hskip 42.67912pt\leq\lvert c_{f^{\prime},k}g_{k,j}\rvert(R+\lambda\sqrt{D(1+\delta_{0})}),$ (72)
with probability $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$, where Chi-Square tail bound is applied in (72). Consequently, for each realization of $\mathcal{U}_{k}^{t}=\mathcal{A}$ and $\mathcal{V}^{t}=\mathcal{B}$, $\bm{F}_{f^{\prime}}^{t,\{k,-j\}}$ is $(\epsilon_{f^{\prime},\{k,j\}}^{\mathcal{A},4},\delta_{f^{\prime},\{k,j\}}^{(4)})$-differentially private, where $\epsilon_{f^{\prime},\{k,j\}}^{\mathcal{A},4}=\hskip-2.84526pt\left[2\log\left(\frac{1.25}{\delta_{f^{\prime},\{k,j\}}^{(4)}}\right)\right]^{\frac{1}{2}}\hskip-2.84526pt\frac{\lvert c_{f^{\prime},k}g_{k,j}\rvert(R+\lambda\sqrt{D(1+\delta_{0})})}{\sqrt{\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}}}\hskip-2.84526pt\triangleq\frac{c_{3}}{\sqrt{\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}}}$. Thus, it holds that
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\},\mathcal{E})$ (73a)
$\leq\hskip-14.22636pt\sum_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \mathcal{B}\in[K],k\in\mathcal{B}\\ \lvert\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}-\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}\rvert\leq r_{3}\end{subarray}}\hskip-14.22636pt\mathrm{P_{r}}(\{j\in\mathcal{A}\}\cap\{k\in\mathcal{B}\})\Bigg(e^{\frac{c_{1}}{\sqrt{\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}}}}\cdot$
$\hskip 28.45274pt\vskip-8.53581pt\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\in\mathcal{A}\}\cap\{k\in\mathcal{B}\})\hskip-1.42262pt+\hskip-1.42262pt\delta_{f^{\prime},\{k,j\}}^{(4)}\Bigg)$ (73b)
$\leq e^{\frac{c_{3}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}-r_{3}}}}\cdot\hskip-14.22636pt\sum_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \mathcal{B}\in[K],k\in\mathcal{B}\\ \lvert\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}-\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}\rvert\leq r_{3}\end{subarray}}\hskip-14.22636pt\mathrm{P_{r}}(\{j\in\mathcal{A}\}\cap\{k\in\mathcal{B}\})\cdot$
$\hskip 28.45274pt\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\in\mathcal{A}\}\cap\{k\in\mathcal{B}\})\hskip-1.42262pt+\hskip-1.42262pt\delta_{f^{\prime},\{k,j\}}^{(4)}$ (73c)
$\leq e^{\frac{c_{3}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}-r_{3}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\},\mathcal{E})\hskip-1.42262pt$
$\hskip 170.71652pt+\hskip-1.42262pt\delta_{f^{\prime},\{k,j\}}^{(4)},$ (73d)
where (73c) holds since $\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}>\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}-r_{3}$, (73d) is similar to (54).
Furthermore, according to (71), we have
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\})$
$\leq$ $\mathrm{P_{r}}(\mathcal{E})\bigg(e^{\frac{c_{3}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}-r_{3}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\},\mathcal{E})$
$\hskip 128.0374pt+\delta_{f^{\prime},\{k,j\}}^{(4)}\bigg)+\delta^{\prime}$ (74a)
$\leq$ $e^{\frac{c_{3}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}-r_{3}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\})$
$\hskip 142.26378pt+\delta_{f^{\prime},\{k,j\}}^{(4)}+\delta^{\prime}.$ (74b)
In addition, we have
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\notin\mathcal{U}_{k}^{t}\}\cup\{k\notin\mathcal{V}^{t}\})$
$\leq e^{\frac{c_{3}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}-r_{3}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\notin\mathcal{U}_{k}^{t}\}\cup\{k\notin\mathcal{V}^{t}\}).$ (75)
Substitute (74b) and (75) into (70), we get
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\in\mathcal{S})\leq e^{\frac{c_{3}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}-r_{3}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S})$
$\hskip 28.45274pt+\mathrm{P_{r}}(\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\})\left(\delta_{f^{\prime},\{k,j\}}^{(4)}+\delta^{\prime}\right),$ (76)
where $\mathrm{P_{r}}(\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\})=(1-p_{k})(1-p_{k,j})$. The proof is thereby completed.
When $\Delta\bm{\Theta}_{j}^{t}=u$ is perturbed to $\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime}$. The $\bm{F}_{f^{\prime}}^{t}$ should not fluctuate significantly. Notably, any relaying neighbor of client $j$ (including client $j$), say client $k$, can convey the perturbation to $\bm{F}_{f^{\prime}}^{t}$ if $c_{f^{\prime}k}\neq 0$. The perturbation does not affect $\bm{F}_{f^{\prime}}^{t}$ only if all relaying neighbors of client $j$ that successfully receive from $j$ either fail to connect to the server or correspond to a zero coefficient in $\bm{c}_{f^{\prime}}$. Denote the event that the perturbation of $\Delta\bm{\Theta}_{j}^{t}$ can cause a perturbation of $\bm{F}_{f^{\prime}}^{t}$ as event $\mathcal{J}$, then $\mathcal{J}=\{\exists k:k\in\mathcal{V}_{f^{\prime}},k\in\mathcal{V}^{t},j\in\mathcal{U}_{k}^{t}\}$.
For $\forall\bm{u},\bm{u}^{\prime}\in\mathbb{B}_{D}(R)$, it holds that
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})$ $=\mathrm{P_{r}}(\mathcal{J})\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},\mathcal{J})$
$\hskip-42.67912pt+\mathrm{P_{r}}(\mathcal{J}^{c})\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime},\mathcal{J}^{c}).$ (77)
Consider the events such that resulting variance $\nu_{F_{f^{\prime}}}^{t}$ lies within the event $\mathcal{E}\triangleq\{\lvert\nu_{F_{f^{\prime}}}^{t}-\bar{\nu}_{F_{f^{\prime}}}^{t}\rvert\leq r_{4}\}$, where $r_{4}$ is chosen such that $\mathrm{P_{r}}(\mathcal{E}^{\mathrm{c}})<\delta^{\prime}$, $\delta^{\prime}\in(0,1)$, then
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u},\mathcal{J})$ $=\mathrm{P_{r}}(\mathcal{E})\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u},\mathcal{J},\mathcal{E})$
$\hskip-45.5244pt+\mathrm{P_{r}}(\mathcal{E}^{c})\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},\mathcal{J},\mathcal{E}^{c})$ (78a)
$\hskip-56.9055pt\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u},\mathcal{J},\mathcal{E})$ (78b)
For each realization of $\mathcal{U}_{k}^{t}=\mathcal{A}$ and $\mathcal{V}^{t}=\mathcal{B}$, similar to (55), the privacy loss is identical to that of the Gaussian mechanism, with variance $\nu_{F_{f^{\prime}}}^{t}$. The $\ell_{2}$-sensitivity is bounded by
$\hskip 11.38109pt\sup_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{B}\in[K],\mathcal{B}\cap\mathcal{V}_{f^{\prime}}=\emptyset,\\ k\in\mathcal{B},\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \lvert\nu_{F_{f^{\prime}}}^{t}-\bar{\nu}_{F_{f^{\prime}}}^{t}\rvert\leq r_{4}\end{subarray}}\hskip 0.0pt\left\lVert\sum_{k\in\mathcal{B}}c_{f^{\prime},k}g_{k,j}(\bm{u}-\bm{u}^{\prime})\right\rVert$
$\hskip 42.67912pt\leq\sum_{k=1}^{K}\lvert c_{f^{\prime},k}g_{k,j}\rvert\cdot 2R.$ (79)
Consequently, for each realization of $\mathcal{U}_{k}^{t}=\mathcal{A}$ and $\mathcal{V}^{t}=\mathcal{B}$, $\bm{F}_{f^{\prime}}^{t}$ is $(\epsilon_{f^{\prime},j}^{\mathcal{A},\mathcal{B},5},\delta_{f^{\prime},j}^{(5)})$-differentially private, where $\epsilon_{f^{\prime},j}^{\mathcal{A},\mathcal{B},5}=\hskip-2.84526pt\left[2\log\left(\frac{1.25}{\delta_{f^{\prime},j}^{(5)}}\right)\right]^{\frac{1}{2}}\hskip-2.84526pt\frac{2R\sum_{k=1}^{K}\lvert c_{f^{\prime},k}g_{k,j}\rvert}{\sqrt{\nu_{F_{f^{\prime}}}^{t}}}\hskip-2.84526pt\triangleq\frac{c_{4}}{\sqrt{\nu_{F_{f^{\prime}}}^{t}}}$.
Thus, it holds that
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u},\mathcal{J},\mathcal{E})$
$=\sum_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{B}\in[K],\mathcal{B}\cap\mathcal{V}_{f^{\prime}}=\emptyset,\\ k\in\mathcal{B},\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \lvert\nu_{F_{f^{\prime}}}^{t}-\bar{\nu}_{F_{f^{\prime}}}^{t}\rvert\leq r_{4}\end{subarray}}\mathrm{P_{r}}(\mathcal{A},\mathcal{B})\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u},\mathcal{J}(\mathcal{A},\mathcal{B}),\mathcal{E})$ (80a)
$\leq\hskip 0.0pt\sum_{\hskip-28.45274pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{B}\in[K],\mathcal{B}\cap\mathcal{V}_{f^{\prime}}=\emptyset,\\ k\in\mathcal{B},\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \lvert\nu_{F_{f^{\prime}}}^{t}-\bar{\nu}_{F_{f^{\prime}}}^{t}\rvert\leq r_{4}\end{subarray}}\mathrm{P_{r}}(\mathcal{A},\mathcal{B})\big(e^{\epsilon_{f^{\prime},j}^{\mathcal{A},\mathcal{B},5}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J}(\mathcal{A},\mathcal{B}),\mathcal{E})$
$\hskip 170.71652pt+\delta_{f^{\prime},j}^{(5)}\big)$ (80b)
$\leq e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t}-r_{4}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J},\mathcal{E})+\delta_{f^{\prime},j}^{(5)}.$ (80c)
Similar to (71) and (74b), it holds that
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u},\mathcal{J})$
$\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\left(e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t}-r_{4}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J},\mathcal{E})+\delta_{f^{\prime},j}^{(5)}\right)$ (81a)
$\leq e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t}-r_{4}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J})+(\delta_{f^{\prime},j}^{(5)}+\delta^{\prime})$ (81b)
Additionally, it holds that
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J}^{c})$
$\leq e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t}-r_{4}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J}^{c})$ (82)
As a result,
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u})$
$\leq e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t}-r_{4}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime})+\mathrm{P_{r}}(\mathcal{J})(\delta_{f^{\prime},j}^{(5)}+\delta^{\prime}),$ (83)
where $\mathrm{P_{r}}(\mathcal{J})=1-\tilde{p}_{j,f^{\prime}}$, and $\tilde{p}_{j,f^{\prime}}$ denotes the probability of client $j$’s contribution not being conveyed by those participating in $\bm{c}_{f^{\prime}}$-combination.
Appendix E Proof of Theorem V.5
Assume $\operatorname{Cov}[\Delta\bm{\Theta}_{k}^{t}])=\mathrm{Diag}(\bm{I}_{D}\zeta^{2})$, now we investigate the privacy protection for the identity of the participating client $j$ provided by other local model updates, in the successfully reconstructed global model $\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t}$. According to the Chi-squared tail bound, $\forall\delta_{0}$, it holds that
$\mathrm{P_{r}}\left(\left\lVert\Delta\bm{\Theta}_{j}^{t}\right\rVert\hskip-1.42262pt\leq\hskip-1.42262pt\zeta\sqrt{D(1+\delta_{0})}\right)\hskip-1.42262pt\geq\hskip-1.42262pt1\hskip-1.42262pt-\hskip-1.42262pte^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}.$ (84)
Thus, the $\ell_{2}$-sensitivity can be bounded by
$\left\lVert\frac{1}{K}\Delta\bm{\Theta}_{j}^{t}\right\rVert\leq\frac{\zeta}{K}\sqrt{D(1+\delta_{0})}$ (85)
with probability $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$. The effective noise $\frac{1}{K}\sum_{k\neq j}\Delta\bm{\Theta}_{k}^{t}\sim\mathcal{N}(0,\frac{K-1}{K}\zeta^{2}\bm{I}_{D})$, therefore, the conclusion of Gaussian mechanism applies. Specifically, $\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t}$ is $(\epsilon^{(6)},\delta^{(6)})$-differentially private in protecting the identity of client $j$ with probability $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$, where
$\epsilon^{(6)}=\frac{\frac{\zeta}{K}\sqrt{D(1+\delta_{0})}}{\frac{\sqrt{K-1}}{K}\zeta}\left[2\log\left(\frac{1.25}{\delta^{(6)}}\right)\right]^{\frac{1}{2}},$
i.e.,
$\epsilon^{(6)}=\sqrt{\frac{D(1+\delta_{0})}{K-1}}\left[2\log\left(\frac{1.25}{\delta^{(6)}}\right)\right]^{\frac{1}{2}}.$
The (84) holds before and after perturbation. Hence, $\mathrm{P_{r}}\left(\bm{u}\in\mathbb{B}_{D}(\zeta\sqrt{D(1+\delta_{0})})\right)\geq 1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$ and $\mathrm{P_{r}}\left(\bm{u^{\prime}}\in\mathbb{B}_{D}(\zeta\sqrt{D(1+\delta_{0})})\right)\geq 1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$. When $\Delta\bm{\Theta}_{j}^{t}$ is perturbed from $\bm{u}$ to $\bm{u}^{\prime}$, the $\ell_{2}$-sensitivity can be bounded by
$\left\lVert\frac{1}{K}(\bm{u}-\bm{u}^{\prime})\right\rVert\leq\frac{2\zeta}{K}\sqrt{D(1+\delta_{0})}.$ (86)
with probability $\left(1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}\right)^{2}$. So $\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t}$ is $(\epsilon^{(7)},\delta^{(7)})$-differentially private in protecting the local data of client $j$ with probability $\left(1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}\right)^{2}$, where
$\epsilon^{(7)}=2\sqrt{\frac{D(1+\delta_{0})}{K-1}}\left[2\log\left(\frac{1.25}{\delta^{(7)}}\right)\right]^{\frac{1}{2}}.$ (87)
Appendix F Choices of Bernstein Parameters
Lemma F.1 (Bernstein’s Inequality).
Let $X$ be a zero-mean real-valued random variable, bounded by $M$, almost surely. Then, for $\forall r\geq 0$, we have
$\mathrm{P_{r}}(X\geq r)\leq\exp\left(-\frac{r^{2}}{\mathbb{E}[X^{2}]+\frac{2}{3}Mr}\right),$ (88)
or
$\mathrm{P_{r}}(\lvert X\rvert\geq r)\leq 2\exp\left(-\frac{r^{2}}{\mathbb{E}[X^{2}]+\frac{2}{3}Mr}\right).$ (89)
By choosing $\delta^{\prime}=2\exp\left(-\frac{r^{2}}{\mathbb{E}[X^{2}]+\frac{2}{3}Mr}\right)$, this is a quadratic in $r$ and $r$ can be solved accordingly, i.e.,
$r=\ln\left(\frac{2}{\delta^{\prime}}\right)\cdot\left(\frac{1}{3}M+\sqrt{\frac{1}{9}M^{2}+\frac{\mathbb{E}[X^{2}]}{\ln\left(\frac{2}{\delta^{\prime}}\right)}}\right).$ (90)
F-A Choice of $r_{1}$ and $r_{2}$
Let $X_{k,2}=\nu_{k}^{t}-\bar{\nu}_{k}$. Then, $\mathbb{E}[X_{k,2}]=0$, and $\max\{X_{k,2}\}\leq\max\{\nu_{k}^{t}\}\leq\max_{k}\{\nu_{k}^{t}\}\triangleq M_{2}$. Due to Bernoulli r.v.s, $M_{2}$ is computable, through e.g., brute-force search. Moreover,
$\mathbb{E}[X_{k,2}^{2}]=\mathbb{E}[(\nu_{k}^{t}-\bar{\nu}_{k})^{2}]=\mathbb{E}[(\nu_{k}^{t})^{2}]-\bar{\nu}_{k}^{2},$ (91)
where $\bar{\nu}_{k}$ and $\mathbb{E}[(\nu_{k}^{t})^{2}]$ are given in (100) and (103b), respectively. Thereby, $r_{2}$ is chosen as
$r_{2}=\ln\left(\frac{2}{\delta^{\prime}}\right)\cdot\left(\frac{1}{3}M_{2}+\sqrt{\frac{1}{9}M_{2}^{2}+\frac{\mathbb{E}[X_{k,2}^{2}]}{\ln\left(\frac{2}{\delta^{\prime}}\right)}}\right).$ (92)
Similarly, let $X_{k,1}=\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}$. Then, $\mathbb{E}[X_{k,1}]=0$, and $\max\{X_{k,1}\}\leq\max\{\nu_{k}^{t,-j}\}\leq\max_{k}\{\nu_{k}^{t,-j}\}\triangleq M_{1}$. By setting $\tau_{k,j}=0$, $M_{1}$ is computable, through e.g., brute-force search. Moreover,
$\mathbb{E}[X_{k,1}^{2}]=\mathbb{E}[(\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j})^{2}]=\mathbb{E}[(\nu_{k}^{t,-j})^{2}]-(\bar{\nu}_{k}^{-j})^{2},$ (93)
where $(\bar{\nu}_{k}^{-j})^{2}$ is given in (102) and $\mathbb{E}[(\nu_{k}^{t,-j})^{2}]$ can be computed by setting $\tau_{k,j}=0$ in (103b). Thereby, $r_{1}$ is chosen as
$r_{1}=\ln\left(\frac{2}{\delta^{\prime}}\right)\cdot\left(\frac{1}{3}M_{1}+\sqrt{\frac{1}{9}M_{1}^{2}+\frac{\mathbb{E}[X_{k,1}^{2}]}{\ln\left(\frac{2}{\delta^{\prime}}\right)}}\right).$ (94)
F-B Choice of $r_{3}$ and $r_{4}$
Let $X_{k,4}=\nu_{F_{f^{\prime}}}^{t}-\bar{\nu}_{F_{f^{\prime}}}$. Then, $\mathbb{E}[X_{k,4}]=0$, and $\max\{X_{k,4}\}\leq\max\{\nu_{F_{f^{\prime}}}^{t}\}\leq\max_{f^{\prime}}\{\nu_{F_{f^{\prime}}}^{t}\}\triangleq M_{4}$. Due to the Bernoulli r.v.s, $M_{4}$ is computable, through e.g., brute-force search. Moreover,
$\mathbb{E}[X_{k,4}^{2}]=\mathbb{E}[(\nu_{F_{f^{\prime}}}^{t}-\bar{\nu}_{F_{f^{\prime}}})^{2}]=\mathbb{E}[(\nu_{F_{f^{\prime}}}^{t})^{2}]-\bar{\nu}_{F_{f^{\prime}}}^{2},$ (95)
where $\bar{\nu}_{F_{f^{\prime}}}$ and $\mathbb{E}[(\nu_{F_{f^{\prime}}}^{t})^{2}]$ are given in (107) and (103b), respectively. Thereby, $r_{4}$ is chosen as
$r_{4}=\ln\left(\frac{2}{\delta^{\prime}}\right)\cdot\left(\frac{1}{3}M_{4}+\sqrt{\frac{1}{9}M_{4}^{2}+\frac{\mathbb{E}[X_{k,4}^{2}]}{\ln\left(\frac{2}{\delta^{\prime}}\right)}}\right).$ (96)
Similarly, let $X_{k,3}=\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}-\bar{\nu}_{F_{f^{\prime}}}^{\{k,-j\}}$. Then, $\mathbb{E}[X_{k,3}]=0$, and $\max\{X_{k,3}\}\leq\max\{\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}\}\leq\max_{f^{\prime}}\{\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}\}\triangleq M_{3}$. By setting $\tau_{k,j}=0$, $M_{3}$ is computable, through e.g., brute-force search. Moreover,
$\mathbb{E}[X_{k,3}^{2}]$ $=\mathbb{E}[(\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}-\bar{\nu}_{F_{f^{\prime}}}^{\{k,-j\}})^{2}]$
$=\mathbb{E}[(\nu_{F_{f^{\prime}}}^{t,\{k,-j\}})^{2}]-(\bar{\nu}_{F_{f^{\prime}}}^{\{k,-j\}})^{2},$ (97)
where $\bar{\nu}_{F_{f^{\prime}}}^{\{k,-j\}}$ and $\mathbb{E}[(\nu_{F_{f^{\prime}}}^{t,\{k,-j\}})^{2}]$ can be computed be setting $\tau_{k,j}=0$ in (107) and (112), respectively. Thereby, $r_{3}$ is chosen as
$r_{3}=\ln\left(\frac{2}{\delta^{\prime}}\right)\cdot\left(\frac{1}{3}M_{3}+\sqrt{\frac{1}{9}M_{3}^{2}+\frac{\mathbb{E}[X_{k,3}^{2}]}{\ln\left(\frac{2}{\delta^{\prime}}\right)}}\right).$ (98)
Appendix G Calculation of Privacy Noise Variance
G-A Calculation of $\bar{\nu}_{k}$ and $\mathbb{E}[(\nu_{k}^{t})^{2}]$
Let us begin by computing the variance of the aggregated secret keys $\bm{M}_{k}^{t}$ at relaying client $k$. It can be verified that $\bm{M}_{k}^{t}=\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{N}_{m}^{t}=\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{k}^{t}$, so $\mathrm{var}(\bm{M}_{k}^{t})=\nu_{k}^{t}\bm{I}_{D}=\lVert\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}\bm{I}_{D}$ in (LABEL:eq:_agg_noise). Similarly, we define $\bm{M}_{k}^{t,-j}=\sum_{m\in\mathcal{U}_{k}^{t}\setminus\{j\}}g_{k,m}\bm{N}_{m}^{t}=\sum_{m\neq j}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{k}^{t}$, and we have $\mathrm{var}(\bm{M}_{k}^{t,-j})=\nu_{k}^{t,-j}\bm{I}_{D}=\lVert\bm{\Lambda}_{k}^{t,-j}\bm{A}\rVert^{2}\bm{I}_{D}$ in (LABEL:eq:_agg_noise_no_j). The $\bm{\Lambda}_{k}^{t}$ depends on $\{\tau_{k,m}\}_{m=1}^{K}$, so the variance $\nu_{k}^{t}$ of the effective noise $\bm{M}_{k}^{t}$ is also a random variable concerning the network topology. It should be noted that due to correlation, the variance of $\bm{M}_{k}^{t}$ does not necessarily increase monotonously with the number of received updates $|\mathcal{U}_{k}^{t}\rvert$ at client $k$. The expected variance $\bar{\nu}_{k}$ is given by
$\bar{\nu}_{k}=\mathbb{E}_{\mathcal{U}_{k}^{t}}[\nu_{k}^{t}]=\mathbb{E}\left[\lVert\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}\right].$ (99)
Furthermore, we have
$\mathbb{E}\left[\lVert\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}\right]=\mathbb{E}\left[\bm{\Lambda}_{k}^{t}\bm{A}\bm{A}^{\top}(\bm{\Lambda}_{k}^{t})^{\top}\right]$
$=\operatorname{Tr}\left(\operatorname{Cov}[\bm{\Lambda}_{k}^{t}]\bm{A}\bm{A}^{\top}\right)+\mathbb{E}\left[\bm{\Lambda}_{k}^{t}\bm{A}\bm{A}^{\top}(\bm{\Lambda}_{k}^{t})^{\top}\right]$
$=\lambda^{2}\sum_{m=1}^{K}(1-p_{k,m})g_{m,k}^{2}+\bm{\Lambda}_{k}^{\mathrm{E}}\bm{A}\bm{A}^{\top}(\bm{\Lambda}_{k}^{\mathrm{E}})^{\top},$ (100)
where $\bm{\Lambda}_{k}^{\mathrm{E}}=\mathbb{E}[\bm{\Lambda}_{k}^{t}]=[(1-p_{k,1})g_{k,1},\cdots,(1-p_{k,K})g_{k,K}]$.
Similarly, we also have
$\bar{\nu}_{k}^{-j}=\mathbb{E}_{\mathcal{U}_{k}^{t}}[\nu_{k}^{t,-j}]=\mathbb{E}\left[\lVert\bm{\Lambda}_{k}^{t,-j}\bm{A}\rVert^{2}\right]\bm{I}_{D},$ (101)
and
$\mathbb{E}\left[\lVert\bm{\Lambda}_{k}^{t,-j}\bm{A}\rVert^{2}\right]$
$=\lambda^{2}\sum_{m=1}^{K}(1-p_{k,m})g_{m,k}^{2}+\bm{\Lambda}_{k}^{\mathrm{E},-j}\bm{A}\bm{A}^{\top}(\bm{\Lambda}_{k}^{\mathrm{E},-j})^{\top},$ (102)
where $\bm{\Lambda}_{k}^{\mathrm{E},-j}$ is identical to $\bm{\Lambda}_{k}^{\mathrm{E}}$ except the $j$-th element is set to 0.
It can be verified that $\nu_{k}^{t}$ can also be written in the form $\nu_{k}^{t}=\sum_{k_{1},k_{2}\in\mathcal{U}_{k}}C^{(k)}_{k_{1},k_{2}}\tau^{t}_{k,k_{1}}\tau^{t}_{k,k_{2}}$, where $C^{(k)}_{k_{1},k_{2}}=g_{k,k_{1}}[\bm{A}\bm{A}^{\top}]_{k_{1}k_{1}}g_{k,k_{2}}$. Following this, we have
$\left(\nu_{k}^{t}\right)^{2}$ $=\hskip-5.69054pt\sum_{k_{1},k_{2},k_{3},k_{4}\in\mathcal{U}_{k}}\hskip-5.69054ptC^{(k)}_{k_{1},k_{2}}C^{(k)}_{k_{3},k_{4}}\tau^{t}_{k,k_{1}}\tau^{t}_{k,k_{2}}\tau^{t}_{k,k_{3}}\tau^{t}_{k,k_{4}},$ (103a)
$\mathbb{E}\left[\left(\nu_{k}^{t}\right)^{2}\right]$ $=\hskip-5.69054pt\sum_{k_{1},k_{2},k_{3},k_{4}\in\mathcal{U}_{k}}\hskip-5.69054ptC^{(k)}_{k_{1},k_{2}}C^{(k)}_{k_{3},k_{4}}\mathbb{E}\left[\tau^{t}_{k,k_{1}}\tau^{t}_{k,k_{2}}\tau^{t}_{k,k_{3}}\tau^{t}_{k,k_{4}}\right],$ (103b)
where $\mathbb{E}\left[\tau^{t}_{k,k_{1}}\tau^{t}_{k,k_{2}}\tau^{t}_{k,k_{3}}\tau^{t}_{k,k_{4}}\right]$ is given by
$\mathbb{E}\left[\tau^{t}_{k,k_{1}}\tau^{t}_{k,k_{2}}\tau^{t}_{k,k_{3}}\tau^{t}_{k,k_{4}}\right]=\prod_{u\in\{k_{1}\}\cup\{k_{2}\}\cup\{k_{3}\}\cup\{k_{4}\}}(1-p_{k,u}),$ (104)
since $\tau^{t}_{k,k_{1}}$ and $\tau^{t}_{k,k_{2}}$ are independent Bernoulli r.v., $(\tau^{t}_{k,k_{1}})^{2}=\tau^{t}_{k,k_{1}}$, and $\{k_{1}\}\cup\{k_{2}\}\cup\{k_{3}\}\cup\{k_{4}\}$ ensures the uniqueness of the links in the product.
G-B Calculation of $\bar{\nu}_{F_{f^{\prime}}}$ and $\mathbb{E}[(\nu_{F_{f^{\prime}}}^{t})^{2}]$
The aggregated secret key according to $\bm{c}_{f^{\prime}}$ is
$\bm{N}_{F_{f^{\prime}}}^{t}$ $=\sum_{k\in\mathcal{V}^{t}}c_{f^{\prime},k}\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{N}_{m}^{t}$ (105a)
$=\sum_{k=1}^{K}c_{f^{\prime},k}\tau_{k}^{t}\sum_{m=1}^{K}\tau_{k,m}^{t}g_{k,m}\bm{N}_{m}^{t}$ (105b)
$=\sum_{k=1}^{K}c_{f^{\prime},k}\tau_{k}^{t}\sum_{m=1}^{K}\tau_{k,m}^{t}g_{k,m}\sum_{l=1}^{L}\alpha_{m,l}\bm{Z}_{l}^{t}$ (105c)
$=\sum_{l=1}^{L}\sum_{k=1}^{K}c_{f^{\prime},k}\tau_{k}^{t}\sum_{m=1}^{K}\tau_{k,m}^{t}g_{k,m}\alpha_{m,l}\bm{Z}_{l}^{t}.$ (105d)
Thus, $\mathrm{var}(\bm{N}_{F_{f^{\prime}}}^{t})=\nu_{F_{f^{\prime}}}^{t}\bm{I}_{D}$, where
$\nu_{F_{f^{\prime}}}^{t}$ $=\sum_{l=1}^{L}\left(\sum_{k=1}^{K}c_{f^{\prime},k}\tau_{k}^{t}\sum_{m=1}^{K}\tau_{k,m}^{t}g_{k,m}\alpha_{m,l}\right)^{2}$ (106a)
$=\sum_{l=1}^{L}\left(\sum_{k=1}^{K}c_{f^{\prime},k}\tau_{k}^{t}\bm{\Lambda}_{k}^{t}\bm{\alpha}_{l}\right)^{2}$ (106b)
$\hskip-8.53581pt=\sum_{l=1}^{L}\bm{\alpha}_{l}^{\top}\left(\sum_{k_{1},k_{2}}c_{f^{\prime},k_{1}}c_{f^{\prime},k_{2}}\tau_{k_{1}}^{t}\tau_{k_{2}}^{t}\bm{\Lambda}_{k_{1}}^{t}\bm{\Lambda}_{k_{2}}^{t}\right)\bm{\alpha}_{l}.$ (106c)
Thus, we have
$\bar{\nu}_{F_{f^{\prime}}}=\mathbb{E}\left[\nu_{F_{f^{\prime}}}^{t}\right]=\sum_{l=1}^{L}\bm{\alpha}_{l}^{\top}\left(\sum_{k_{1},k_{2}}c_{f^{\prime},k_{1}}c_{f^{\prime},k_{2}}\tau_{k_{1}}^{t}\tau_{k_{2}}^{t}\Sigma_{k}\right)\bm{\alpha}_{l},$ (107)
where $\Sigma_{k}=\mathbb{E}\left[\bm{\Lambda}_{k_{1}}^{t}\bm{\Lambda}_{k_{2}}^{t}\right]$ and is given by
$[\Sigma_{k}]_{ij}=\begin{cases}g_{k,i}^{2}(1-p_{k,i}),\;\;\;\;\;\text{if}\;i=j\\ \left[(\bm{\Lambda}_{k}^{\mathrm{E}})^{\top}\bm{\Lambda}_{k}^{\mathrm{E}}\right]_{ij}\;\;\;\;\;\text{if}\;i\neq j\end{cases}.$ (108)
Next, let us calculate $\mathbb{E}\left[(\nu_{F_{f^{\prime}}}^{t})^{2}\right]$. The (106a) can be rewritten as
$\nu_{F_{f^{\prime}}}^{t}=\sum_{l=1}^{L}\left(\sum_{k_{1},m_{1}}c_{f^{\prime},k_{1}}g_{k_{1},m_{1}}\tau_{k_{1}}^{t}\tau_{k_{1},m_{1}}^{t}\alpha_{m_{1},l}\right)^{2}$ (109a)
$=\sum_{l=1}^{L}\sum_{k_{1},m_{1}}\sum_{k_{2},m_{2}}c_{f^{\prime},k_{1}}c_{f^{\prime},k_{2}}g_{k_{1},m_{1}}g_{k_{2},m_{2}}\alpha_{m_{1},l}\alpha_{m_{2},l}$
$\hskip 113.81102pt\cdot\tau_{k_{1}}^{t}\tau_{k_{1},m_{1}}^{t}\tau_{k_{2}}^{t}\tau_{k_{2},m_{2}}^{t}.$ (109b)
Let $A_{k_{1},m_{1}}^{f^{\prime}}=c_{f^{\prime},k_{1}}g_{k_{1},m_{1}}$, $t(k_{1},m_{1})=\tau_{k_{1}}^{t}\tau_{k_{1},m_{1}}^{t}$, and $H_{m_{1},m_{2}}=\sum_{l=1}^{L}\alpha_{m_{1},l}\alpha_{m_{2},l}$ (109b) can be simplified to
$\nu_{F_{f^{\prime}}}^{t}=\sum_{k_{1},m_{1}}\sum_{k_{2},m_{2}}A_{k_{1},m_{1}}^{f^{\prime}}A_{k_{2},m_{2}}^{f^{\prime}}H_{m_{1},m_{2}}t(k_{1},m_{1})t(k_{2},m_{2}).$ (110)
Thus,
$(\nu_{F_{f^{\prime}}}^{t})^{2}=\hskip-5.69054pt\sum_{k_{1},m_{1}}\sum_{k_{2},m_{2}}\sum_{k_{3},m_{3}}\sum_{k_{4},m_{4}}A_{k_{1},m_{1}}^{f^{\prime}}A_{k_{2},m_{2}}^{f^{\prime}}A_{k_{3},m_{3}}^{f^{\prime}}A_{k_{4},m_{4}}^{f^{\prime}}$
$\hskip 14.22636pt\cdot H_{m_{1},m_{2}}H_{m_{3},m_{4}}t(k_{1},m_{1})t(k_{2},m_{2})t(k_{3},m_{3})t(k_{4},m_{4}).$ (111)
Thus,
$\mathbb{E}\left[(\nu_{F_{f^{\prime}}}^{t})^{2}\right]\hskip-2.84526pt=\hskip-5.69054pt\sum_{k_{1},m_{1}}\hskip-1.42262pt\sum_{k_{2},m_{2}}\hskip-1.42262pt\sum_{k_{3},m_{3}}\hskip-1.42262pt\sum_{k_{4},m_{4}}A_{k_{1},m_{1}}^{f^{\prime}}A_{k_{2},m_{2}}^{f^{\prime}}A_{k_{3},m_{3}}^{f^{\prime}}A_{k_{4},m_{4}}^{f^{\prime}}$
$\hskip 5.69046pt\cdot H_{m_{1},m_{2}}H_{m_{3},m_{4}}\mathbb{E}\left[t(k_{1},m_{1})t(k_{2},m_{2})t(k_{3},m_{3})t(k_{4},m_{4})\right],$ (112)
where
$\mathbb{E}\left[t(k_{1},m_{1})t(k_{2},m_{2})t(k_{3},m_{3})t(k_{4},m_{4})\right]$
$=\prod_{\begin{subarray}{c}u\in\{k_{1}\}\cup\{k_{2}\}\\ \cup\{k_{3}\}\cup\{k_{4}\}\end{subarray}}(1-p_{u})\prod_{\begin{subarray}{c}(u,v)\in\{(k_{1},m_{1})\}\cup\{(k_{2},m_{2})\}\\ \cup\{(k_{3},m_{3})\}\cup\{(k_{4},m_{4})\}\end{subarray}}(1-p_{u,v}),$ (113)
since $\mathbb{E}\left[t(k_{1},m_{1})\right]=(1-p_{k_{1}})(1-p_{k_{1},m_{1}})$, $t(k_{1},m_{1})=t(k_{2},m_{2})$ for $k_{1}=k_{2}$, $m_{1}=m_{2}$, and $\cup$ ensures the uniqueness of the set elements.
Appendix H Complete Proof of Theorem VI.1
Under Assumption VI.1, it holds that
$\mathbb{E}\left[\mathcal{L}(\bm{\Theta}^{t_{s}})\right]-\mathbb{E}\left[\mathcal{L}(\bm{\Theta}^{t_{s}-1})\right]$
$\leq\mathbb{E}\hskip-1.42262pt\left[\left\langle\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1}),\bm{\Theta}^{t_{s}}\hskip-1.42262pt-\hskip-1.42262pt\bm{\Theta}^{t_{s}-1}\right\rangle\right]\hskip-2.84526pt+\hskip-2.84526pt\frac{G}{2}\mathbb{E}\hskip-1.42262pt\left[\lVert\bm{\Theta}^{t_{s}}\hskip-1.42262pt-\hskip-1.42262pt\bm{\Theta}^{t_{s}-1}\rVert^{2}\right].$ (114)
Next, we bound the two terms in (114), respectively.
H-A Bounding the First Term
For the first term, we have
$\mathbb{E}\left[\left\langle\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1}),\bm{\Theta}^{t_{s}}-\bm{\Theta}^{t_{s}-1}\right\rangle\right]$
$=\mathbb{E}\hskip-2.84526pt\left[\hskip-2.84526pt\left\langle\hskip-2.84526pt\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1}),-\eta\frac{1}{K}\sum_{k=1}^{K}\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1},\bm{\xi}_{k}^{t_{s},i_{s}})\hskip-2.84526pt\right\rangle\hskip-2.84526pt\right]$
$=-\eta\hskip-2.84526pt\sum_{i_{s}=1}^{R_{t_{s}}I}\hskip-2.84526pta_{i_{s}}\mathbb{E}\hskip-2.84526pt\left[\hskip-2.84526pt\left\langle\hskip-2.84526pt\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1}),\frac{1}{K}\hskip-2.84526pt\sum_{k=1}^{K}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1})\hskip-2.84526pt\right\rangle\hskip-2.84526pt\right]$ (115a)
$\leq-\frac{1}{2}\eta\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\mathbb{E}\left[\left\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\right\lVert^{2}\right]$
$\hskip 14.22636pt+\frac{1}{2}\eta\hskip-2.84526pt\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\hskip-2.84526pt\left[\left\lVert\nabla\mathcal{L}(\bm{\Theta}_{k}^{t_{s}-1})\hskip-2.84526pt-\hskip-2.84526pt\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s},i_{s}-1})\right\lVert^{2}\right]$
$\hskip 14.22636pt-\frac{1}{2}\eta\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\mathbb{E}\left[\left\lVert\frac{1}{K}\sum_{k=1}^{K}\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s},i_{s}-1})\right\lVert^{2}\right]$ (115b)
$\leq-\frac{1}{2}\eta\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\mathbb{E}\left[\left\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\right\lVert^{2}\right]$
$\hskip 14.22636pt+\frac{1}{2}\eta\frac{G^{2}}{K}\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\sum_{k=1}^{K}\mathbb{E}\left[\left\lVert\bm{\Theta}_{k}^{t_{s}-1}-\bm{\Theta}^{t_{s},i_{s}-1}\right\lVert^{2}\right],$ (115c)
where (115a) follows Assumption VI.2, (115b) is due to the convexity of $\ell_{2}$-norm, and (115c) is due to Assumption VI.1.
H-B Bounding the Second Term
For the second term, it holds that
$\frac{G}{2}\mathbb{E}\left[\lVert\bm{\Theta}^{t_{s}}-\bm{\Theta}^{t_{s}-1}\rVert^{2}\right]$
$=\frac{G}{2}\mathbb{E}\left[\left\lVert-\eta\frac{1}{K}\sum_{k=1}^{K}\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1},\bm{\xi}_{k}^{t_{s},i_{s}})\right\rVert^{2}\right]$
$\leq\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\left[\left\lVert\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1},\bm{\xi}_{k}^{t_{s},i_{s}})\right\rVert^{2}\right]$ (116a)
$\leq\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\mathbb{E}\hskip-1.42262pt\left[\left\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1},\bm{\xi}_{k}^{t_{s},i_{s}})\right\rVert^{2}\right]$ (116b)
$=\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\Bigg(\mathbb{E}\bigg[\bigg\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1},\bm{\xi}_{k}^{t_{s},i_{s}})$
$\hskip 14.22636pt-\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1})\bigg\rVert^{2}\bigg]+\mathbb{E}\left[\left\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1})\right\rVert^{2}\right]\Bigg)$ (116c)
$\leq\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\Bigg(\sigma^{2}+\mathbb{E}\bigg[\bigg\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1})$
$\hskip 14.22636pt-\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s}-1})\bigg\rVert^{2}\bigg]+\mathbb{E}\left[\left\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s}-1})\right\rVert^{2}\right]\Bigg)$ (116d)
$\leq\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\sigma^{2}+\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\cdot$
$\bigg(G^{2}\mathbb{E}\bigg[\bigg\lVert\bm{\Theta}_{k}^{t_{s},i_{s}-1}-\bm{\Theta}^{t_{s}-1}\bigg\rVert^{2}\bigg]+\mathbb{E}\left[\left\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s}-1})\right\rVert^{2}\right]\Bigg)$ (116e)
$\leq\frac{\eta^{2}G}{2}R_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\left(\sigma^{2}+\beta^{2}\mathbb{E}\left[\left\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\right\rVert^{2}\right]+\kappa^{2}\right)$
$+\frac{\eta^{2}G^{3}}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\mathbb{E}\bigg[\bigg\lVert\bm{\Theta}_{k}^{t_{s},i_{s}-1}-\bm{\Theta}^{t_{s}-1}\bigg\rVert^{2}\bigg],$ (116f)
where (116a) and (116b) use Jensen’s Inequality, (116c) is due to unbiased estimation, (116d) follows Assumption VI.2, (116e) follows Assumption VI.1, and (116f) follows Assumption VI.3.
H-C Immediate Result
Substitute (115c) and (116f) into (114), then substitute Lemma VI.2, we obtain the immediate result for one round,
$\mathbb{E}\left[\left\langle\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1}),\bm{\Theta}^{t_{s}}-\bm{\Theta}^{t_{s}-1}\right\rangle\right]$
$\leq\left(-\frac{1}{2}\eta R_{t_{s}}\lVert\bm{a}\rVert_{1}+\eta^{2}GR_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\beta^{2}\right)\mathbb{E}\left[\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]$
$\hskip 14.22636pt+\eta^{2}GR_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\sigma^{2}+\eta^{2}GR_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\kappa^{2}$
$\hskip 14.22636pt+\left(\eta^{2}G^{3}R_{t_{s}}\lVert\bm{a}\rVert_{1}+\frac{1}{2}\eta G^{2}\right)\frac{2\eta^{2}R_{t_{s}}^{3}\lVert\bm{a}\rVert_{1}^{3}}{1-2\eta^{2}G^{2}R_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}}\cdot$
$\hskip 14.22636pt\hskip 14.22636pt\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\left[\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right],$ (117)
Using Assumption VI.3, this further becomes
$\mathbb{E}\left[\left\langle\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1}),\bm{\Theta}^{t_{s}}-\bm{\Theta}^{t_{s}-1}\right\rangle\right]$
$\leq\left(-\frac{1}{2}\eta R_{t_{s}}\lVert\bm{a}\rVert_{1}+\eta^{2}GR_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\beta^{2}\right)\mathbb{E}\left[\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]$
$\hskip 14.22636pt+\eta^{2}GR_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\sigma^{2}+\eta^{2}GR_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\kappa^{2}$
$\hskip 14.22636pt+\left(\eta^{2}G^{3}R_{t_{s}}\lVert\bm{a}\rVert_{1}+\frac{1}{2}\eta G^{2}\right)\frac{2\eta^{2}R_{t_{s}}^{3}\lVert\bm{a}\rVert_{1}^{3}}{1-2\eta^{2}G^{2}R_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}}\cdot$
$\hskip 14.22636pt\hskip 14.22636pt\left(\beta^{2}\mathbb{E}\left[\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]+\kappa^{2}\right),$ (118)
H-D Average over Successful Rounds
After minor rearrangement, and averaging the immediate result over all $T_{s}$ successful rounds, we get
$\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}\bigg(\frac{1}{2}R_{t_{s}}\lVert\bm{a}\rVert_{1}-\eta GR_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\beta^{2}$-
$\hskip 2.84526pt\frac{2\eta R_{t_{s}}^{3}\lVert\bm{a}\rVert_{1}^{3}(\eta^{2}G^{3}R_{t_{s}}\lVert\bm{a}\rVert_{1}+\frac{1}{2}\eta G^{2})\beta^{2}}{1-2\eta^{2}G^{2}R_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}}\bigg)\mathbb{E}\left[\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]$
$\leq\frac{1}{\eta T_{s}}(\mathcal{L}(\bm{\Theta}^{0})-\mathcal{L}^{\star})+\frac{1}{2}\eta G\lVert\bm{a}\rVert_{1}^{2}\sigma^{2}\cdot\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}+$
$\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}\hskip-2.84526pt\bigg(\hskip-2.84526pt\eta G\lVert\bm{a}\rVert_{1}^{2}R_{t_{s}}^{2}\hskip-2.84526pt+\hskip-2.84526pt\frac{2\eta R_{t_{s}}^{3}\lVert\bm{a}\rVert_{1}^{3}(\eta^{2}G^{3}R_{t_{s}}\lVert\bm{a}\rVert_{1}\hskip-2.84526pt+\hskip-2.84526pt\frac{1}{2}\eta G^{2})}{1-2\eta^{2}G^{2}R_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}}\hskip-2.84526pt\bigg)\kappa^{2}.$ (119)
If the learning rate is chosen as $\eta=\frac{1}{L}\sqrt{\frac{K}{T}}$, by neglecting the non-dominant terms in the coefficients, (119) becomes
$\bigg(\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}\bigg)\cdot\frac{1}{2}\lVert\bm{a}\rVert_{1}\cdot\mathbb{E}\left[\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]$
$\leq\frac{G}{\sqrt{\frac{K}{T}}\cdot T_{s}}(\mathcal{L}(\bm{\Theta}^{0})-\mathcal{L}^{\star})+\frac{1}{2}\sqrt{\frac{K}{T}}\lVert\bm{a}\rVert_{1}^{2}\sigma^{2}\bigg(\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}\bigg)$
$\hskip 14.22636pt+\sqrt{\frac{K}{T}}\lVert\bm{a}\rVert_{1}^{2}\kappa^{2}\bigg(\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}\bigg).$ (120)
H-E Relaxation with Probability Guarantee
By (127), (120) can be further bounded as follows with probability $0.9974$,
$\bigg(\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}\bigg)\cdot\frac{1}{2}\lVert\bm{a}\rVert_{1}\cdot\mathbb{E}\left[\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]$
$\leq\frac{G(\mathcal{L}(\bm{\Theta}^{0})-\mathcal{L}^{\star})}{\sqrt{\frac{K}{T}}\left(\frac{T}{e_{R}}-3\sqrt{\frac{\nu_{R}T}{e_{R}^{3}}}\right)}$
$+\frac{\lVert\bm{a}\rVert_{1}^{2}}{2}\sqrt{\frac{K}{T}}\left(\hskip-2.84526pte_{R^{2}}+3\sqrt{\frac{\nu_{R^{2}}}{\frac{T}{e_{R}}-3\sqrt{\frac{\nu_{R}T}{e_{R}^{3}}}}}\right)\hskip-2.84526pt\left(\sigma^{2}+\kappa^{2}\right).$ (121)
Substitute $e_{R},\nu_{R},e_{R^{2}},\nu_{R^{2}}$, we get
$\frac{1}{T_{s}}\hskip-2.84526pt\sum_{t_{s}=1}^{T_{s}}\mathbb{E}\left[\lVert\Delta\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]\hskip-1.42262pt\leq\hskip-1.42262pt\frac{1}{T_{s}}\hskip-2.84526pt\sum_{t_{s}=1}^{T_{s}}\hskip-2.84526ptR_{t_{s}}\mathbb{E}\left[\lVert\Delta\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]$
$\leq\frac{2G(\mathcal{L}(\bm{\Theta}^{0})-\mathcal{L}^{\star})}{\lVert\bm{a}\rVert_{1}\left((1-P_{O})\sqrt{KT}-3\sqrt{P_{O}(1-P_{O})}\right)}$
$\hskip 14.22636pt+2\lVert\bm{a}\rVert_{1}\sqrt{\frac{K}{T}}(\sigma^{2}+2\kappa^{2})\cdot\Bigg(\frac{1-P_{O}}{P_{O}}\mathrm{Li}_{-2}(P_{O})$
$\hskip 14.22636pt+3\sqrt{\frac{\frac{1-P_{O}}{P_{O}}\mathrm{Li}_{-4}(P_{O})-\left(\frac{1-P_{O}}{P_{O}}\mathrm{Li}_{-2}(P_{O})\right)^{2}}{(1-P_{O})T-3\sqrt{P_{O}(1P_{O})T}}}\Bigg).$ (122)
By retaining only the dominant terms and substituting $\mathrm{Li}_{-2}(P_{O})=\frac{P_{O}(1+P_{O})}{(1-P_{O})^{3}}$, we complete the proof.
Appendix I Renewal Theory
According to the algorithm’s stopping rule, training terminates at the first round in which the global model is successful and the cumulative number of training rounds reaches the threshold $T$. Hence, the total number of the actually executed training rounds is
$T_{s}=\min\left\{n:\sum_{t_{s}=1}^{n}R_{t_{s}}\geq T\right\},$ (123)
where $R_{t_{s}}$ is the number of consecutive training rounds between two adjacent successful global model reconstructions. The $R_{t_{s}}$ renews after each successful reconstruction of the global model, thus $R_{t_{s}}\overset{i.i.d.}{\sim}\mathrm{Geo}(1-P_{O})$ for $\forall t_{s}$. Consequently, we have $e_{R}=\mathbb{E}[R_{t_{s}}]=\frac{1}{1-P_{O}}$ and $\nu_{R}=\mathrm{var}(R_{t_{s}})=\frac{P_{O}}{(1-P_{O})^{2}}$. We can also compute the expectation and variance of $R_{t_{s}}^{2}$, i.e.,
$e_{R^{2}}=\mathbb{E}[R_{t_{s}}^{2}]=\hskip-5.69054pt\sum_{R_{t_{s}}=1}^{+\infty}\hskip-2.84526ptR_{t_{s}}^{2}P_{O}^{R_{t_{s}}-1}(1-P_{O})=\hskip-2.84526pt\frac{1-P_{O}}{P_{O}}\mathrm{Li}_{-2}(P_{O}),$ (124a)
$\nu_{R^{2}}=\mathrm{var}(R_{t_{s}}^{2})=\mathbb{E}[R_{t_{s}}^{4}]-e_{R^{2}}^{2}$
$\hskip 14.22636pt=\frac{1-P_{O}}{P_{O}}\mathrm{Li}_{-4}(P_{O})-\left(\frac{1-P_{O}}{P_{O}}\mathrm{Li}_{-2}(P_{O})\right)^{2},$ (124b)
The polylogarithm function $\mathrm{Li}_{-v}(z)=\sum_{k=1}^{\infty}k^{v}z^{k}$ can be computed by $\mathrm{Li}_{-v}(z)=\left(z\frac{\partial(-\ln{(1-z)})}{\partial z}\right)^{v}\frac{z}{1-z}$.
By examining the step-by-step proof of the asymptotic properties of the standard renewal process, we find that these properties also hold for the expression in (123). Therefore, $T_{s}\approx\frac{T}{e_{R}}$. Then, according to CLT,
$\frac{T_{s}-\frac{T}{e_{R}}}{\sqrt{T}}\overset{d}{\rightarrow}\mathcal{N}(0,\frac{\nu_{R}}{e_{R}^{3}}),$ (125)
where $\overset{d}{\rightarrow}$ indicates convergence in distribution.
According to the strong law of large numbers, the empirical average of any function $f(R)$ over the first $T_{s}$ renewals converges to its expected value almost surely, i.e., $\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}\rightarrow\mathbb{E}[R_{t_{s}}^{2}]$. According to CLT,
$\sqrt{T_{s}}\left(\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}-e_{R^{2}}\right)\overset{d}{\rightarrow}\mathcal{N}(0,\nu_{R^{2}}).$ (126)
By calculating $\mathrm{cov}\left(\frac{T_{s}-\frac{T}{e_{R}}}{\sqrt{T}},\sqrt{T_{s}}\left(\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}-e_{R^{2}}\right)\right)$, we find that its value converges to 0, and the variance of the product decays as $\mathcal{O}(\frac{1}{T})$. Thus, for large $T$, (125) and (126) are nearly uncorrelated, although they are both generated from $R_{t_{s}}$. Since they are jointly Gaussian, this also implies asymptotic independence. Let $\mathcal{C}_{1}\triangleq\left\{\frac{T_{s}-\frac{T}{e_{R}}}{\sqrt{T}}>-3\sqrt{\frac{\nu_{R}}{e_{R}^{3}}}\right\}$, $\mathcal{C}_{2}\triangleq\left\{\sqrt{T_{s}}\left(\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}-e_{R^{2}}\right)<3\sqrt{\nu_{R^{2}}}\right\}$. According to $3\sigma$-rule, we have
$\mathrm{P_{r}}\left(\mathcal{C}_{1},\mathcal{C}_{2}\right)\approx 0.9987^{2}\approx 0.9974$ (127)
Equivalently, $\mathcal{C}_{1}=\left\{T_{s}>\frac{T}{e_{R}}-3\sqrt{\frac{\nu_{R}T}{e_{R}^{3}}}\right\}$, and $\mathcal{C}_{2}\triangleq\left\{\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}<e_{R^{2}}+3\sqrt{\frac{\nu_{R^{2}}}{T_{s}}}\right\}$.